A Complete Bibliography of Publications in Archimedes: New Studies in the History and Philosophy of Science

and Technology

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

02 May 2020 Version 1.03

Title word cross-reference

β [Fra05]. F [Jäh19j].

-sum [Jäh19j].

15th [Bal06b, Cha06a]. 1620s [Ari06]. 1672- [Ano05b]. 16th [Mes13m, NB06]. 17th [DGP99, Kal06]. 1830s [Goo17, JM17q]. 1840s [FL15, JM17c, JM17n, JM17q]. 1890s [FL15]. 18th [AA06, Bal06a, BS06, Del06, DGP99, Dup14, NGRR06, SM06]. 19th [Buc96, Che05, Sib03].

20th [Buc96].

Abbe [Fef97]. Aberdeen [Edd07]. Aberratione [Ano05b]. Abnormality [McC14q]. Abraham [Gin17k]. Abuse [Rut10]. Academic [LF15]. Academies [Pep06]. Acceleration [Ste10]. Accident [Gal00]. Account [Sib03]. Accounting [Pin10]. Achromatic [Sha05]. Acoustical [Pan09a]. Acoustics [Hie14l, Hie14r]. Acquired [dL12a]. Acreage [Jäh19f]. Act [Cam15, Kel15, Vir15a]. Action [Cor04a]. Activities [Hom06]. Activity [Pri14, MT09m]. Acts [CV15a, CV15b]. Actual [Art15]. ad [Jac17]. Adaptation [Swe12]. Additio [dC03]. Adjusting [Jon05]. Adolescence [Jak00]. Adoption [Jon10a]. Aerodynamics [And00, Vin00]. Aeronautical [Has00]. Aeronautics [Cro00]. Aerospace [Jak00]. Aesthetic [Bia16c, HG08a, HG08m]. Affair [Hon97, dL12d, dL12i]. Affinities [Ham17]. After [Hei08a, Hei08b, Jäh19j, Kno07, Mel16]. Age [ML12, Pep06]. Agents [Par03]. Agricola [Edd07]. Agricultural [Hor15, Lav15, Phi15, Wil15]. Agriculture [Arc15, PK15b, Tre15, Vet15]. Air [Cha17e, Kis14]. Aircraft [Rol00]. Airframe [Fer00]. Airplanes [Sch00]. Airport [Dou00]. Airports [Dou00]. al [Nia14h, Ray16i, Ray16k]. al-Dīn [Nia14h]. al-Haytham [Ray16i, Ray16k]. Albertus [Rut19a, Rut19l, Rut19o, Rut19p]. Alchemical [Cha07, Mar14, Nev14, Pri14]. Alchemists [Hau14]. Alchemy [Dup14, Kie14, Mar14, NP05]. Aleksandr [Gor12]. Alexandria [Gou10i]. Alfonsine [CG03h, CG03c]. Algorithms [Che15]. all’Antica [Ben14]. Almagest [Swe05]. Altered [Pan09a]. Ambrose [Hon97]. American [Sow15, Bil00, Cro00, Dou00, Fin15, Ham06, MT09k, Pan09d, Rol00]. Among [Gra17]. Amount [Jon05]. Ampère [Ste03]. Amsterdam [Mel17]. Anachronism [Cha07]. Analysis [Art15, Bor17b, Gla14, Gro15, Hey14, NP05, Rab15, Wes14]. Analyticité [Kno15]. Analyticity [Kno15]. Analyzing [Kel15, Ste03]. Anatomized [Mac17]. Anatomy [Kas18m]. Ancient [Che15, HG08b, Jon10a, JPS16, Ros06, Rit16]. Anderson [Edd07]. Andrea [And14]. Andronov [Gin17a, Gin17r]. Animal [Arc15, Mat15]. Annalen [JM17m, JM17l, JM17n]. Antiquity [Dup14, Jon10b]. Antonio [Hup14]. Antwerp [Mes13a, Mes13m]. Any [SS16]. Apothecary [Kle07]. Apothecary-Chemists [Kle07]. Appear [Jon05]. Appendix [Goe14a, Kas18a]. Application [Duc12l]. Applications [HG08n]. Applied [LF15, Sow15].

Appraisal [Nic06l]. Apprentice [Ack13q]. Approach [Bre03, Vet15]. Approaches [Goe14c]. April [Pri07a]. Aquinas [Rut19a, Rut19l]. Arabic [Nia14e]. Arc [Gin17f]. Archimedes [Cha17b]. Architects [Dou00]. Architectura [Cac14]. Architecture [And14, Cam14, Ger14a, Hey14, HG08m, Kis14, Kla14]. Aristotelian [Rut19n, Rut19p]. Aristotle [Hei08a, Hei08c, Hei08j]. Arithmetic [Mes13a, Mes13b, Ray16b]. Art [Dup14, Kie14, Mar14, Mes13c]. Artificial [Hau14, Tre15]. Artificialis [Ray16l]. Artillery [Bre06]. Artisanal [Ber14b]. Artisans [Pri14]. Artistic [Nev14]. Aspects [Kie14, McC14a, Dou00]. Assessment [Gla14]. Association [HG08b]. Assyriologists [Pro16]. Astorga [Goe14a]. Astrologica [Rut19r]. Astrological [Rut19p]. Astrology [Rut19c, Rut19a, Rut19b, Rut19e, Rut19f, Rut19n, Rut19o, Rut19q, Rut19r]. Astronomer [Cam03]. Astronomers [Ste00r]. Astronomiae [Rut19e, Rut19l]. astronomicae [Rut19l]. Astronomical [CG03a, Nia14f, Rot11, Ste16, Tih10]. Astronomy [Ari06, Bar11, CG03h, Cha06a, NB06, Oss16, Ros06, dJ16]. Athenian [YKS14]. Atmospheric [GR00]. Attitudes [Ros06]. Audiencia [AA06]. Augustine [Jac17]. Austria [Joh06]. Austrian [dS10]. Autism [McC14b, McC14c]. Autonomy [Sch06a, Sch06c]. Aversion [Gla14]. Aviation [Bil00]. Axiom [Boa17]. Axiomatic [Cor04a]. Axiomatization [Cor04b, Cor04d]. Axioms [Bia16a].

Babylonian [Oss16, Ste00o, dJ16]. Back [Ack13a, Ano97a, Ano00a, Ano05e, Ano05f, Ano06a, Ano06b, Ano06c, Ano07a, Ano09a, Ano10a, Ano10b, Ano11a, Ano14a, Ano15a, Ano15b, Ano17a, Ano17b, Bia16b, CG03b, Cha17a, Cor04c, Duc12a, Gin17b, Gin17c, Goe14b, Gol19a, Gou10a, Hei08d, HG08c, Jäh19a, JM17a, Kas18b, Kim02a, MT09a, McC12a, Mes13d, Nia14a, Nic06a, Pan09b, Ray16a, Rut19d, Siv15a, Ste00a, Wüt11a, dL12b]. Background [Cha17i, Cor04l, McC14c, Nia14g]. Bacon [Pri07a, Rut19b, Rut19e, Rut19o, Rut19q]. Bacteriology [Fin15]. Bad [Smi05]. Balance [Cha17h]. Baldassarre [Hup14]. Ballistics [Mes13e]. Barcelona [NGRR06]. Barrow [Pro15a]. Bartolomeo [Cri17]. Based [Che15]. Basis [Hie14t]. Bay [Hub17]. Be [Jon05]. Beat [Gre05]. Beauty [Cam14, McO17]. Beddoes [Lev07, Ste17]. Bedford [McC14f]. Before [Dou00, JM17g]. Beginning [Kim02b]. Beginnings [Jak00]. Being [Rab15]. Beliefs [Ray16j]. Bench [HRR03]. Beneficent [Swe12]. Berlin [Hie14a, Hie14c, Hie14m, Hie14p]. Better [Gli15]. Between [Boa17, Bro11, Goe14c, GBR15b, MT09c, MT09n, McC14o, Pec06, Rab15, Rot11, Smi10, Šub10, dL12c, dS10, GBR15a, SG06]. Beyond [BF05a, Oss16, Cha17b]. Biancani [dC03]. Bible [Rup12]. Biblical [Jac17]. bibliothecae [Mac17]. Biconvex [Bor17b]. Big [Ede17]. Bildungsbürger [Ano09g]. Binocular [Ray16i, Ray16o]. Bioenergetics [Ack13q]. Biographies [Goe14c]. Biological [Tre15]. Biran [Sch10]. Biringuccio [Ber14b]. Blaming [Cro00]. Blasius [Mel17]. Bloch [Gin17k]. Blondel [Gin17k, Gin17f]. Bodies [Cha17h]. Body [Gou10h]. Boerhaave [Boa17, Kno07, Pow07]. Bogolyubov [Gin17n]. Bologna [Ben14]. Bones [Cer99]. Book [Kle10, MT09p]. Books [Lev07, Mes13a, Nev14]. Booties [dL12g]. Borel [MT09o]. Born [Ede17]. Borodin [Gor12]. Borrowed [Duc12h]. Botanical [Nic06c, Nic06b, Nic06d, Nic06h, Nic06j, Wil15]. Botanists [Nic06d]. Botany [SM06, Wil15]. Bottari [Gla14]. Boundaries [Loe03, dS10]. Boyle [Cha17c]. Bradley [Pin10]. Brahe [Hei08h].

Brandenburg [Kin06]. Brazil [Cam03]. Breathing [Kis14]. Breeding [Der15, Gli15, Har15, OP15]. Brie [Ack13m]. Brief [Duc12m]. Bring [Gou10i]. Britain [Buc96, Has00, McO17]. British [Coc06, Ole97, Smi10]. Brothers [McC14o]. Bruckner [Hie14u]. Building [BS06, Ben14]. Buitenzorg [Wil15]. Buried [RD03]. Business [Duc12i]. Buying [Jac97].

c [Smi10]. ca

[BHZ17, Goe14h, Kle10, Rut19r]. Caius [Gra17, Nut17]. Calculating [Kis14]. Caledonia [Woo06]. California [Pri07a, Pri07a]. Cambridge [McC12b, Gra17]. Campaign [Bon11c, Fin15]. Can [McO17, Pan17, SG06]. Canada [Zel17]. Candide [Woo06]. Cannot [Jäh19j]. Cantabrigian [Sch06b]. Cape [Goo17]. Capital [Hie14a]. Caracas [AA06]. Care [Siv15e, Siv15f]. Career [Cor04b, Mau12]. Careers [JM17c]. Carl [Fef97, Rhe03]. Cartan [Gin17u]. Carter [Ray16q]. Cartesian [Ano05j]. Case [Gol19b, Gol19m, Gol19u, Gol19v, Gol19w, Gol19y, Gol19z, Gol19-27, Gol19-28, Gol19-29, Gol19-30, Gol19-31, Gol19-32, Gol19-33, Gol19-34, Gol19-36, Gol19-37, Gol19-38, Gol19-39, Gol19-40, Gol19-41, Gol19-42, Gol19-43, Gol19-44, Gol19-45, Gol19-47, Gol19-48, Gol19-49, Gol19x, Gol19-50, Gol19-51, Gol19-52, Gol19-53, Gol19-54, Gol19-55, Gol19-56, Gol19-58, Gol19-59, Gol19-60, Gol19-61, Gol19-62, Gol19-63, Gol19-64, Gol19-65, Gol19-66, Gol19-67, Gol19-69, Gol19-70, Gol19-71, Gol19-72, Gol19-73, Gol19-74, Gol19-75, Gol19-76, Gol19-77, Gol19-35, Gol19-78, Gol19-80, Gol19-81, Gol19-82, Gol19-83, Gol19-84, Gol19-85, Gol19-86, Gol19-87, Gol19-88, Gol19-89, Gol19-91, Gol19-46, Gol19-57, Gol19-68, Gol19-79, Gol19-90, Gol19c, Gol19d, Gol19e, Gol19f, Gol19g].

Case [Gol19h, Gol19i, Gol19j, Gol19k, Gol19l, Gol19n, Gol19o, Gol19p, Gol19q, Gol19r, Gol19s, Gol19t, Hor15, MT09b, Rut10, SM06]. Cases [Che15, Pro15a, Šub10, Gol19-96]. Casino [Ber14a]. Cassirer [Bia16h, Bia16j]. Catalogus [Mac17]. Causality [Hei10a]. Causes [Ano09d, Duc12h, Hie14s]. Causis [Rut19p]. Cavendish [Kim02b, Kim02c, Kim02f, Kim02g, Kim02h, McC14d, McC14f, McC14o, McC14s, McC14u]. Cavendishes [McC14e]. Cazre [Pal03]. Celestial [Bro11, Cam03, Hat11]. Cells [Par03]. Censored [dC03]. Centuries [Bro11, DGP99, Ray16o]. Century [Bal06b, Cor04l, Dup14, Lee15, Mes13m, NGRR06, Gol19-96, Nic06d, And00, AA06, Bal06a, Bar11, BS06, BVL07, Bos10, Buc96, Cha06a, Cha17l, Che05, Del06, Dij05, Dij17, Fei03b, GR00, Güt15, Ham06, Jac17, Jon10b, Kal06, Kim12, Kle07, NB06, OP15, Pan09a, Phi10, Pri07a, Pri07b, SM06, Sib03, SCD99, Siv15e, Sow15]. Certainty [Sch09]. Ceti [Hat11]. Chain [RD03]. Challenge [Cha17l]. Change [Bon11a, Bro11, Hat11, Ano09d]. Changes [Pri07b]. Changing [NP05, Sup00]. Characteristics [dL12a]. Characterization [JM17s]. characters [Teb15]. Charles [Buc12a, Buc12b, ML12, Mau12]. Chemical [Boa17, Bre06, Edd07, Ede17, JM06, Lev07, MJ06]. Chemicals [Cha06b]. Chemist [Gor12]. Chemistry [And17, BS06, BVL07, Bro06, Kni17, Kno07, Lev07, Pan06, Pow07, Pri07a]. Chemists [Kle07]. Chevreul [Ari06]. chez [Kno15]. Chicken [Der15]. Chief [Tod03]. China [BE15, Gol19-96, Kim12, Lav15, Siv15e, Ste00b]. Chinese [Che15, Siv15f, Ste00c, Ste00s, Teb15]. Christiaan [Dij05]. Christoph [Gor03]. Chromatic [Sha05]. Chronology [Duc12m]. Chymical [NP03]. Chymistry [Pri07b]. Chymists [Pri14]. Circle [Hie14m, Hie14p]. Circles [Pri14]. City [Dou00]. Civil [Ano06d]. Clapham [McC14f]. Classic [Gin17t]. Classical [Ano09b, Ano09h, Sch09, Siv15b]. Classification [Rob15]. Classroom [McV17]. Clausewitz [Lan17]. Clausius [JM17h]. Climate [FL15]. Coignet [Mes13f]. Cold [dL12c, Gol19-96, dL12d]. Coleridge [Sch17]. colleagues [MT09j]. Collection [Mac17]. Collectors [Pan17]. Combination [Hie14r]. Comets [Hei08c, Hei08f, Hei08e, Hei08g, Hei08j, Hei08k]. Commentaries [Jac17]. Commentary [CG03a, Rut19p]. Commercial [Mes13m]. Common [McC14f]. Community [Fin17, Ham06, MT09k]. Comparability [Lev15]. Comparative [Joh06, MJ06]. Comparison [Nia14h]. Comparisons [Sch00]. Complex [Der15]. Compte [Ack13m]. Computer [GM03]. Conceive [Jäh19j]. Concept [Bia16g, Cam14, Cha17f, Gin17a, Gin17s, Goe14e, HG08b, HG08j, HG08d, Kas18m, Ste06b]. Conception [Ano09d, Ano09b, Ano09c]. Concepts [Hau14, Pot06]. Concluding [JM17b, Nia14b]. Conclusion [Ano05g, Duc12b, Jäh19b, Kas18c, McC14g, Mes13g, ZH17a]. Conclusions [Ack13b, Siv15c]. Conditions [Ano09d]. Conference [Gin17d]. Confession [Rot11]. Configuration [Nia14h]. Configurations [Rut19f]. Conflict [Fei03a]. Conjecture [Ray16h, Ray16q]. Connectedness [JM17h]. Connecting [JM17c]. Connection [BB16, Rot11]. Connections [JM17e, Zel17]. Conrad [Bla17]. Consensus [Ole97]. Conservation [Hie14g]. Conservative [Gre05]. Consilia [Cri17]. Consolidation [Jäh19d]. Constructing [Pan09c, SCD99]. Construction [Duc12k, Nic06b, Nic06d, Ray16b]. Contemporary [Tih10]. Content [Nia14e, Nic06c]. Contents [Loe03].

Context [Ara06, Cah97, CG03d, HH06, Kas18j, Kas18k, Nic06l, Pan17, Ric06, Sch06b, SS06a, SS06b, Zel17]. Contexts [HG08b, Pot06, Ric06]. Continua [Cha17m]. Continuity [Bar11, Bon11a, Bro11, Hat11]. Continuous [Cac14]. Contours [Ano09e]. Contributions [Pri07a]. Copenhagen [BB16, Jäh19d, Jäh19c]. Copernican [Bon11c, Wil12]. Copernicus [AA06, Ros06]. Copies [Mil17]. Coproduction [Wil15]. Copying [Mil17]. Cornelius [Tes11]. Correns [Rhe03]. Correspondence [Goe14h, Jäh19d, Jäh19e, Jäh19i, Jäh19k, Pro16]. Cosmic [Hat11]. Cosmological [Bar11, Bro11]. Cosmology [Bon11a]. Cosmos [Hei08f, Rot11]. Cotton [Hor15]. Coulston [Buc12a, Buc12b]. courbes [Kno15]. Creation [Cah97]. Creativity [Kim02h]. Credibility [Buc96, Hon97]. Crisis [NGRR06]. Criticism [Jon10b]. Crossroads [Gin17l]. Cultivated [Güt15]. Cultivation [Hor15, NB06]. Cultural [MT09c, Phi10]. Culture [Eam17, Rut19c, Sch00, Woo06, Zel17]. Cultures [Lee15, Mil17, Ole97]. Cuneiform [AS16, Pro16, Ste16]. Cunitia [Swe12]. Cunitz [Swe12]. curve [Kno15]. Curvilinear [Ray16h]. Cycle [Ack13l, Ack13m, Ack13o, GM03].

Dīn [Nia14h]. damage [Gol19-96]. Dangin [AS16]. Dante [Cer99]. Dark [Jac97]. Data [Sha03, Sup00]. Dati [Goe14h]. Daubenmire [Sow15]. David [Cor04d, Gra11]. Death [Fin17]. Debate [Ber14b]. Decay [Fra05]. Dedication [Bla17]. Deductions [Duc12c]. Defending [Gou10b, Rut19e]. Definition [HG08j]. Definitions [Bia16a, Siv15h]. Degree [Rob15]. del [Ben14]. Della [Bel17, Bor17b, BHZ17, Eam17, Smi17, ZH17b, Rut10, Hee17]. Demonstrative [Duc12c]. Denmark [BB16]. Descartes [Cam14, Cha17d, Gre05, Hee17, Wil12]. Describing [Che15]. Design [Dou00, Fri14]. Designs [Dou00, Gin17k]. Desire [Hir17]. Determinant [Nia14e]. Development [JM17q, Kim02g, Rol00, Sch06a, Wüt11b]. Developments [Mel16]. Diagrams [Bor17b, Hee17, Wüt11g, Wüt11h, Wüt11i]. Did [Har15, HG08k]. Differential [Rob15]. Difficulty [Rab15]. Digital [ML12]. Dilemmas [JM06]. Dilthey [Azz10, Pin10]. Dirac [Wüt11c]. Direct [Ack13m, Ack13n]. Directions [Hie14p]. Directorship [Kim02i]. Disarmament [JM06]. Disciplinary [Rut19f, dS10]. Disciplines [Rut19f]. Discourses [Rup12]. Discovering [Hat11, Pot06, Ste06b]. Discovery [Ara06, Frä12, GM03, How06, HH06, Nic06l, Pec06, RD03, SS06b, Sch06c, SG06]. Discussion [Bia16c, Blo00, Gra11, Ste00d]. discussions [Gol19-96]. Dismal [Sch17]. disorders [Gol19-96]. Dispersion [Sha05]. Displays [Pan17]. Dissertation [Jäh19j]. Distinction [How06, Pec06, Phi10, Sch06b, SS06a, SS06b, SG06]. Distinctions [Phi10]. Disunity [BF05a]. Diversity [OP15]. Divination [Rut19a]. Divine [Kas18l]. Division [Mar14]. Doctrine [Hei10b]. Dome [Gla14]. Dr. [Lev07]. Draughtsmen [Nic06d]. Drawing [Güt15]. Drawings [Hup14]. Drift [Ano06o]. Dryland [Vet15]. Duhem [Sch06a]. Dunthorne [Ste10]. DuPont [Smi06]. During [AA06, BS06, Cha06a, Del06, DGP99, Gav99]. Durkheim [Sch10]. Dutch [Hom06, Ver11]. Dutch/Shell [Hom06]. Dwight [Wes14]. Dyeing [Mar14]. Dynamo [Gin17f].

Early [Bon11a, Cor04b, Hie14b, HG08b, Kel15, Kla14, McC12c, Pri07b, Row16, Ste00r, Buc96, FL15, Arc15, Che05, Dij17, FNB06, Lee15, Rut10, Sib03]. Earn [dL12g]. Earth [JM17p]. East [Ste00m]. Eccentricity [McC14h]. Ecclesiastical [Gra17]. Eclipse [Ste00o, Ste00r, Ste00s].

Ecliptic [Swe10]. Ecole [GG12]. Ecological [Ack13c, Ack13n, Ack13o, Vet15]. Ecologist [Ack13s]. Ecology [Ack13m, Ack13n, Ack13p, Sow15]. Economics [Sch17, dS10]. Ecosystems [Pot06]. Edenda [dC03]. Edge [Sha05, Ste17]. Education [Ano06d, Goe14e, NGRR06]. Educational [Phi10]. Edward [Ano06d, Ham06]. Effect [Jäh19f, Ste17]. Efficacy [Siv15g]. Effort [Cha06b]. Egypt [Rit16]. Ehrenfried [Wes14]. Eighteenth [BVL07, Nic06d, Kle07, Pri07a, Pri07b, SCD99]. Eighteenth-century [Nic06d, Kle07, Pri07a, Pri07b, SCD99]. Einstein [Bia16h, Cor04e, RS03]. Electricities [Ste06b]. Electrodynamics [Wüt11i]. Electromagnetic [Cor04h]. Élémentaires [Sch10]. Elementorum [Rut19p]. Elements [Boa17, Gou10h, Pow07, Cha17b]. Eleventh [Siv15e]. Eleventh-Century [Siv15e]. Elimination [Jes15]. Elite [Siv15i]. Embodied [Are10]. Embryology [Hir17]. Emergence [Has00, Kim02c, SM00]. Empire [LF15]. Empiricism [Ueb10]. Encounter [Hie14c]. Encounters [Hie14d]. End [Kim02d]. Energy [Ack13d, Hie14g]. Engagement [Cha17d]. Engine [SM00]. Engineering [And00, Fer00, Gin17q, Vin00]. Engineers [Dou00]. England [McO17]. English [Goe14a]. Enharmonium [Hie14e, Hie14f]. Enlightenment [Gav99, Pes06]. Enochian [Mar14]. Entomology [LF15]. Enumeration [Teb15, Vir15b]. Enumerations [Rob15]. Environment [Hub17]. Epilogue [Cor04f]. Epistemological [Ber14b, Gla14, Phi10]. Equation [Gin17u, Wüt11c]. Equations [Rob15]. Equilibrium [Cha17o]. Equinoxes [Swe10]. equipollence [Kno15, Kno15]. Era [Bor06, Hei08g, Kim02d, Fin15]. Erich [Gli15]. Erklären [Are10, Fee10b, Fee10a, Lea10]. Ernst [Cha07, Fef97]. Error [Buc05]. Errors [Ray16p, RS03]. Erya’s [Teb15]. Especial [GG12]. Essays [BS17, MK17b, Buc12a]. Establishing [JM17d]. Estimates [Jon05]. Ether [Ano06o]. Euclid [Gou10f]. Euclidean [Bia16d, Bia16h]. Euler [HG08k]. Europe [Piñ06, Rut10, Ste00p]. European [Gav99]. Evaluating [Hie14f]. Eve [MT09i]. Everyday [McC14i]. Evolution [And00]. Exchange [Ack13d, Fer00]. Exchanges [MT09o]. Existence [Gin17u]. Expanding [Pan09d]. Experience [Coc06, Hie14b, Kle10, Pin10]. Experiences [Sib03]. Experiment [Cha17j, Lav15, Vin00]. Experimental [Bre03, Loe03]. Experimenting [Cha17e]. Experiments [Eam17, FL15, Rhe03]. Expert [Hie14f]. Expertise [Bre06, Hor15, Ste06a]. Explaining [Mül10]. Explanation [Hei10a, Sch10]. Exploring [Loe03]. Explosives [Coc06, Ste06a]. Extractions [Kel15]. Extraordinary [McC14s]. Extreme [Hub17]. Eye [Ano03c].

F [Sow15]. Fabri [Pal03]. Fabricius

[Gra11]. faciles [Swe12]. Facing [Duc12c]. Fact [Ray16c]. Factory [MJ06]. Facts [And17]. Faculty [Piñ06]. Family [McC14d, Mes13f]. Fantasies [And17]. Fantoni [Rut10]. Faraday [Pan09e, Ste03]. Fashioning [Cha17f]. Fate [Rut19a]. Feeding [Mat15]. Fellow [Pro16]. Feltre [Goe14e, Goe14d, Goe14i]. Fertilizer [Jäh19f, Tre15]. Feynman [Wüt11c, Wüt11d, Wüt11g, Wüt11h, Wüt11i]. Fiction [Ray16c]. Field [Wüt11d, Zel17, dJ16, Jak00]. Fienus [Hir17]. Fifteenth [Bar11]. Filippo [Rut10]. Fire [Boa17, Dij17, Ede17]. First [Gin17d, Gin17k, Kim02f, Lav15, Coc06, Pri07a]. Fishing [Hub17]. Fit [Jon05]. Five [Rup12]. Flashback [Jac17]. Fleming [Hon97]. Flight [GR00, Jak00, SS16, Sup00]. Floating [Cha17h]. Florence [Kie14]. Food [Sch17]. Forerunner [Sch06b]. Forest [How06]. Forestry [FL15]. Forests [FL15]. Forgotten [RD03]. Form [Hau14]. Formation [Duc12l, Ste06b]. Formes [Sch10]. Formulation [Jäh19d]. Fortifications [Mes13e]. Fortschritte [JM17l]. Foundation [Ras06, Sch06c]. Foundations [Cor04g, JM17e, Rut19m, Rut19p]. Fourfold [Mar14]. Frame [Jon10a]. France [BVL07, Gin17m, Gin17p, Gin17r, Hor15, MT09c, MT09i, OP15, Sch10]. Francesco [Bel17]. Francis [Pri07a]. Franck [Jäh19f]. Francois [AS16]. Fraunhofer [Jac97]. Frederick [McC14o]. Frederik [Mel17]. Free [Ack13e, Wüt11e]. Freedom [Ric06]. French [Bre06, Cha06b, HG08m, Kno15, Lev07, MT09d, MT09j]. Fresco [Ray16c]. Fritz [Jäh19j]. Front [Ack13f, Ack13g, Ack13h, Ack13i, Ack13j, Ack13k, Ano97b, Ano99, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano03a, Ano03b, Ano05h, Ano05i, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano07b, Ano09f, Ano10c, Ano10d, Ano11b, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano16, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Bia16e, CG03e, Cha17g, Cor04i, Duc12d, Duc12e, Duc12f, Duc12g, Gin17g, Gin17h, Gin17i, Gin17j, Goe14f, Gol19-92, Gol19-93, Gol19-94, Gou10d, Hei08i, Hie14h, Hie14i, Hie14j, Hie14k, HG08e, HG08f, HG08g, HG08h, Jäh19g, JM17f, Kas18d, Kas18e, Kas18f, Kas18g, Kas18h, Kim02e]. Front [MT09e, MT09f, MT09g, MT09h, McC12d, McC12e, McC12f, McC14j, McC14k, McC14l, McC14m, Mes13h, Nia14c, Nic06e, Pan09f, Ray16d, Ray16e, Ray16f, Ray16g, Rut19g, Rut19h, Rut19i, Rut19j, Rut19k, Siv15d, Ste00e, Ste00f, Ste00g, Ste00h, Ste00i, Ste00j, Wüt11f, dL12e].

Frontier [Ham06]. Frontline [MJ06]. Fronts [Hom06]. Functions [Gin17e]. Fundamental [Sch06b].

G. [GBR15a]. Galileo [Cam14, Cha17h, Fei03a, Lev15, Pal03, RD03, Wil12, dC03]. Gallery [Kie14]. Galvani [Bre03]. Garden [Wil15]. Gases [Ste06a]. Gauging [Mes13o]. Geisteswissenschaft [Phi10]. Gemma [Tes11]. General [Bia16h, RS03]. Generalising [Kas18l]. Generality [Rob15]. Generation [Goe14e]. genesi [Jac17]. Genesis [Old12, RS03, Wüt11g]. Genetic [Der15, dL12d, Har15]. Genre [BS06]. Genres [BE15]. Geography [Güt15, Mit10, Phi15]. Geology [Goo17, Old12, Edd07]. Geometric [Fri14]. Geometrical [Ger14a, Hee17]. Geometrie [Cor04d]. Geometries [Bia16d]. Geometry [Bia16g, Bia16h, Bia16j, Hey14, YKS14]. Georg [Cha07]. George [NP03]. Georgics [Edd07]. Gérard [Gin17f]. Gérard-Lescuyer [Gin17f]. Gerardus [Mel17]. German [Hie14a, Bos10, JM17g, Kle10, Nev14, Ole97, Tre15, dS10]. Germano [Sch06b]. Germany [Buc96, Joh06, JM17r, JM17s, Kle07, Mat15, Phi10, Phi15, SS16]. Gessner [Bla17]. Giambattista [Bel17, BHZ17, Eam17, Smi17, ZH17b]. Gillispie [Buc12a, Buc12b, ML12, Mau12]. Giovanni [Gla14, Rut10]. Giuseppe [dC03]. Glass [Jac97]. Glossary [CG03f]. Goals [Wüt11b]. God [Duc12j]. Goldsmiths [Hau14, Pri14]. Good [Che05]. Göttingen [JM17p, Row16]. Governmental [LF15]. Granada [AA06]. Grassi [Fei03a]. Graz [Row16]. Great [McC14s, Wüt11c, Gin17k, Vet15]. Greco [Ste00k]. Greek [DGP99]. Grey [McC14d]. Grienberger [Fei03a, Gor03]. Ground [Sup00]. Grounds [Fei03a]. Group [Hie14c]. Growth [Cro00]. Grundlagen [Cor04d]. GTR [Cor04j]. Guglielmo [Hon97].

H [BS17]. Hadamard [Gin17l, MT09o]. Half [Cha06a, NB06]. Hall [McC14e]. Hamiltonian [Wüt11j]. Handy [Swe05]. Hanging [RD03]. Hans [GM03, Sch06c]. Happen [HG08k]. Harmonium [Hie14o]. Harmony [Hie14s, HG08b]. Hauck [Ray16h]. Hay’a [Nia14e]. Haytham [Ray16i, Ray16k]. Health [Siv15e, Siv15f]. Heart [Gre05]. Heaven [Kim12, Kim12]. Heavens [Nia14h]. Helmholtz [Ano09d, Ano09g, Ano09h, Ano09i, Bia16f, Bia16j, Hie14c, Hie14l, Hie14m, Pan09g, Sch09]. Help [Kel15]. Henry [Gou10i, McC14o, McC14s]. Henseleit [GM03]. Herculaneum [Kle12]. Heritage [Ano06l, Bil00]. Herman [Pow07]. Hermann [Pan09g, Sch09]. Hermeneutic [Kle10]. Herschel [Goo17]. Heuristic [Nic06l]. Heyns [Mes13l]. Hieremiae [Mac17]. High [Ste06a]. Hilbert [Cor04b, Cor04d, Cor04e, Cor04j]. Hippocrates [Kno07]. Hippocratical [Kno07]. Hippolyte [Mül10]. Historian [Nut17]. Historians [Gra17]. Historical [Cha17i, Fee10b, Fee10a, Mül10, Old12, SS06b, dS10]. Historicism [Bos10, BF05a]. Historiography [Høy16]. History [Arc15, Buc12a, CV15a, CV15b, FL15, Gal00, Gil12, Gin17m, Goe14c, Gou10b, Gou10g, Gra17, Hei08j, HRR03, HG08o, JM17r, Kim02h, MK17b, Mau12, Mül10, Old12, PK15b, Sch06b, SS16, Zel17]. Holker [McC14e]. Holwarda [Ver11]. Homchen. [Azz10]. Home [McC12g]. Honor [MK17b, Buc12a]. Honore [Pal03]. Honour [BS17]. Hope [Goo17]. Houses [McC14f]. Hudson [Hub17]. Human [Kas18j, Kas18k, Lea10]. Humanism [Goe14d]. Humanist [Goe14c]. Humanitatis [Goe14d]. Humanities [Smi10]. Hund [Jäh19f]. Hundred [Cha17f, Cha17n]. Huygens [Dij05]. Hydrostatics [Cha17b, Cha17d, Cha17i, Cha17j, Cha17k, Cha17m, Cha17n].

Hypatia [Gou10b, Gou10h]. Hypotheses [Bia16a, Swe05]. Hypothesis [Ano06l, Ray16b]. Hypothetization [Ano09i].

Ibn [Ray16i, Ray16k]. Ideas

[Bro11, Gin17l, Ham06, Kim12, Sch06b]. Igne [Wes14]. II [Kas18j, Dou00, Duc12l, Sch00]. III [Duc12m, Kas18k]. Illa [dC03]. Illustrated [Mel17]. Illustrations [Nic06c, Nic06b, Nic06d, Nic06h, Nic06j]. Images [NG99]. Imagination [Hir17]. Imagines [Rut19l]. Immanuel [HG08k]. Impact [Bro11]. Imperial [BE15]. Implementation [Jäh19i]. Implications [Smi17]. Imponderable [Boa17]. Importance [Kno07]. Imprecisions [Rag10]. Improvement [Fin15, Lav15]. Incompatibility [Gin17q]. Individuality [Bos10]. Inductive [Sch06c]. Industrial [Cha06b, Kin06]. Industry [Kim12, Lee15, MJ06, Pan06]. Inextricability [Ara06]. Infinite [Art15, Lev15]. Infinitesimals [Jes15]. Infinities [Lev15]. Influenced [Azz10]. Influences [Pan09h]. Infra [Loe03]. Infra-Red [Loe03]. Innovation [Buc17, Gre05, Hor15]. Inoculation [Fin15]. Insect [LF15]. Insights [Pro16, RS03]. Instead [Lea10]. Institute [Pri07a]. Institutional [Rut19m, MT09m]. Instructional [Vir15a]. Instrument [Cah97, Mes13i]. Instrumentation [Sup00]. Instruments [Boa17, Dij17, Pow07]. Integrate [Che15]. Intellectual [Row16, Ano06l]. Intensities [Jäh19k]. Intensity [Jäh19c]. Interactions [Rab15]. Interest [McO17]. International [Bil00, Gin17d]. Interpretation [Bia16d, Bos10, Ver11, Wüt11h]. Interpretive [Tod03]. Interrelations [GBR15a, GBR15b]. Interventionism [MT09n]. Interventions [Hau14].

Intonation [Hie14o]. Introduction [Ano05j, Ano06m, Ano06n, Ano09j, Bon11b, Bor17a, BF05a, CG03g, Cor04k, Fee10b, Ger14b, Goe14g, Gol19-95, Hie14n, HG08i, Jäh19h, MK17a, Mes13j, Nic06f, PK15a, SS06a, Ste00l, Wüt11h, dL12d]. Invariances [Hee17]. Invention [Rol00]. Inventor [Pro15a]. Investigating [Ack13l]. Investigation [Ack13q, Sha05]. Investigative [Hol03]. Investigator [Tod03]. Iran [Nia14d]. Isaac [Duc12i]. Islamic [Rag10, Ste00m]. Issue [BE15]. Italian [Bor06, HG08m, MT09n]. Italy [Fin17, Hor15, MT09c, MT09i, Pan06, Pep06]. Itinerary [Duc12m]. IV [Fri14, Kas18l]. Ivan [Tod03].

J [Kim02f, Kim02g]. J. [Kim02f, Kim02g]. Jacques [Ari06]. James [Edd07, Goe14a, Lea10, Mil17, YKS14]. Japan [Ste00n, Arc15, Hie14d, Lee15]. Java [Wil15]. Jesuit [Cam03, Fei03b, Lei03, Pal03]. Jesus [Gor03]. Jewish [MT09k]. Johannes [Gra11, Hie14m, Hie14p, Ver11]. John [McC12l, Goo17, Gra17, Hon97, Nut17, Wes14]. Joseph [Jac97]. Journal [McC14n]. Journals [JM17m]. Journey [Row16]. Journeys [JPS16]. Just [Hie14o]. Justification [Ara06, How06, HH06, Nic06l, Pec06, SS06b, Sch06c, Ste06b, SG06]. Justifying [Pot06].

Kaiser [Ras06]. kaiwu [Kim12]. Kant [Bia16c, Bia16f, HG08k]. Kantianism [Šub10]. Keeping [Sib03]. Kepler [Rot11, Bon11c, Gra11, Hee17, Swe10]. Kindred [Ano06o]. Kirchhoff [JM17h]. Kirchmöser [Kin06]. Kirwan [Old12]. Knowledge [FL15, Fin17, How06, Nev14, Phi15, Ray16j, Rut19r, SCD99, Vet15]. Koenig [Pan09a]. Konopinski [Fra05]. Krebs [GM03]. Kreide [Azz10]. Kries [Hei10a]. Krylov [Gin17n]. Kuhlmann [Lan06]. Kuhn [HH06]. Kurd [Azz10]. Kurt [GM03].

Laboratories [Bre06, Dup14, Kie14]. Laboratory [Ack13e, Ack13l, GM03, Hol03, JM17q, Loe03, NP03, Sib03, Tod03, Wes14, Kim02h]. Lalande [Ste10]. Langley [Loe03]. Language [Nia14e, Nic06k]. Laplacian [Hei08f, Hei08g]. Laßwitz [Azz10]. Late [BS06, Cor04l, McC12h, SM06, BE15, Fin17, Mes13m, Ste00o, Ste00p]. Latitude [Swe05]. Latter [NB06]. Laurel [Mes13l]. Laws [JM17c]. Leadership [Kim02g, Kim02h]. Learn [Pan17]. Learning [Gou10e, Sir17]. Lecture [And17]. Lectures [BVL07, Gin17s, Gou10i, Kal06]. Legacy [CG03h, Hie14l, Ray16k]. Legendre [HG08j]. Legume [Fin15]. Leibniz [Bee15, Art15, GBR15a, GBR15b, Gro15, Jes15, Kno15, Lev15, Pro15a, Rab15]. Lenses [Ano05g, Bor17b, Dij05, Sha05]. Leonardo [Ber14b, Goe14h]. Leonhard [HG08k]. Lescuyer [Gin17f]. Letter [Goe14a]. Letters [McC12i, McC14o, Mil17]. Levere [BS17, Ham17]. Liénard [Gin17u]. Life [Ack13l, Ack13m, Ack13o, Gil12, McC14i, McC14p, Mes13m, Nia14i, PK15b, Sir17, Šub10]. Ligamentum [Rut19n]. Light [Ano05c, Dij17, Frä12]. Likely [Mau07]. Limits [Duc12c, Ste06b]. Lindstedt [Gin17q]. Line [Güt15]. Lineage [Cha07]. Lineages [Gou10e]. Linear [Ray16j]. Lines [Jac97]. Links [Nic06g]. Liquids [Cha17l, Cha17m, Cha17o]. List [Pro15b]. Lists [Teb15]. Literature [BS06, BE15]. litteram [Jac17]. Lives [Gou10f]. Local [FL15, Ole97]. Logical [Ueb10]. Lone [Tod03]. Longitudes [Jon10a]. Longomontanus [Swe10]. Loss [Sch09]. Lost [How06, JM06]. Lowthers [McC14e]. Luc [Old12]. Lucretius [Kle12]. Luigi [Bre03]. Lumière [Ano05d]. Lysenko [dL12d, dL12h].

Machine [Gin17f, Mil17]. Magi [Dij17]. Magic [Jäh19k, Rut19q, Rut19r]. Magician [ZH17b]. Magnetism [JM17p]. Magnitude [Bia16i]. Magnus [Rut19a, Rut19l, Rut19o, Rut19p]. main [Duc12i]. Maine [Sch10]. Make [Jon05]. Makers [Mes13i]. Making [Ano06p, Hie14b, HG08d, Nic06h]. Man [Gra17, Frä12]. Management [Wil15]. Managerial [Tod03]. Managing [Bre06]. Mandel’shtam [Gin17o]. Manner [Kno07]. Mannheim [Šub10]. Manufacture [Coc06, Fer00, Jac97]. Mapping [Güt15, Loe03, Mes13k]. Maps [Mit10]. Marco [Ber14a]. Marconi [Hon97]. Maria [Swe12]. Market [Cha07, Pan09d, Pan09h]. Martii [Mac17]. Masaccio [Ray16c]. Maskelyne [Hon97]. Master [Ack13m, And14, Buc12a]. Masterpiece [Boa17]. Material [Ber14a, Kie14]. Materials [Nic06i]. Maternal [Hir17]. mathematicae [Gou10c]. Mathematical [Bee15, Bor06, Che15, Dij05, Duc12c, Ger14a, Gla14, Gou10b, Gro15, HG08l, Kal06, Kel15, Lei03, Mes13m, Oss16, Pro15b, Pro16, Rut19f]. Mathematician [Bee15, JPS16]. Mathematicians [JM17i, MT09j]. Mathematics [AS16, CL06, GBR15a, GBR15b, Gor03, Gou10g, GG12, HG08n, Høy16, Hup14, Kla14, Mel16, Mes13m, Rab15, Rut19o, SS16, Ros06]. Mathematicus [Goe14d]. Mathematizing [Rut19o]. Matter [Ack13a, Ack13d, Ack13f, Ack13g, Ack13h, Ack13i, Ack13j, Ack13k, Ano97a, Ano97b, Ano99, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano03a, Ano03b, Ano05e, Ano05f, Ano05h, Ano05i, Ano06a, Ano06b, Ano06c, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano07a, Ano07b, Ano09a, Ano09f, Ano10a, Ano10b, Ano10c, Ano10d, Ano11a, Ano11b, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano16, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Art15, Bia16b, Bia16e, CG03b, CG03e, Cha17a, Cha17g, Cor04c, Cor04i, Duc12a, Duc12d, Duc12e, Duc12f, Duc12g]. Matter [Gin17b, Gin17c, Gin17g, Gin17h, Gin17i, Gin17j, Goe14b, Goe14f, Gol19a, Gol19-92, Gol19-93, Gol19-94, Gou10a, Gou10d, Ham17, Hei08d, Hei08i, Hie14h, Hie14i, Hie14j, Hie14k, HG08c, HG08e, HG08f, HG08g, HG08h, Jäh19a, Jäh19g, JM17a, JM17f, Kas18b, Kas18d, Kas18e, Kas18f, Kas18g, Kas18h, Kim02a, Kim02e, MT09a, MT09e, MT09f, MT09g, MT09h, McC12a, McC12d, McC12e, McC12f, McC14j, McC14k, McC14l, McC14m, Mes13d, Mes13h, Nia14a, Nia14c, Nic06a, Nic06e, Pan09b, Pan09f, Ray16a, Ray16d, Ray16e, Ray16f, Ray16g, Rut19d, Rut19g, Rut19h, Rut19i, Rut19j, Rut19k, Siv15a, Siv15d, Ste00a, Ste00e, Ste00f, Ste00g, Ste00h, Ste00i, Ste00j, Wüt11a, Wüt11f, dL12b, dL12e]. Matthew [Gra17]. Maurolico [Bel17]. Max [Hei10a]. Mayer [Ste10]. McCook [Jak00]. Means [Wüt11b]. Measurement [Hes99]. Measurements [Bia16d, Che05]. Mechanical [Cor04h, Kal06]. Mechanics [Cha17l, Gin17n, Hey14, Wüt11j]. Mechanism [Ano09c, Ano09h, Ano09i, Ano09k, Cha17c, Sch09]. Mediaeval [Nev14]. Medical [Bal06b, BE15, Gol19-96, Gra17, Mac17, Ste17]. Medicine [Cri17, Jac17, Kno07, Lev07, MK17b, Piñ06, Siv15b, Siv15f]. Medieval [Kel15, Ste00p]. Meets [Cam14]. Men [Mil17, Mil17]. Mendelism [Gli15, Har15]. Mendelist [dL12f]. Mengoli [Pro15a]. Menstruums [Boa17]. Mesa [Ray16b]. Mesopotamia [Pro15b, Ste00q]. Mesopotamian [Høy16, Mel16]. Metallogenetical [Hau14]. metaphysicae [Rut19n]. Method [Ack13m, Cor04a, Gin17n, Gin17o, Gin17q, Gin17t].

Méthode [Sch10]. Methodenstreit [dS10]. Methodological [Duc12m]. Methodology [Duc12i, Duc12k, Duc12l, Duc12m]. Methods [Ack13n, Der15, JM17j, Nic06i]. Metrical [Bia16g]. Metropolis [Mes13m]. Michell [McC12l]. Michurinism [dL12f]. Microbial [Lee15]. Microbiology [Ack13c, Ack13o, Ack13r]. Microscope [Ack13l]. Microscopical [Fef97]. Mid [BVL07, Sow15]. Mid-Eighteenth-Century [BVL07]. Mid-Twentieth-Century [Sow15]. Mie [Cor04e]. Milestones [Wüt11b]. Military [Bro06, Pan06, Ras06]. Mill [Hei10a, Sch17]. Minkowski [Cor04m]. Mira [Hat11]. Mirandola [Rut10]. Misrepresentation [Buc17]. Missionary [Cam03]. Mistake [Buc05]. Mition [RD03]. Mobilization [Cha06b, Joh06, Ste06a]. Model [Duc12k, Duc12l]. Models [GM03, Nia14h]. Modern [Ano09e, Arc15, Bon11a, Bor06, FL15, FNB06, HG08b, Jak00, JPS16, Kla14, NG99, Ros06, Rut10, Sch09, Pep06]. Modesty [Gor03]. Mold [Lee15]. Mongols [Nia14d]. Montagnana [Cri17]. Montpellier [McV17]. Moody [Ack13q]. Moon [Ste10]. Morality [Buc17]. Morganists [dL12f]. Morley [Ano06d, Ham06]. Morleys [Ano06q]. Morphology [Wil15]. Motion [Pal03]. Motions [Gin17e]. Müller [Hie14m, Hie14p]. Multiplet [Jäh19k]. Multitude [Lev15]. Munitions [Bro06, Joh06]. Muscular [Bre03]. Museums [Pan17]. Music [Hie14b, Hie14d, Hie14t]. Music-Making [Hie14b]. Musical [Hie14s].

Nancy [MK17b]. Naples [Eam17]. Napoleonic [Hor15]. Narrating [Sib03]. Narratives [Pri07a]. National [Hom06, Sch00]. Natural [Arc15, CL06, FL15, Hau14, HG08o, McC14p, Rut19n, Rut19p, Rut19r, Smi10, Duc12i, Kim12].

Natural-Philosophical [Duc12i]. Naturalis [Ray16l, Rut19n]. Naturalis/Artificialis [Ray16l]. Naturalist [Fin17]. Nature [Ack13e, Ack13l, Ano09d, Ano09e, Bre03, Güt15, Kle10, Lan17, Nic06d, Pot06, Sch09, Sup00, Tre15]. Naturphilosophie [Lan17]. Naturwissenschaft [Phi10]. Naturwissenschaft/Geisteswissenschaft [Phi10]. Navigation [Mes13c]. Nazi [SS16]. Near [Ste00m]. Neo [Šub10]. Neo-Kantianism [Šub10]. Network [Jäh19e]. Neugebauer [BB16, Høy16, JPS16, Mel16, Oss16, Pro16, Rit16, Row16, SS16, Ste16]. Newton [AA06, Cam14, Cha17m, Cha17n, Duc12c, Duc12i, Duc12h, Duc12m, Hei08b, Hei08h, Kas18m, Kas18i, Sha03, Sha05, Smi05, Wil12]. Newtonian [Hei08e, Hei08k]. Nicolaas [Mel17]. Nineteenth [Cor04l, Bos10, Güt15, Ham06, Jon10b, OP15, Pan09a, Phi10]. Nineteenth-Century [Bos10, Güt15, Ham06, OP15, Pan09a, Phi10]. ninety [Gol19-96]. Nobody [Pri07b]. Nomenclature [Cam15]. Non [Bia16d, Bia16h, Hei08k, dC03]. Non-Euclidean [Bia16d, Bia16h]. Nonlinear [Gin17d, Gin17m, Gin17n]. Normality [McC14q]. North [Dij17, Pan09d]. Notebook [RS03]. Notebooks [GM03, Hol03, HRR03, Loe03, NP03, Sha03, Tod03]. Notes [Gin17a, Par03, RD03]. Noticed [Pri07b]. Notions [Rab15]. Nova [Gra11]. Novel [Cha17f]. Novelties [Bro11]. Nueva [AA06]. Number [Bia16i, Bia16j, Gol19b, Gol19m, Gol19u, Gol19v, Gol19w, Gol19y, Gol19z, Gol19-27, Gol19-28, Gol19-29, Gol19-30, Gol19-31, Gol19-32, Gol19-33, Gol19-34, Gol19-36, Gol19-37, Gol19-38, Gol19-39, Gol19-40, Gol19-41, Gol19-42, Gol19-43, Gol19-44, Gol19-45, Gol19-47, Gol19-48, Gol19-49, Gol19x, Gol19-50, Gol19-51, Gol19-52, Gol19-53, Gol19-54, Gol19-55, Gol19-56, Gol19-58, Gol19-59, Gol19-60, Gol19-61, Gol19-62, Gol19-63, Gol19-64, Gol19-65, Gol19-66, Gol19-67, Gol19-69, Gol19-70, Gol19-71, Gol19-72, Gol19-73, Gol19-74, Gol19-75, Gol19-76, Gol19-77, Gol19-35, Gol19-78, Gol19-80, Gol19-81, Gol19-82, Gol19-83, Gol19-84, Gol19-85, Gol19-86, Gol19-87, Gol19-88, Gol19-89, Gol19-91, Gol19-46, Gol19-57, Gol19-68, Gol19-79, Gol19-90, Gol19c, Gol19d, Gol19e]. Number [Gol19f, Gol19g, Gol19h, Gol19i, Gol19j, Gol19k, Gol19l, Gol19n, Gol19o, Gol19p, Gol19q, Gol19r, Gol19s, Gol19t]. Numbers [Sib03]. Nutrition [Mat15]. Nymphs [Mes13l].

oberen [Azz10]. Objectivity [SS16]. Objects [Ger14a, Pot06]. Obliquity [Swe10]. Observationes [Mel17]. Observations [JM17b, Ste00r, Swe10]. Obstacles [Gla14]. Office [Vet15]. Oil [Arc15]. Old [Frä12, Jäh19i]. Omnis [Goe14d]. One [Cha17f, Cha17n, MT09o]. Ontology [Are10]. Opening [Kim12]. Opera [Rut19q]. Operating [Hom06]. Operations [Che15, Kel15]. Operator [Wüt11j]. Opinions [Hie14f]. Opposition [Ueb10]. Optical [Bor17b, Jac97, Sha03]. Optician [ZH17b]. Optics [BHZ17, Dij05, Duc12c, Fef97, Ray16o]. Orbs [Bar11]. Order [Jon05]. Ordering [Kel15]. Orders [Kla14]. Ordinary [JM17k]. Organic [Tre15]. Organization [Rob15]. Organizing [Smi06]. Origin [Jäh19d, Wüt11h]. Origins [Ede17, Tre15]. Orthodox [Ueb10]. Orville [Cro00]. Oscillations [Gin17a, Gin17m, Gin17p, Gin17s]. Other [JM17m]. Otto [BB16, JPS16, Rit16, SS16]. Our [Jon05]. Ourselves [Jon05]. Outbreaks [LF15]. Outsider [Høy16]. Oxford [Gou10i]. Oxygen [Ano06r].

Padua [Del06]. Painted [And14]. Palazzo [Ben14]. Panofsky [Ray16h]. Papalexi [Gin17o]. Papendick [Hie14q]. Paper [Blo00, Bor17b]. Papyrus [Tih10]. Paradigm [Gin17p]. Paris [Gin17d, Ari06, Pan09a]. Parker [Gra17]. Pasadena [Pri07a]. Pascal [Cha17o]. Patent [Cro00]. Patents [Rol00]. Pater [Goe14d]. Path [HG08a, HG08l]. Pathways [Hol03]. Paul [Fri14]. Pavlov [Tod03]. Pear [OP15]. Peas [Gli15]. Peasants [Gli15]. Peculiarities [McC14s]. Peking [Frä12]. Pérès [MT09b]. perfect [Ano05j]. Perhaps [Sch06b]. Period [FNB06]. Periodic [Gin17u, Gin17e]. Periphery [Gav99]. Persian [Nia14e]. Person [McC14r]. Personality [McC14a, McC14s]. Perspectiva [Ray16l, Rut19o]. Perspective [Jon10b, Ray16h, Ray16j, Ray16b, Ray16m, Ray16n, Ray16o, Ray16p, Ray16q, And14]. Perspectives [Fee10b, Fee10a, Fei03b, Joh06, MJ06, PK15b, SS06b]. Persuading [MT09k]. Peruzzi [Hup14]. Pest [LF15]. Pesticide [Arc15]. Peter [Gla14, Mes13l]. Peurbach [Bar11]. Pharmaceuticals [Cha06b]. Phase [Duc12k, Duc12l]. Phenomena [Duc12c, Duc12j]. Philological [Pro16]. Philosopher [Bee15]. philosophiae [Rut19n]. Philosophical [Duc12i, Gro15, Rut19p, SS06b]. Philosophy [Ano09e, CL06, Duc12i, GBR15a, GBR15b, McC14p, Mül10, Rut19n, Sch06c, Sch09, Šub10, Rab15]. Phocylides [Ver11]. Photometry [Che05]. Physical [Bia16d, Boa17, Buc17, Hei08j, JM17m, JM17l]. Physicians [MK17b]. Physicist [JM17r]. Physicists [JM17g, JM17i]. Physics [Cor04d, Cor04g, Hei08e, HG08n, JM17d, JM17j, JM17k, JM17p, JM17n, JM17o, JM17q, JM17r, JM17s]. Physik [Cor04d]. Physiological [Hie14l, Hie14r, Hie14s, Hie14t]. Physiologist [Ack13s]. Physiology [Ack13q, Bre03, Hie14p, Hie14g]. Picard [MT09o]. Pico [Rut10]. Picture [Kle12, Rut19o]. Pierpont [Loe03]. Pierre [Pal03]. Pigs [dL12g]. Pilgrimages [Cam03]. Pillars [Cor04e]. Pisum [Rhe03]. Place [Ano06p, Sow15]. Places [Phi15]. Plains [Vet15]. Planetary [Swe05]. Planets [Swe05]. Planners [Dou00]. Plant [Ack13q, Ack13s, Güt15, Har15, Phi15, Kin06]. Plants [Güt15, Phi15]. Plato [Gou10c]. Play [Bre03]. Players [dJ16]. Pocket [Hof03]. Podestà [Ben14]. Poetry [Kni17, Lan17]. Poggendorff [JM17n]. Poincaré [Rob15, Gin17o, Gin17q, Gin17f]. Point [Ray16m, Ray16n]. Points [Rob15]. Poison [Ede17]. Pol [Gin17o, Gin17k, Gin17r, Gin17s, Gin17t, Gin17u]. Policy [Cha06b, Del06]. Politics [Buc17, Phi10, Rut19c, dL12d, dS10]. Polytechnique [GG12]. Pools [Rol00]. Popular [Azz10, Siv15j]. Porta [Bel17, Bor17b, BHZ17, Eam17, Hee17, Smi17, ZH17b]. Portraits [Goe14e]. Portugal [Lei03, SCD99]. Portuguese [CL06]. Position [Piñ06]. Positions [JM17o]. Possession [SS16]. Post [Hei08g]. Posterity [Fei03a]. Poudres [Bre06]. Powder [Kin06, Smi06]. Powers [Ber14a]. Pozzo [And14]. Practical [Gli15, Hup14, Mes13m]. Practice [Edd07, Har15, Lei03, Loe03, Mat15, OP15, Ste03, dS10, Gol19-96]. Practices [Kal06, Kle10, MK17b, Wes14]. Practicing [Jäh19i, Kno07]. Praise [Goe14c]. Prato [Goe14a, Goe14h]. Praxis [Are10]. Precession [Swe10]. Precision [Ben14, Hes99, Ole97]. Predetermined [Jon05]. Predictions [Ste00r]. Premodern [Nev14]. Prescribe [Che15]. Press [Mil17, dL12g]. Pressure [Cha17c, Cha17f, Cha17n]. Principal [Nia14f]. Principia [Duc12k, Duc12l, Duc12m]. Principle [Jäh19d, Jäh19e, Jäh19i, Jäh19k]. Principles [Edd07, Pow07]. Prisms [Sha05]. Private [Sha03]. Problem [Jäh19c]. Problems [Gor03, Ros06, Smi17]. Processes [Ber14b, Gin17d]. Produced [BE15]. Production [Joh06, Kim12]. Professional [Gil12]. Professors [MK17b, Pan06]. Professorships [JM17k]. Programme [SG06]. Progress [Del06, Sch06a]. Progressive [Fin15]. Project [Fri14]. Projective [Bia16g]. Projet [Ano05k]. Prologue [CV15a]. Promoting [JM17p]. Promotion [Fin15]. Propaganda [MT09d]. Propagation [Wüt11e]. Properties [Ray16m]. Propitia [Swe12]. Proportion [Cac14]. Proposal [Fra05]. Proprietatum [Rut19p]. Protagonists [MT09l]. Protestant [Kle10]. Prototype [Gin17u]. Prout [Ano06l]. Provenance [Pan17]. Providence [Rut19a]. Prussian [Kin06]. Pseudo [Rut19p]. Pseudo-Aristotelian [Rut19p]. Psychologique [Sch10]. Psychologism [Pec06]. Ptolemy [Gou10i, Hei10b, Jon10a, Jon10b, Mit10, Rag10, Rut10, Swe05, Swe10, Tih10]. Public [BVL07, LF15, Sha03]. Publications [Buc12b, Cha07]. Pumpkins [Gli15]. Purpose [Nia14g]. Puzzle [Kle12]. Puzzling [Gou10f]. Pythagoras [Gou10c].

Quantum [Jäh19e, Jäh19i, Wüt11i, Wüt11j]. Quasi [Gin17e]. Quasi-periodic [Gin17e]. Queen [Cri17]. Quest [Duc12c]. Question [Siv15g]. Qutb [Nia14h].

Radical [Ham06]. Radio [Gin17q].

Ramsauer [Jäh19f]. Ramus

[Gou10b, Gou10g]. Rationalities [Are10]. Rayleigh [Kim02i]. Reacting [dL12h]. Reactions [Jäh19c, Rag10]. Reader [Pro15a, dC03]. Reading [Cha07, Ric06]. Reality [Bar11]. Reasons [BF05b]. Reassessment [BHZ17, ZH17a]. Reception [Ack13r, Hei10b, Ste16]. Recipe [Nev14]. Reconstructed [GM03]. Reconstructing [RS03]. Records [Ste00o, Ste00s, Ste03].

Recurrent [Gin17e]. Red [Loe03, dL12g]. redditae [Swe12]. Redefined [Cha07]. Rediscovery [Gou10b]. Reductionism [Cor04h]. Reference [GG12, Jon10a]. Reflection [Tod03]. Reflections [Mau07, Che15]. Reform [Phi10]. Reforms [JM17q]. Refraction [Bel17, Hee17]. refractione [Bor17b, Smi17]. refutata [Wes14]. Refutation [Fra05]. Regarding [Ray16c, Ray16h, Ray16j, Ray16b]. Region [Loe03]. Regional [Are10]. Regionalizing [Vet15]. Regions [DGP99]. Reich [Hie14a]. Reiche [Jäh19j]. Reichenbach [Ric06, Sch06c]. Rejection [Jon10a, Ray16n]. Relation [Art15]. Relations [MT09c, Smi10]. Relationship [Bia16f]. Relativity [Bia16h, Cor04m, RS03]. Relaxation [Gin17p, Gin17s]. Religieuse [Sch10]. Religion [Rut19b, Siv15j, Siv15k]. Religions [Siv15i]. Religious [Kla14, Kle10, Rot11]. Remarks [Nia14b]. Renaissance [Cha06a, Fin17, Eam17, Gou10b, Kie14, Piñ06, Rut10, Ste00p]. Rending [Gou10h]. Renovation [Pes06]. Replicas [Pan17]. Representation [Wüt11b]. Republic [Ver11]. Reputation [Pan09c]. Research [Ano09g, Bor06, Ham06, Has00, HRR03, Jak00, JM17m, JM17l, JM17n, Old12, Pan17, Par03, Ste03, Tod03, dL12d]. Researches [McC12c, McC12h]. Reserve [Ano06d]. Resistance [Ole97]. Resources [Arc15]. Response [Gin17r]. Responses [LF15, Pal03]. Results [Jäh19j]. Revealed [Rob15]. Reverend [McC12l]. Reviewing [SS16]. Revisiting [SS06a, SS06b]. Revolution [Cha17k, Pri07b, Wil12]. Revolutionary [HG08j, HG08d]. Revolutions [HG08k, Lev07, Rut19p]. Reworking [HRR03]. Rexford [Sow15]. Rickert [Pin10]. Right [BF05b]. Rise [FL15, LF15]. Robert [Ack13m]. Roger [Rut19b, Rut19e, Rut19o, Rut19q]. Role [GG12, Kim12, Nic06j, Rol00]. Roman [Ste00k]. Romance [BS17]. Romanticism [McO17]. Rome [Bal06a, Fri14]. Room [Kis14]. Root [Kel15]. Rose [OP15]. Royal [Hom06, Kin06]. Rudolph [Pan09a]. Rudolphine [Swe12]. Rudophinae [Swe12]. Rule [Jäh19j]. Russia [Bro06, FL15]. Russian [Ack13p, Ack13r, LF15]. Ruysch [Mel17].

S.J. [And14]. Saint [Gla14]. Saint-Peter [Gla14]. Salamanca [Cha06a]. Salamance [Bal06b]. Sammlung [Hie14q]. Samuel [Loe03]. San [Ber14a]. Sandy [Jäh19f]. Sangallo [Hup14]. Sanskrit [Kel15]. Sapientia [Rut19r]. sapientiae [Rut19q]. Sassolo [Goe14a, Goe14h]. Savile [Gou10i, Gou10b]. Scandal [Fin15]. Scandinavia [Hes99]. Scattering [Wüt11e]. Schedule [Hof03]. Scholae [Gou10c]. Scholar [Ano03c]. School [Mes13b, dS10, Gin17o, Kim02c, Kim02g, dS10]. Schwann [Par03]. Science [Ack13n, Ack13p, And00, Ano09d, Ano09b, Azz10, Bal06b, Buc12a, Buc17, BS17, CV15a, CV15b, Che05, Dij05, Ede17, Fei03b, FNB06, Fin15, Ham06, HRR03, Hub17, JPS16, Mat15, Mau12, Mül10, NG99, OP15, Pal03, Pan17, Ras06, Rup12, Sch17, Sch06a, Sch06c, Šub10, Woo06, Zel17, SG06, Ras06]. Science* [Gil12]. Sciences [Bal06a, DGP99, Gav99, Ger14a, Pep06, PK15b, SCD99, Smi10]. Scientific [Ack13n, Are10, BS06, Buc96, Cah97, Cam14, Cha17k, Cha07, Del06, Eam17, FL15, HG08b, HG08j, HG08d, Hor15, NGRR06, Ric06, dS10]. Scientist [Ano09g, Ham06, McC14s]. Scientists [McC14q]. Scorn [dL12g]. Scottish [Woo06]. Search [SS16]. Second [Cha06a, Goe14e, JM17r, Pro15a]. Secular [Ste10]. Seeing [Ano03c, Güt15]. Self [Gin17a]. Self-Oscillations [Gin17a]. Semantic [Teb15]. Seminar [JM17q]. Seminary [Gin17l]. Sensations [Hie14t, Pan09a, Pan09g]. Sensibilities [HG08m]. Sensorium [Kas18l, Kas18j, Kas18k, Kas18m, Kas18i]. Sergei [Ack13o]. Series [Gin17f, Pro15b]. Series-Dynamo [Gin17f]. Service [Bre06]. Setting [Hie14m, YKS14]. Seventeenth [Dij17, Bro11, Cha17l, Dij05, Fei03b, Kim12, Ray16o]. Seventeenth-Century [Cha17l, Kim12]. Several [Hom06]. Shīrāzī [Nia14e, Nia14h, Nia14i]. Shaping [dS10]. Shared [Hau14]. Shell [Hom06]. Should [Jon05]. Shuwei [Gol19-96]. Significance [NP05]. Simmel [Šub10]. Simple [Gin17t, Rab15]. Simplicity [Wil12]. Singing [Gin17f]. Singular [Rob15]. Siraisi [MK17b]. Sixteenth [Bar11, Bro11]. Skating [Sha05]. Smoke [Ede17]. Smokeless [Smi06]. Social [Šub10]. Society [Ric06, Rut19c, Siv15f, Gor03]. Sociology [SG06]. Soil [Ack13p, Fin15]. Solar [Jac97, Loe03, Swe10]. Solution [Gin17t, Gin17u]. Some [How06, Jac17, Jon05, Kel15, Kie14, Rab15, Siv15h]. Something [Duc12h]. Sommerfeld [Jäh19k]. Song [Kim12]. Sound [Pan09e]. Sources [Che15, Goe14i, Nia14f, Smi17]. Sovereignty [Hub17]. Soviet [Ack13r]. Space [Bia16g, Bia16j]. Spain [NG99, Bro11]. Spanish [BS06, NB06, Pes06]. Speaking [DGP99]. Spectrum [Jac97, Loe03]. Speculum [Rut19e, Rut19l]. Speech [Cam15, Kel15, Vir15a]. Speeds [Vin00]. Sphere [Ari06]. Spirit [Goe14d]. Spirits [Ano06o]. Spitzer [dL12i]. Spread [Hes99]. Square [Kel15, McC14f]. Stahl [Cha07]. Standard [Jac97]. Standards [Buc96, Ole97]. Stansel [Cam03]. Star [Bon11c, Tes11]. Stargazing [Mes13n]. Starkey [NP03]. Stars [Ver11]. State [Siv15k]. Station [Wil15]. Stations [Lav15]. Stevin [Cha17b, Cha17i, Cha17n]. Story [Mau07]. Strategies [Bla17]. Strategist [Ano09g]. Strong [SG06]. Structure [Teb15]. Structured [Pro15b]. Structures [Gla14]. Struggle [Wüt11c]. Stuart [YKS14]. Student [Mau12]. Studies [Lee15, Ray16o, Rut10, SM06]. Study [Hor15, Nia14g, Sch09, Zel17]. Studying [Ste03]. Style [Tod03]. Successive [Wüt11e]. Suits [Cro00]. sum [Jäh19j]. Summetria [HG08d]. Summus [Goe14d]. Sunrise [Ste00c]. Sunset [Ste00c]. Supersonic [Vin00]. Survey [Fri14]. Symbolic [Sch00]. Symmetry [HG08b, HG08j, HG08n, HG08d, HG08o, McO17]. Synthesis [Ack13q]. Synthetic [Ray16q]. Syntony [Hon97]. System [Ray16n]. Systematization [Wüt11d].

Tables [CG03c, Swe05, Swe12]. Tabulae [Swe12]. Taine [Mül10]. talismans [Rut19l]. Teacher [Mes13b]. Teaching [Bal06b, JM17q, Kal06]. Technical [Buc96, CG03f, Ras06, Ste06a]. Techniques [Kim12]. Technological [Joh06]. Technology [Bil00, Buc17, Dup14, Pri07a]. telescopio [Bor17b]. Temporal [Ber14a]. Ten [Kim02f]. Terms [CG03f, Hei10b]. Test [Sup00]. Tetrabiblos [Rut10]. Text [CG03i, Pro15b, dC03]. Textbooks [BS06]. Texts [Che15, CV15b, Kas18i, Kel15, Mit10, Nia14h, Pro16, Ste16, Vir15a]. Textual [BE15, CV15a, CV15b, Vir15b]. Theater [Eam17]. Theatre [And17]. Their [Bel17, Gra11, GM03]. Them [Jon05]. Theodor [Par03]. Theology [Kim12, Kle10, Rut19a, Rut19b]. Theology/Religion [Rut19b]. Theoretical [JM17j, JM17k, JM17o, JM17r, JM17s, Wüt11d]. théorie [Kno15]. Theories [AA06, Bel17, Gin17l, Hei08j, Hei08k, JM17c, Swe05]. Theorists [Hie14d]. Theory [Cam15, Cor04e, Duc12l, Fra05, Gin17m, Har15, Hei08c, Hie14t, Jäh19i, Kel15, Smi17, Swe10, Vin00, Vir15a, Wüt11d, Kno15]. Therapy [Siv15i, Siv15j, Siv15k]. Thermal [Wes14]. Thermodynamics [Ack13q, Ack13o]. These [McO17]. Things [Hau14, Kim12]. Thirteenth [Jac17, Ray16o]. Thirteenth-Century [Jac17]. Thomas [Lev07, Hir17, HH06, Rut19a, Rut19l]. Thomson [Kim02f, Kim02g]. Thornhill [McC12j, McC12l]. Thought [GBR15b, Sch17, Smi10]. Thoughts [How06, Old12]. Three [Ano09k]. Thriller [Kle12]. Thureau [AS16]. Thureau-Dangin [AS16]. Tiangong [Kim12]. Tiermärchen [Azz10]. Time [Gro15, Ros06]. Timed [Ste00s]. Times [HG08b, Ste00c, Ste00r]. Toledo [CG03c]. Tolerance [Ben14, Ole97]. Tone [Hie14t, Pan09g]. Tones [Hie14r]. Tools [Bor17b]. Total [Smi06]. Town [Kin06]. Tractatus [Ano05a]. Tradition [Mar14, Mit10, Nic06g]. Traditional [Der15, Lee15]. Traditions [Ano09k, Kim02b]. Training [Gor12, Pan09i]. Traité [Ano05d]. Trans [Hom06]. Trans-National [Hom06]. Transcendental [Bia16c]. Transform [Har15]. Transformation [Ack13d, Ack13s, Fef97, Jäh19i]. Transformations [JPS16, Pan09j, Wil15]. Transforming [Kin06]. Transition [Der15, Sch09]. Transitions [McC12k]. Translated [Goe14a]. Translating [Ack13m, Oss16]. translation [Gol19-96]. Transmission [Nev14]. Traveler [McC14t]. Treat [Duc12j]. Treatise [And14]. Treatment [HG08o]. Tree [Mes13l, Pro15b]. Tree-Structured [Pro15b]. Trees [McC14d]. Trevor [BS17, Ham17]. Trinity [Ray16c]. Triode [Gin17k]. Tropics [Wil15]. Trouble [McO17]. Truth [Rot11]. Tschermak [Gli15]. Tschirnhaus [Wes14]. Tulp [Mel17]. Tunnel [Has00]. Turbofan [SM00]. twelfth [Gol19-96]. twelfth-century [Gol19-96]. Twentieth [Lee15, Sow15, And00, GR00]. Twentieth-Century [Lee15, And00]. Two [Cor04e, Pal03, Ray16m, Ray16n, Rut10, Ste06b]. Two-Point [Ray16m, Ray16n]. Tycho [Swe10].

U.S. [Ste06a]. Uffizi [Kie14]. Uhlenbeck [Fra05]. Ultramicroscope [Cah97]. Uncovering [Duc12k, Duc12l, Duc12m]. Understanding [Goe14c, Hei10a, Lea10, Ray16p, Sch10]. Unedited [dC03]. Unified [Cor04e]. Uniqueness [Gin17u]. Unity [Pin10]. Universities [Bor06, CL06, FNB06, JM17d, NB06, Pep06, Ros06, Rut19m, Woo06]. University [JM17q, NGRR06, Pes06, SM06, Ari06, Bal06a, Bal06b, Cha06a, Del06, Gra17, Kal06, SM06]. Unpublished [Tih10]. Until [Høy16]. Uppsala [Kal06]. Uptake [Ack13n]. Urania [Swe12, Swe12]. Urban [Fri14]. Urea [GM03]. USA [Der15]. USDA [Vet15]. Use [Jon10b, Rut10, SG06, Wüt11h]. Using [Hee17, Jäh19k].

Valencia [Piñ06, SM06]. Valentin [Cam03]. Vannoccio [Ber14b]. Variation [Cac14]. Variety [Kal06]. Vauban [Lan17]. Venetian [Del06]. verba [Rut19q]. Versa [Pot06]. Verstehen [Are10, Fee10b, Fee10a, Lea10, Ueb10]. Versus [Sha03, HH06, Sch06a]. Vestiges [Kle10]. via [Hei10a]. Vice [Pot06]. Viceroyship [AA06]. Victors [JM06]. Videtur [dC03]. vie [Sch10]. Vienna [Hie14u]. View [Bia16h]. Viewpoints [Siv15h]. Views [McC14u]. Village [Kin06]. Vinci [Ber14b]. Vinogradskii [Ack13n, Ack13o, Ack13r, Ack13s]. Vision [Ray16i, Ray16o, Smi17]. Visits [BB16]. Visual [Che05, Nic06k]. Vitae [Goe14e, Goe14d]. Vito [MT09m]. Vitruvius [Cac14, Cam14]. Vittorino [Goe14a, Goe14e, Goe14d, Goe14i]. Volterra [MT09j, MT09o, MT09n, MT09m]. vs [Nia14e].

W [GBR15a]. Walked [Høy16]. Walther [Wes14]. Wand [Jäh19k]. Wanderers [How06]. War [Ano06d, Cha06b, Dou00, JM06, Lan17, Lan06, MJ06, MT09o, Sch00, Smi06, Ste06a, dL12c, MT09j, Coc06, Gin17k, dL12d]. Ward [Pin10]. Warfare [Ede17]. Wartime [Pan06]. Was [Smi05]. Watt [Mil17]. Waves [Ano05g, Ano05c, Dij05]. Weaknesses [SG06]. Wearing [dL12g]. Weber [Hei10a, JM17h, Šub10]. Weekday [Gor12]. Weighing [McC12l]. Weltanschauung [SS16]. Werke [Cah97]. West [Sow15]. Western [Ano06d]. Whale [Arc15]. Whewell [Sch06b]. Whiggery [Old12]. Whipple [Hei08j]. White [Ray16q]. Who [Dou00]. Wider [Kas18k]. Wilbur [Cro00]. Wilhelm [Ras06, Azz10]. William [Lea10, Sch06b]. Wind [Has00]. Wine [Mes13o]. Wings [Vin00]. Wisdom [Kni17]. Within [Pri14, Buc05]. Without [Wüt11i, Wüt11j, Pow07]. Wooden [Sch00]. Work [Gin17r, Jon10b, Kim12, Rob15]. Works [Ber14b, Nia14e]. Workshop [Pan09a, Pan09j, Pri07a]. World [Coc06, McC12l, Mes13k, Ste00k, dL12c, Dou00, MT09j, MT09o, Sch00, Smi06]. Wright [Cro00]. Writing [Azz10]. Wrong [BF05b, Che05, Smi05]. WW1 [MT09i].

Xu [Gol19-96].

Years [Cha17f, Cha17n, Kim02f, Par03]. Yingxing [Kim12]. Younger [Hup14].

Zeiss [Cah97, Fef97]. Zoological [Cam15]. Zürich [RS03].

References

Aparicio:2006:TCN

[AA06] Luis Carlos Arboleda Aparicio and Diana Soto Arango. The theories of Copernicus and Newton in the Viceroyship of Nueva Granada and the Audiencia De Caracas during the 18th Century. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 19, pages 289–309. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3975-1_19.

Ackert:2013:BM

[Ack13a] Lloyd Ackert. Back matter. In Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 179–191. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-94-007-5198-9/1.pdf.

Ackert:2013:C

[Ack13b] Lloyd Ackert. Conclusions. In Lloyd Ackert, editor, Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850– 1950, volume 34 of Archimedes:

New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 173–178. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-007-5198-9_11.

Ackert:2013:EM

[Ack13c] Lloyd Ackert. Ecological microbiology. In Lloyd Ackert, editor, Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 125–137. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-5198-9_8.

Ackert:2013:EMT

[Ack13d] Lloyd Ackert. The exchange of matter and the transformation of energy. In Lloyd Ackert, editor, Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 23–36. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-5198-9_2.

Ackert:2013:FNL

[Ack13e] Lloyd Ackert. Free nature in the laboratory. In Lloyd Ackert, editor, Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 59–70. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-5198-9_4.

Ackert:2013:FMa

[Ack13f] Lloyd Ackert. Front matter. In Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xviii. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bfm:978-94-007-5198-9/1.pdf.

Ackert:2013:FMb

[Ack13g] Lloyd Ackert. Front matter. In Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 1. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL

https://link.springer.com/
content/pdf/bfm:978-94-007-
5198-9/1/1.pdf.

Ackert:2013:FMc

[Ack13h] Lloyd Ackert. Front matter. In Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 37. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL

https://link.springer.com/
content/pdf/bfm:978-94-007-
5198-9/2/1.pdf.

Ackert:2013:FMd

[Ack13i] Lloyd Ackert. Front matter. In Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 71. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL

https://link.springer.com/
content/pdf/bfm:978-94-007-
5198-9/3/1.pdf.

Ackert:2013:FMe

[Ack13j] Lloyd Ackert. Front matter. In Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 105. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL

https://link.springer.com/
content/pdf/bfm:978-94-007-
5198-9/4/1.pdf.

Ackert:2013:FMf

[Ack13k] Lloyd Ackert. Front matter. In Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 139. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL

https://link.springer.com/
content/pdf/bfm:978-94-007-
5198-9/5/1.pdf.

Ackert:2013:LNI

[Ack13l] Lloyd Ackert. The laboratory is nature: Investigating the cycle of life under the microscope. In Lloyd Ackert, editor, Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 39–58. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-5198-9_3.

Ackert:2013:MBC

[Ack13m] Lloyd Ackert. The master of Brie– Compte–Robert and his “direct method:” translating the cycle of life into ecology. In Lloyd Ackert, editor, Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 107–124. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-94-007-5198-9_7.

Ackert:2013:SEE

[Ack13n] Lloyd Ackert. Science is ecological and ecology is scientific: The Uptake of Vinogradskii’s direct methods. In Lloyd Ackert, editor, Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 141–160. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-5198-9_9.

Ackert:2013:SVC

[Ack13o] Lloyd Ackert. Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2013. ISBN 94-007-5197-4 (hardcover), 94-007-5198-2 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xviii + 191 pp. LCCN QH31.W566 A25 2013e.

Ackert:2013:SSR

[Ack13p] Lloyd Ackert. Soil science and Russian ecology. In Lloyd Ackert, editor, Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 93–103. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-5198-9_6.

Ackert:2013:STB

[Ack13q] Lloyd Ackert. A synthesis of thermodynamics and bioenergetics in plant physiology: The investigation of a moody apprentice. In Lloyd Ackert, editor, Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 3–22. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-5198-9_1.

Ackert:2013:VRR

[Ack13r] Lloyd Ackert. Vinogradskii’s reception in Russian and Soviet microbiology. In Lloyd Ackert, editor, Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 161–172. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-94-007-5198-9_10.

Ackert:2013:VTP

[Ack13s] Lloyd Ackert. Vinogradskii’s transformation from plant physiologist to ecologist, 1890–1920. In Lloyd Ackert, editor, Sergei Vinogradskii and the Cycle of Life: From the Thermodynamics of Life to Ecological Microbiology, 1850–1950, volume 34 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 73–92. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5198-2. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-5198-9_5.

Anderson:2000:EAT

[And00] John D. Anderson, Jr. The evolution of aerodynamics in the Twentieth-Century: Engineering or science? In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 241– 256. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-011-4379-0_9.

Andersen:2014:MPA

[And14] Kirsti Andersen. The master of painted architecture: Andrea Pozzo, S.J. and his Treatise on Perspective. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 165–188. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-05998-3_7.

Anderson:2017:FFC

[And17] Robert G. W. Anderson. Facts or fantasies in the chemistry lecture theatre? In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 95– 115. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-58436-2_6.

Anonymous:1997:BM

[Ano97a] Anonymous. Back matter. In Jed Z. Buchwald, editor, Scientific Credibility and Technical Standards in 19th and early 20th century Germany and Britain, volume 1 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 177–182. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 1997. ISBN 94-009-1784-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bbm:978-94-009-1784-2/1.pdf.

Anonymous:1997:FM

[Ano97b] Anonymous. Front matter. In Jed Z. Buchwald, editor, Scientific Credibility and Technical Standards in 19th and early 20th century Germany and Britain, volume 1 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i– ix. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 1997. ISBN 94-009-1784-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-94-009-1784-2/1.pdf.

Anonymous:1999:FM

[Ano99] Anonymous. Front matter. In Kostas Gavroglu, editor, The Sciences in the European Periphery During the Enlightenment, volume 2 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xi. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 94-011-4770-1. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-011-4770-5/1.pdf.

Anonymous:2000:BM

[Ano00a] Anonymous. Back matter. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 361– 383. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-94-011-4379-0/1.pdf.

Anonymous:2000:FMa

[Ano00b] Anonymous. Front matter. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i– xvi. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bfm:978-94-011-4379-0/1.pdf.

Anonymous:2000:FMb

[Ano00c] Anonymous. Front matter. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 1. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (electronic).

                                        URL https://link.springer.
                                        com/content/pdf/bfm:978-94-
                                        011-4379-0/1/1.pdf.

Anonymous:2000:FMc

[Ano00d] Anonymous. Front matter. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 181. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (electronic).

                                        URL https://link.springer.
                                        com/content/pdf/bfm:978-94-
                                        011-4379-0/2/1.pdf.

Anonymous:2000:FMd

[Ano00e] Anonymous. Front matter. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 257. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (electronic).

URL https://link.springer.
com/content/pdf/bfm:978-94-
011-4379-0/3/1.pdf.

Anonymous:2000:FMe

[Ano00f] Anonymous. Front matter. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 347. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (electronic).

URL https://link.springer.
com/content/pdf/bfm:978-94-
011-4379-0/4/1.pdf.

Anonymous:2003:FMa

[Ano03a] Anonymous. Front matter. In Mordechai Feingold, editor, The New Science and Jesuit Science: Seventeenth Century Perspectives, volume 6 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–ix. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2003. ISBN 94-017-0361-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-017-0361-1/1.pdf.

Anonymous:2003:FMb

[Ano03b] Anonymous. Front matter. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xv. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-0-306-48152-9/1.pdf.

Anonymous:2003:SSE

[Ano03c] Anonymous. The scholar’s seeing eye. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 15, pages 309–325. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.

springer.com/chapter/10.
1007/0-306-48152-9_15.

Anonymous:2005:T

[Ano05a] Anonymous. 1653 — ‘Tractatus’.

In Fokko Jan Dijksterhuis, editor, Lenses and Waves: Christiaan Huygens and the Mathematical Science of Optics in the Seventeenth Century, volume 9 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 11–51. Springer Science + Business Media, Inc., Dordrecht, The Netherlands, 2005. ISBN 1-4020-2698-6. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/1-4020-2698-8_2.

Anonymous:2005:A

[Ano05b] Anonymous. 1655–1672-“de aberratione”. In Fokko Jan Dijksterhuis, editor, Lenses and Waves: Christiaan Huygens and the Mathematical Science of Optics in the Seventeenth Century, volume 9 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 53–106. Springer Science + Business Media, Inc., Dordrecht, The Netherlands, 2005. ISBN 1-4020-2698-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-2698-8_3.

Anonymous:2005:WL

[Ano05c] Anonymous. 1677–1679 — waves of light. In Fokko Jan Dijksterhuis, editor, Lenses and Waves: Christiaan Huygens and the Mathematical Science of Optics in the Seventeenth Century, volume 9 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 159–211. Springer Science + Business Media, Inc., Dordrecht, The Netherlands, 2005. ISBN 1-4020-2698-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-2698-8_5.

Anonymous:2005:TL

[Ano05d] Anonymous. 1690 — traité de la lumière. In Fokko Jan Dijksterhuis, editor, Lenses and Waves: Christiaan Huygens and the Mathematical Science of Optics in the Seventeenth Century, volume 9 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 213–253. Springer Science + Business Media, Inc., Dordrecht, The Netherlands, 2005. ISBN 1-4020-2698-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-2698-8_6.

Anonymous:2005:BMa

[Ano05e] Anonymous. Back matter. In Fokko Jan Dijksterhuis, editor, Lenses and Waves: Christiaan Huygens and the Mathematical Science of Optics in the Seventeenth Century, volume 9 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 265–288. Springer Science + Business Media, Inc., Dordrecht, The Netherlands, 2005. ISBN 1-4020-2698-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bbm:978-1-4020-2698-0/1.pdf.

Anonymous:2005:BMb

[Ano05f] Anonymous. Back matter. In Jed Z. Buchwald and Allan Franklin, editors, Wrong for the Right Reasons, volume 11 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 229. Springer, Dordrecht, The Netherlands, 2005. ISBN 1-4020-3048-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bbm:978-1-4020-3048-2/1.pdf.

Anonymous:2005:CLW

[Ano05g] Anonymous. Conclusion: Lenses & waves. In Fokko Jan Dijksterhuis, editor, Lenses and Waves: Christiaan Huygens and the Mathematical Science of Optics in the Seventeenth Century, volume 9 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 255–266. Springer Science + Business Media, Inc., Dordrecht, The Netherlands, 2005. ISBN 1-4020-2698-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-2698-8_7.

Anonymous:2005:FMa

[Ano05h] Anonymous. Front matter. In Fokko Jan Dijksterhuis, editor, Lenses and Waves: Christiaan Huygens and the Mathematical Science of Optics in the Seventeenth Century, volume 9 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–viii. Springer Science + Business Media, Inc., Dordrecht, The Netherlands, 2005. ISBN 1-4020-2698-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-1-4020-2698-0/1.pdf.

Anonymous:2005:FMb

[Ano05i] Anonymous. Front matter. In Jed Z. Buchwald and Allan Franklin, editors, Wrong for the Right Reasons, volume 11 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–viii. Springer, Dordrecht, The Netherlands, 2005. ISBN 1-4020-3048-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-1-4020-3048-2/1.pdf.

Anonymous:2005:IPC

[Ano05j] Anonymous. Introduction — ‘the perfect Cartesian’. In Fokko Jan Dijksterhuis, editor, Lenses and Waves: Christiaan Huygens and the Mathematical Science of Optics in the Seventeenth Century, volume 9 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–10. Springer Science + Business Media, Inc., Dordrecht, The Netherlands, 2005. ISBN 1-4020-2698-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-2698-8_1.

Anonymous:2005:P

[Ano05k] Anonymous. The ‘projet’ of 1672.

In Fokko Jan Dijksterhuis, editor, Lenses and Waves: Christiaan Huygens and the Mathematical Science of Optics in the Seventeenth Century, volume 9 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 107–158. Springer Science + Business Media, Inc., Dordrecht, The Netherlands, 2005. ISBN 1-4020-2698-6. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/1-4020-2698-8_4.

Anonymous:2006:BMa

[Ano06a] Anonymous. Back matter. In Ralph R. Hamerla, editor, An American Scientist on the Research Frontier: Edward Morley, Community, and Radical Ideas in Nineteenth-Century Science, volume 13 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 231–261. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4089-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.

                                        springer.com/content/pdf/
                                        bbm:978-1-4020-4089-4/1.pdf.

Anonymous:2006:BMb

[Ano06b] Anonymous. Back matter. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 231–233. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bbm:978-1-4020-4251-5/1.pdf.

Anonymous:2006:BMc

[Ano06c] Anonymous. Back matter. In Roy Macleod and Jeffrey Allan Johnson, editors, Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 247–279. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5490-4. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bbm:978-1-4020-5490-7/1.pdf.

Anonymous:2006:EME

[Ano06d] Anonymous. Edward Morley: Education, civil war, and the Western Reserve. In Ralph R. Hamerla, editor, An American Scientist on the Research Frontier: Edward Morley, Community, and Radical Ideas in Nineteenth-Century Science, volume 13 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 47–91. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4089-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-4089-X_3.

Anonymous:2006:FMa

[Ano06e] Anonymous. Front matter. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–viii. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-1-4020-3975-1/1.pdf.

Anonymous:2006:FMb

[Ano06f] Anonymous. Front matter. In Ralph R. Hamerla, editor, An American Scientist on the Research Frontier: Edward Morley, Community, and Radical Ideas in Nineteenth-Century Science, volume 13 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xiii. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4089-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-1-4020-4089-4/1.pdf.

Anonymous:2006:FMc

[Ano06g] Anonymous. Front matter. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages v–vi. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-1-4020-4251-5/1.pdf.

Anonymous:2006:FMd

[Ano06h] Anonymous. Front matter. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 1. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-1-4020-4251-5/1/1.
                                        pdf.

Anonymous:2006:FMe

[Ano06i] Anonymous. Front matter. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 55. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-1-4020-4251-5/2/1.
pdf.

Anonymous:2006:FMf

[Ano06j] Anonymous. Front matter. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 117. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-1-4020-4251-5/3/1.
pdf.

Anonymous:2006:FMg

[Ano06k] Anonymous. Front matter. In Roy Macleod and Jeffrey Allan Johnson, editors, Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages ii–xix. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5490-4.

ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-1-4020-5490-7/1.pdf.

Anonymous:2006:IHP

[Ano06l] Anonymous. Intellectual Heritage, Prout’s hypothesis. In Ralph R. Hamerla, editor, An American Scientist on the Research Frontier: Edward Morley, Community, and Radical Ideas in Nineteenth-Century Science, volume 13 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 153–179. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4089-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-4089-X_6.

Anonymous:2006:Ia

[Ano06m] Anonymous. Introduction. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–7. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-3975-1_1.

Anonymous:2006:Ib

[Ano06n] Anonymous. Introduction. In Ralph R. Hamerla, editor, An American Scientist on the Research Frontier: Edward Morley, Community, and Radical Ideas in Nineteenth-Century Science, volume 13 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–8. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4089-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/1-4020-4089-X_1.

Anonymous:2006:KSE

[Ano06o] Anonymous. Kindred spirits: The ether drift. In Ralph R. Hamerla, editor, An American Scientist on the Research Frontier: Edward Morley, Community, and Radical Ideas in Nineteenth-Century Science, volume 13 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 123–152. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4089-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-4089-X_5.

Anonymous:2006:MP

[Ano06p] Anonymous. Making a place. In Ralph R. Hamerla, editor, An American Scientist on the Research Frontier: Edward Morley, Community, and Radical Ideas in Nineteenth-Century Science, volume 13 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 93–121.

Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4089-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-4089-X_4.

Anonymous:2006:M

[Ano06q] Anonymous. The Morleys. In Ralph R. Hamerla, editor, An American Scientist on the Research Frontier: Edward Morley, Community, and Radical Ideas in Nineteenth-Century Science, volume 13 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 9–45. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4089-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/1-4020-4089-X_2.

Anonymous:2006:O

[Ano06r] Anonymous. Oxygen. In Ralph R. Hamerla, editor, An American Scientist on the Research Frontier: Edward Morley, Community, and Radical Ideas in Nineteenth-Century Science, volume 13 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 181–229. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4089-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/1-4020-4089-X_7.

Anonymous:2007:BM

[Ano07a] Anonymous. Back matter. In Lawrence M. Principe, editor, New Narratives in Eighteenth-Century Chemistry: Contributions from the First Francis Bacon Workshop, 21–23 April 2005, California Institute of Technology, Pasadena, California, volume 18 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 195–200. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2007. ISBN 1-4020-6278-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-1-4020-6278-0/1.pdf.

Anonymous:2007:FM

[Ano07b] Anonymous. Front matter. In Lawrence M. Principe, editor, New Narratives in Eighteenth-Century Chemistry: Contributions from the First Francis Bacon Workshop, 21–23 April 2005, California Institute of Technology, Pasadena, California, volume 18 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i– xi. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2007. ISBN 1-4020-6278-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-1-4020-6278-0/1.pdf.

Anonymous:2009:BM

[Ano09a] Anonymous. Back matter. In Gregor Schiemann, editor, Hermann von Helmholtz’s Mechanism: The Loss of Certainty: a Study on the Transition from Classical to Modern Philosophy of Nature, volume 17 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 249–283. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 1-4020-5630-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bbm:978-1-4020-5630-7/1.pdf.

Anonymous:2009:CCS

[Ano09b] Anonymous. The classical conception of science. In Gregor Schiemann, editor, Hermann von Helmholtz’s Mechanism: The Loss of Certainty: a Study on the Transition from Classical to Modern Philosophy of Nature, volume 17 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 23–31. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 1-4020-5630-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-1-4020-5630-7_3.

Anonymous:2009:CM

[Ano09c] Anonymous. The conception of mechanism. In Gregor Schiemann, editor, Hermann von Helmholtz’s Mechanism: The Loss of Certainty: a Study on the Transition from Classical to Modern Philosophy of Nature, volume 17 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 15–22. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 1-4020-5630-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-1-4020-5630-7_2.

Anonymous:2009:CCC

[Ano09d] Anonymous. Conditions and causes for the Change in Helmholtz’s conception of science and nature. In Gregor Schiemann, editor, Hermann von Helmholtz’s Mechanism: The Loss of Certainty: a Study on the Transition from Classical to Modern Philosophy of Nature, volume 17 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 229–248. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 1-4020-5630-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-1-4020-5630-7_9.

Anonymous:2009:CMP

[Ano09e] Anonymous. Contours of modern philosophy of nature. In Gregor Schiemann, editor, Hermann von Helmholtz’s Mechanism: The Loss of Certainty: a Study on the Transition from Classical to Modern Philosophy of Nature, volume 17 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 43–54. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 1-4020-5630-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-1-4020-5630-7_5.

Anonymous:2009:FM

[Ano09f] Anonymous. Front matter. In Gregor Schiemann, editor, Hermann von Helmholtz’s Mechanism: The Loss of Certainty: a Study on the Transition from Classical to Modern Philosophy of Nature, volume 17 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–x. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 1-4020-5630-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-1-4020-5630-7/1.pdf.

Anonymous:2009:HBS

[Ano09g] Anonymous. Helmholtz, a Bildungsbürger, scientist, and research strategist. In Gregor Schiemann, editor, Hermann von Helmholtz’s Mechanism: The Loss of Certainty: a Study on the Transition from Classical to Modern Philosophy of Nature, volume 17 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 57–74. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 1-4020-5630-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-1-4020-5630-7_6.

Anonymous:2009:HCM

[Ano09h] Anonymous. Helmholtz’s classical mechanism. In Gregor Schiemann, editor, Hermann von Helmholtz’s Mechanism: The Loss of Certainty: a Study on the Transition from Classical to Modern Philosophy of Nature, volume 17 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 75–157. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 1-4020-5630-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-1-4020-5630-7_7.

Anonymous:2009:HHM

[Ano09i] Anonymous. The hypothetization of Helmholtz’s mechanism. In Gregor Schiemann, editor, Hermann von Helmholtz’s Mechanism: The Loss of Certainty: a Study on the Transition from Classical to Modern Philosophy of Nature, volume 17 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 159– 227. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 1-4020-5630-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-5630-7_8.

Anonymous:2009:I

[Ano09j] Anonymous. Introduction. In Gregor Schiemann, editor, Hermann von Helmholtz’s Mechanism: The Loss of Certainty: a Study on the Transition from Classical to Modern Philosophy of Nature, volume 17 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–11. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 1-4020-5630-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-1-4020-5630-7_1.

Anonymous:2009:TTM

[Ano09k] Anonymous. Three traditions in mechanism. In Gregor Schiemann, editor, Hermann von Helmholtz’s Mechanism: The Loss of Certainty: a Study on the Transition from Classical to Modern Philosophy of Nature, volume 17 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 33–41. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 1-4020-5630-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-1-4020-5630-7_4.

Anonymous:2010:BMa

[Ano10a] Anonymous. Back matter. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 311–320. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0.

ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bbm:978-90-481-3540-0/1.pdf.

Anonymous:2010:BMb

[Ano10b] Anonymous. Back matter. In Alexander Jones, editor, Ptolemy in Perspective: Use and Criticism of his Work from Antiquity to the Nineteenth Century, volume 23 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 217–229. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-2788-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-90-481-2788-7/1.pdf.

Anonymous:2010:FMa

[Ano10c] Anonymous. Front matter. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–vi. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-90-481-3540-0/1.pdf.

Anonymous:2010:FMb

[Ano10d] Anonymous. Front matter. In Alexander Jones, editor, Ptolemy in Perspective: Use and Criticism of his Work from Antiquity to the Nineteenth Century, volume 23 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages ii–xv. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-2788-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-90-481-2788-7/1.pdf.

Anonymous:2011:BM

[Ano11a] Anonymous. Back matter. In Patrick J. Boner, editor, Change and Continuity in Early Modern Cosmology, volume 27 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 177–181. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 94-007-0037-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-94-007-0037-6/1.pdf.

Anonymous:2011:FM

[Ano11b] Anonymous. Front matter. In Patrick J. Boner, editor, Change and Continuity in Early Modern Cosmology, volume 27 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xii. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 94-007-0037-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-007-0037-6/1.pdf.

Anonymous:2012:FMa

[Ano12a] Anonymous.

Front matter. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xix. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-007-2627-7/1.pdf.

Anonymous:2012:FMb

[Ano12b] Anonymous. Front matter. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 1. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-007-2627-7/1/1.
pdf.

Anonymous:2012:FMc

[Ano12c] Anonymous.

Front matter. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 47. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-007-2627-7/2/1.
pdf.

Anonymous:2012:FMd

[Ano12d] Anonymous. Front matter. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 79. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-007-2627-7/3/1.
pdf.

Anonymous:2012:FMe

[Ano12e] Anonymous. Front matter. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 135. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.

springer.com/content/pdf/
bfm:978-94-007-2627-7/4/1.
pdf.

Anonymous:2012:FMf

[Ano12f] Anonymous.

Front matter. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 165. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-007-2627-7/5/1.
pdf.

Anonymous:2012:FMg

[Ano12g] Anonymous.

Front matter. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 215. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-007-2627-7/6/1.
pdf.

Anonymous:2014:BM

[Ano14a] Anonymous. Back matter. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 313–318. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bbm:978-3-319-05998-3/1.pdf.

Anonymous:2014:FMa

[Ano14b] Anonymous. Front matter. In Sven Dupré, editor, Laboratories of Art: Alchemy and Art Technology from Antiquity to the 18th Century, volume 37 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xxiii. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05065-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-3-319-05065-2/1.pdf.

Anonymous:2014:FMb

[Ano14c] Anonymous. Front matter. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xiv. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-3-319-05998-3/1.pdf.

Anonymous:2014:FMc

[Ano14d] Anonymous. Front matter. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 43–46. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/content/
pdf/bfm:978-3-319-05998-3/1/
1.pdf.

Anonymous:2014:FMd

[Ano14e] Anonymous. Front matter. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 59–60. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/content/
                                        pdf/bfm:978-3-319-05998-3/2/
                                        1.pdf.

Anonymous:2014:FMe

[Ano14f] Anonymous. Front matter. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 135–136. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/content/
                                        pdf/bfm:978-3-319-05998-3/3/
                                        1.pdf.

Anonymous:2014:FMf

[Ano14g] Anonymous. Front matter. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 189–191. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/content/
pdf/bfm:978-3-319-05998-3/4/
1.pdf.

Anonymous:2014:FMg

[Ano14h] Anonymous. Front matter. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 243–245. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/content/
pdf/bfm:978-3-319-05998-3/5/
1.pdf.

Anonymous:2015:BMa

[Ano15a] Anonymous. Back matter. In Norma B. Goethe, Philip Beeley, and David Rabouin, editors, G. W. Leibniz, Interrelations between Mathematics and Philosophy, volume 41 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 207–210. Springer Netherlands, Dordrecht, The Netherlands, 2015. ISBN 94-017-9664-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bbm:978-94-017-9664-4/1.pdf.

Anonymous:2015:BMb

[Ano15b] Anonymous. Back matter. In Karine Chemla and Jacques Virbel, editors, Texts, Textual Acts and the History of Science, volume 42 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 421–430. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-16444-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bbm:978-3-319-16444-1/1.pdf.

Anonymous:2015:FMa

[Ano15c] Anonymous. Front matter. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–vii. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bfm:978-3-319-12185-7/1.pdf.

Anonymous:2015:FMb

[Ano15d] Anonymous. Front matter. In Norma B. Goethe, Philip Beeley, and David Rabouin, editors, G. W. Leibniz, Interrelations between Mathematics and Philosophy, volume 41 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–x. Springer Netherlands, Dordrecht, The Netherlands, 2015. ISBN 94-017-9664-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-94-017-9664-4/1.pdf.

Anonymous:2015:FMc

[Ano15e] Anonymous. Front matter. In Norma B. Goethe, Philip Beeley, and David Rabouin, editors, G. W. Leibniz, Interrelations between Mathematics and Philosophy, volume 41 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 1. Springer Netherlands, Dordrecht, The Netherlands, 2015. ISBN 94-017-9664-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-94-017-9664-4/1/1.
pdf.

Anonymous:2015:FMd

[Ano15f] Anonymous. Front matter. In Norma B. Goethe, Philip Beeley, and David Rabouin, editors, G. W. Leibniz, Interrelations between Mathematics and Philosophy, volume 41 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 73. Springer Netherlands, Dordrecht, The Netherlands, 2015. ISBN 94-017-9664-5.

ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-94-017-9664-4/2/1.
                                        pdf.

Anonymous:2015:FMe

[Ano15g] Anonymous. Front matter. In Norma B. Goethe, Philip Beeley, and David Rabouin, editors, G. W. Leibniz, Interrelations between Mathematics and Philosophy, volume 41 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 135. Springer Netherlands, Dordrecht, The Netherlands, 2015. ISBN 94-017-9664-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-94-017-9664-4/3/1.
                                        pdf.

Anonymous:2015:FMf

[Ano15h] Anonymous. Front matter. In Karine Chemla and Jacques Virbel, editors, Texts, Textual Acts and the History of Science, volume 42 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–ix. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-16444-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-3-319-16444-1/1.pdf.

Anonymous:2015:FMg

[Ano15i] Anonymous. Front matter. In Karine Chemla and Jacques Virbel, editors, Texts, Textual Acts and the History of Science, volume 42 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 47. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-16444-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/content/
pdf/bfm:978-3-319-16444-1/1/
1.pdf.

Anonymous:2015:FMh

[Ano15j] Anonymous. Front matter. In Karine Chemla and Jacques Virbel, editors, Texts, Textual Acts and the History of Science, volume 42 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 219. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-16444-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/content/
pdf/bfm:978-3-319-16444-1/2/
1.pdf.

Anonymous:2016:FM

[Ano16] Anonymous.

Front matter. In Alexander Jones, Christine Proust, and John M. Steele, editors, A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science, volume 45 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i– xi. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-25865-6. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-3-319-25865-2/1.pdf.

Anonymous:2017:BMa

[Ano17a] Anonymous. Back matter. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 299– 307. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-3-319-56514-9/1.pdf.

Anonymous:2017:BMb

[Ano17b] Anonymous. Back matter. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 301– 310. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-3-319-58436-2/1.pdf.

Anonymous:2017:FMa

[Ano17c] Anonymous. Front matter. In Arianna Borrelli, Giora Hon, and Yaakov Zik, editors, The Optics of Giambattista Della Porta (ca. 1535–1615): A Reassessment, volume 44 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–v. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-50215-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-3-319-50215-1/1.pdf.

Anonymous:2017:FMb

[Ano17d] Anonymous. Front matter. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xvi. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-3-319-56514-9/1.pdf.

Anonymous:2017:FMc

[Ano17e] Anonymous. Front matter. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 29. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-3-319-56514-9/1/1.
                                        pdf.

Anonymous:2017:FMd

[Ano17f] Anonymous. Front matter. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 77. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-3-319-56514-9/2/1.
                                        pdf.

Anonymous:2017:FMe

[Ano17g] Anonymous. Front matter. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 153. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-3-319-56514-9/3/1.
pdf.

Anonymous:2017:FMf

[Ano17h] Anonymous. Front matter. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i– ix. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-3-319-58436-2/1.pdf.

Arabatzis:2006:ICD

[Ara06] Theodore Arabatzis. On the inextricability of the context of discovery and the context of justification. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 13, pages 215–230. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-4251-5_13.

Arch:2015:WOP

[Arc15] Jakobina Arch. Whale oil pesticide: Natural history, animal resources, and agriculture in Early Modern Japan. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 93–111. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-12185-7_6.

Arens:2010:EVE

[Are10] Katherine Arens. Erklären, Verstehen, and embodied rationalities: Scientific praxis as regional ontology. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 141–159.

Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-3540-0_8.

Ariew:2006:SJD

[Ari06] Roger Ariew. The sphere of Jacques du Chevreul: Astronomy at the University of Paris in the 1620s. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 99–109. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3975-1_7.

Arthur:2015:LAI

[Art15] Richard T. W. Arthur. Leibniz’s actual infinite in relation to his analysis of matter. In Norma B. Goethe, Philip Beeley, and David Rabouin, editors, G. W. Leibniz, Interrelations between Mathematics and Philosophy, volume 41 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 137–156. Springer Netherlands, Dordrecht, The Netherlands, 2015. ISBN 94-017-9664-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-94-017-9664-4_7.

Andre-Salvini:2016:FTD

[AS16] Béatrice André-Salvini. François Thureau-Dangin and cuneiform mathematics. In Alexander Jones, Christine Proust, and John M. Steele, editors, A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science, volume 45 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 197–206. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-25865-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-25865-2_6.

Azzouni:2010:HWD

[Azz10] Safia Azzouni. How Wilhelm Dilthey influenced popular science writing: Kurd Laßwitz’s “Homchen. Ein Tiermärchen aus der oberen Kreide”. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 61–79. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.

springer.com/chapter/10.
1007/978-90-481-3540-0_4.

Baldini:2006:SUR

[Bal06a] Ugo Baldini. The sciences at the University of Rome in the 18th Century. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 14, pages 201–230. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3975-1_14.

Ballester:2006:MSM

[Bal06b] Luis García Ballester. Medical science and medical teaching at the University of Salamance in the 15th century. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 37–64. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3975-1_4.

Barker:2011:RPO

[Bar11] Peter Barker. The reality of Peurbach’s orbs: Cosmological continuity in Fifteenth and Sixteenth Century astronomy. In Patrick J. Boner, editor, Change and Continuity in Early Modern Cosmology, volume 27 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 7–32. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 94-007-0037-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-0037-6_2.

Brack-Bernsen:2016:ONV

[BB16] Lis Brack-Bernsen. Otto Neugebauer’s visits to Copenhagen and his connection to Denmark. In Alexander Jones, Christine Proust, and John M. Steele, editors, A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science, volume 45 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 107–126. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-25865-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-25865-2_3.

Bretelle-Establet:2015:ITG

[BE15] Florence Bretelle-Establet. The issue of textual genres in the medical literature produced in Late Imperial China. In Karine Chemla and Jacques Virbel, editors, Texts, Textual Acts and the History of Science, volume 42 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 87–142. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-16444-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-16444-1_3.

Beeley:2015:LPM

[Bee15] Philip Beeley. Leibniz, philosopher mathematician and mathematical philosopher. In Norma B. Goethe, Philip Beeley, and David Rabouin, editors, G. W. Leibniz, Interrelations between Mathematics and Philosophy, volume 41 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 23–48. Springer Netherlands, Dordrecht, The Netherlands, 2015. ISBN 94-017-9664-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-017-9664-4_2.

Belle:2017:FMG

[Bel17] Riccardo Bellé. Francesco Maurolico, Giambattista Della Porta and their theories on refraction. In Arianna Borrelli, Giora Hon, and Yaakov Zik, editors, The Optics of Giambattista Della Porta (ca. 1535–1615): A Reassessment, volume 44 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 169– 200. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-50215-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-50215-1_8.

Benelli:2014:PPB

[Ben14] Francesco Benelli. The Palazzo del Podestà in Bologna: Precision and tolerance in a building all’Antica. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 61–77. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-05998-3_3.

Beretta:2014:MTP

[Ber14a] Marco Beretta. Material and temporal powers at the Casino di San Marco (1574–1621). In Sven Dupré, editor, Laboratories of Art: Alchemy and Art Technology from Antiquity to the 18th Century, volume 37 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 129–156. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05065-6. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-05065-2_6.

Bernardoni:2014:APE

[Ber14b] Andrea Bernardoni. Artisanal processes and epistemological debate in the works of Leonardo Da Vinci and Vannoccio Biringuccio. In Sven Dupré, editor, Laboratories of Art: Alchemy and Art Technology from Antiquity to the 18th Century, volume 37 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 53–78. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05065-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-05065-2_3.

Buchwald:2005:IBD

[BF05a] Jed Z. Buchwald and Allan Franklin. Introduction: Beyond disunity and historicism. In Jed Z. Buchwald and Allan Franklin, editors, Wrong for the Right Reasons, volume 11 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–16. Springer, Dordrecht, The Netherlands, 2005. ISBN 1-4020-3048-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-3048-7_1.

Buchwald:2005:WRR

[BF05b] Jed Z. Buchwald and Allan Franklin, editors. Wrong for the Right Reasons, volume 11 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer, Dordrecht, The Netherlands, 2005. ISBN 1-4020-3047-9 (hardcover), 1-4020-3048-7 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). viii + 228 pp. LCCN Q175 .W7916 2005.

Borrelli:2017:OGD

[BHZ17] Arianna Borrelli, Giora Hon, and Yaakov Zik, editors. The Optics of Giambattista Della Porta (ca. 1535–1615): A Reassessment, volume 44 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-50214-X (hardcover), 3-319-50215-8 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). v + 205 pp. LCCN PQ4630.P6; QC352 .O68 2017.

Biagioli:2016:AHD

[Bia16a] Francesca Biagioli. Axioms, hypotheses, and definitions. In Space, Number, and Geometry from Helmholtz to Cassirer, volume 46 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 51–80. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-31779-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-31779-3_3.

Biagioli:2016:BM

[Bia16b] Francesca Biagioli. Back matter. In Space, Number, and Geometry from Helmholtz to Cassirer, volume 46 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 229–239. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-31779-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bbm:978-3-319-31779-3/1.pdf.

Biagioli:2016:DKT

[Bia16c] Francesca Biagioli. The discussion of Kant’s transcendental aesthetic. In Space, Number, and Geometry from Helmholtz to Cassirer, volume 46 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 23–50. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-31779-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-31779-3_2.

Biagioli:2016:ENE

[Bia16d] Francesca Biagioli. Euclidean and non-Euclidean geometries in the interpretation of physical measurements. In Space, Number, and Geometry from Helmholtz to Cassirer, volume 46 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 151–188. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-31779-2. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-31779-3_6.

Biagioli:2016:FM

[Bia16e] Francesca Biagioli. Front matter. In Space, Number, and Geometry from Helmholtz to Cassirer, volume 46 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xx. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-31779-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-3-319-31779-3/1.pdf.

Biagioli:2016:HRK

[Bia16f] Francesca Biagioli. Helmholtz’s relationship to Kant. In Space, Number, and Geometry from Helmholtz to Cassirer, volume 46 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–21. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-31779-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-31779-3_1.

Biagioli:2016:MPG

[Bia16g] Francesca Biagioli. Metrical projective geometry and the concept of space. In Space, Number, and Geometry from Helmholtz to Cassirer, volume 46 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 117–149. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-31779-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-31779-3_5.

Biagioli:2016:NEG

[Bia16h] Francesca Biagioli. Non-Euclidean geometry and Einstein’s general relativity: Cassirer’s view in 1921. In Space, Number, and Geometry from Helmholtz to Cassirer, volume 46 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 189–228. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-31779-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-31779-3_7.

Biagioli:2016:NM

[Bia16i] Francesca Biagioli. Number and magnitude. In Space, Number, and Geometry from Helmholtz to Cassirer, volume 46 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 81–116. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-31779-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-31779-3_4.

Biagioli:2016:SNG

[Bia16j] Francesca Biagioli. Space, Number, and Geometry from Helmholtz to Cassirer, volume 46 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2016. ISBN 3-319-31777-6 (hardcover), 3-319-31779-2 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xx + 239 pp. LCCN QA685 .B49 2016.

Bilstein:2000:AAT

[Bil00] Roger E. Bilstein. American aviation technology: An international heritage. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 207– 222. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-011-4379-0_7.

Blair:2017:DSC

[Bla17] Ann Blair. The dedication strategies of Conrad Gessner. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 197–237. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-56514-9_10.

Bloor:2000:DP

[Blo00] David Bloor. Discussion paper. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 14, pages 349–360. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-011-4379-0_14.

Boantza:2017:EIM

[Boa17] Victor D. Boantza. Elements, instruments, and menstruums: Boerhaave’s imponderable fire between chemical masterpiece and physical axiom. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 9–46. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-58436-2_2.

Boner:2011:CCE

[Bon11a] Patrick J. Boner, editor. Change and Continuity in Early Modern Cosmology, volume 27 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2011. ISBN 94-007-0036-9 (hardcover), 94-007-0037-7 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xii + 181 pp. LCCN QB981 .C5325 2011.

Boner:2011:I

[Bon11b] Patrick J. Boner. Introduction. In Patrick J. Boner, editor, Change and Continuity in Early Modern Cosmology, volume 27 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–6. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 94-007-0037-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-0037-6_1.

Boner:2011:KCC

[Bon11c] Patrick J. Boner. Kepler’s Copernican campaign and the new star of 1604. In Patrick J. Boner, editor, Change and Continuity in Early Modern Cosmology, volume 27 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 93–114. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 94-007-0037-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-94-007-0037-6_6.

Borgato:2006:MRI

[Bor06] Maria Teresa Borgato. Mathematical research in Italian universities in the modern era. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 127–139. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-3975-1_9.

Borrelli:2017:I

[Bor17a] Arianna Borrelli. Introduction. In Arianna Borrelli, Giora Hon, and Yaakov Zik, editors, The Optics of Giambattista Della Porta (ca. 1535–1615): A Reassessment, volume 44 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–10. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-50215-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-50215-1_1.

Borrelli:2017:ODP

[Bor17b] Arianna Borrelli. Optical diagrams as “paper tools”: Della Porta’s analysis of biconvex lenses from de refractione to de telescopio. In Arianna Borrelli, Giora Hon, and Yaakov Zik, editors, The Optics of Giambattista Della Porta (ca. 1535–1615): A Reassessment, volume 44 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 57–96. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-50215-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-50215-1_4.

Bos:2010:IIN

[Bos10] Jacques Bos. Individuality and interpretation in Nineteenth-Century German historicism. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 207–220. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-3540-0_11.

Bresadola:2003:PNL

[Bre03] Marco Bresadola. At play with nature: Luigi Galvani’s experimental approach to muscular physiology. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 67–92. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/0-306-48152-9_4.

Bret:2006:MCE

[Bre06] Patrice Bret. Managing chemical expertise: The laboratories of the French Artillery and the Service des Poudres. In Roy Macleod and Jeffrey Allan Johnson, editors, Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 203–219. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5490-4. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-5490-4_12.

Brooks:2006:MMC

[Bro06] Nathan M. Brooks. Munitions, the military, and chemistry in Russia. In Roy Macleod and Jeffrey Allan Johnson, editors, Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 75–101. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5490-4. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-5490-4_6.

Brotons:2011:CCC

[Bro11] Víctor Navarro Brotóns. Continuity and change in cosmological ideas in Spain between the Sixteenth and Seventeenth Centuries: The impact of celestial novelties. In Patrick J. Boner, editor, Change and Continuity in Early Modern Cosmology, volume 27 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 33–50. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 94-007-0037-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-0037-6_3.

Belmar:2006:SCT

[BS06] Antonio García Belmar and José Ramón Bertomeu Sánchez. Spanish chemistry textbooks during late 18th Century: Building up a new genre of scientific literature. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 16, pages 241–257. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-3975-1_16.

Buchwald:2017:RSE

[BS17] Jed Buchwald and Larry Stewart, editors. The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58435-9 (hardcover), 3-319-58436-7 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). ix + 310 pp. LCCN Q125 .R733 2017.

Buchwald:1996:SCT

[Buc96] Jed Z. Buchwald, editor. Scientific Credibility and Technical Standards in 19th and early 20th Century Germany and Britain, volume 1 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 1996. ISBN 0-7923-4762-5 (hardcover), 94-009-1784-8 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). ix + 182 pp. LCCN ????

Buchwald:2005:EWM

[Buc05] Jed Z. Buchwald. An error within a mistake? In Jed Z. Buchwald and Allan Franklin, editors, Wrong for the Right Reasons, volume 11 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 185–208. Springer, Dordrecht, The Netherlands, 2005. ISBN 1-4020-3048-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3048-7_9.

Buchwald:2012:MSH

[Buc12a] Jed Z. Buchwald, editor. A Master of Science History: essays in honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2012. ISBN 94-007-2626-0 (hardcover), 94-007-2627-9 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xix + 439 pp. LCCN Q125 .M37 2012.

Buchwald:2012:PCC

[Buc12b] Jed Z. Buchwald. Publications by Charles Coulston Gillispie. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 3–13. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2627-7_1.

Buchwald:2017:PMI

[Buc17] Jed Buchwald. Politics, morality, innovation, and misrepresentation in physical science and technology. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 201–217. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-58436-2_12.

Bensaude-Vincent:2007:PLC

[BVL07] Bernadette Bensaude-Vincent and Christine Lehman. Public lectures of chemistry in Mid-Eighteenth-Century France. In Lawrence M. Principe, editor, New Narratives in Eighteenth-Century Chemistry: Contributions from the First Francis Bacon Workshop, 21–23 April 2005, California Institute of Technology, Pasadena, California, volume 18 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 77–96. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2007. ISBN 1-4020-6278-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-1-4020-6278-0_5.

Cache:2014:PCV

[Cac14] Bernard Cache. Proportion and continuous variation in Vitruvius’s de architectura. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 47–58. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-05998-3_2.

Cahan:1997:ZWU

[Cah97] David Cahan. The Zeiss Werke and the ultramicroscope: The creation of a scientific instrument in context. In Jed Z. Buchwald, editor, Scientific Credibility and Technical Standards in 19th and early 20th century Germany and Britain, volume 1 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 67–115. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 1997. ISBN 94-009-1784-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-009-1784-2_3.

Camenietzki:2003:CPV

[Cam03] Carlos Ziller Camenietzki. The celestial pilgrimages of Valentin Stansel (1621–1705), Jesuit astronomer and missionary in Brazil. In Mordechai Feingold, editor, The New Science and Jesuit Science: Seventeenth Century Perspectives, volume 6 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 249–270. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2003. ISBN 94-017-0361-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-017-0361-1_6.

Camerota:2014:SCB

[Cam14] Filippo Camerota. A scientific concept of beauty in architecture: Vitruvius meets Descartes, Galileo, and Newton. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 215–241. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-05998-3_10.

Cambefort:2015:ZNS

[Cam15] Yves Cambefort. Zoological nomenclature and speech act theory. In Karine Chemla and Jacques Virbel, editors, Texts, Textual Acts and the History of Science, volume 42 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 143–181. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-16444-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-16444-1_4.

Cerruti:1999:DB

[Cer99] Luigi Cerruti. Dante’s bones.

In Kostas Gavroglu, editor, The Sciences in the European Periphery During the Enlightenment, volume 2 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 95–178. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 94-011-4770-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-011-4770-5_4.

Chabas:2003:AC

[CG03a] José Chabás and Bernard R. Goldstein. Astronomical commentary. In The Alfonsine Tables of Toledo, volume 8 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 135–224. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2003. ISBN 94-017-0213-6. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-017-0213-3_4.

Chabas:2003:BM

[CG03b] José Chabás and Bernard R. Goldstein. Back matter. In The Alfonsine Tables of Toledo, volume 8 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 307–341. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2003. ISBN 94-017-0213-6. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bbm:978-94-017-0213-3/1.pdf.

Chabas:2003:ATT

[CG03c] José Chabás and Bernard R. Goldstein. The Alfonsine Tables of Toledo, volume 8 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2003. ISBN 90-481-6404-4 (hardcover), 94-017-0213-6 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xiii + 341 pp. LCCN QB4.

Chabas:2003:C

[CG03d] José Chabás and Bernard R. Goldstein. Context. In The Alfonsine Tables of Toledo, volume 8 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 225–241.

Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2003. ISBN 94-017-0213-6. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-017-0213-3_5.

Chabas:2003:FM

[CG03e] José Chabás and Bernard R. Goldstein. Front matter. In The Alfonsine Tables of Toledo, volume 8 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xiii. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2003. ISBN 94-017-0213-6. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-017-0213-3/1.pdf.

Chabas:2003:GTT

[CG03f] José Chabás and Bernard R. Goldstein. A glossary of technical terms. In The Alfonsine Tables of Toledo, volume 8 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 95–133. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2003. ISBN 94-017-0213-6. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-017-0213-3_3.

Chabas:2003:I

[CG03g] José Chabás and Bernard R. Goldstein. Introduction. In The Alfonsine Tables of Toledo, volume 8 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–8. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2003. ISBN 94-017-0213-6. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-017-0213-3_1.

Chabas:2003:LAA

[CG03h] José Chabás and Bernard R. Goldstein. The legacy of Alfonsine astronomy. In The Alfonsine Tables of Toledo, volume 8 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 243–306. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2003. ISBN 94-017-0213-6. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-017-0213-3_6.

Chabas:2003:T

[CG03i] José Chabás and Bernard R. Goldstein. Text. In The Alfonsine Tables of Toledo, volume 8 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 9–94. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2003. ISBN 94-017-0213-6. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-017-0213-3_2.

Chabas:2006:USR

[Cha06a] José Chabás. The University of Salamanca and the renaissance of astronomy during the second half of the 15th Century. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 29–36. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3975-1_3.

Chauveau:2006:MIP

[Cha06b] Sophie Chauveau. Mobilization and industrial policy: Chemicals and pharmaceuticals in the French war effort. In Roy Macleod and Jeffrey Allan Johnson, editors, Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 21–30. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5490-4. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-5490-4_2.

Chang:2007:GES

[Cha07] Kevin Chang. Georg Ernst Stahl’s alchemical publications: Anachronism, reading market, and a scientific lineage redefined. In Lawrence M. Principe, editor, New Narratives in Eighteenth-Century Chemistry: Contributions from the First Francis Bacon Workshop, 21–23 April 2005, California Institute of Technology, Pasadena, California, volume 18 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 23–43. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2007. ISBN 1-4020-6278-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-6278-0_2.

Chalmers:2017:BM

[Cha17a] Alan F. Chalmers. Back matter. In One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 193– 197. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56529-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-3-319-56529-3/1.pdf.

Chalmers:2017:BAS

[Cha17b] Alan F. Chalmers. Beyond Archimedes: Stevin’s Elements of hydrostatics. In One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 27–48. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56529-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-56529-3_3.

Chalmers:2017:BMP

[Cha17c] Alan F. Chalmers. Boyle on mechanism and pressure. In One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 111– 133. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56529-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-56529-3_8.

Chalmers:2017:DEH

[Cha17d] Alan F. Chalmers. Descartes’ engagement with hydrostatics. In One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 59– 82. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56529-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-56529-3_5.

Chalmers:2017:EA

[Cha17e] Alan F. Chalmers. Experimenting with air. In One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 99–109. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56529-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-56529-3_7.

Chalmers:2017:FNC

[Cha17f] Alan F. Chalmers. Fashioning a novel concept of pressure: One hundred years. In One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 161–171. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56529-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-56529-3_10.

Chalmers:2017:FM

[Cha17g] Alan F. Chalmers. Front matter. In One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–ix. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56529-X. ISSN 1385-0180 (print), 2215-0064 (electronic). LCCN

????  URL https://link.
springer.com/content/pdf/
bfm:978-3-319-56529-3/1.pdf.

Chalmers:2017:GFB

[Cha17h] Alan F. Chalmers. Galileo, floating bodies and the balance. In One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 49– 57. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56529-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-56529-3_4.

Chalmers:2017:HBS

[Cha17i] Alan F. Chalmers. The historical background to Stevin’s hydrostatics. In One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 15–25. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56529-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-56529-3_2.

Chalmers:2017:HE

[Cha17j] Alan F. Chalmers. Hydrostatics and experiment. In One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 173–186. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56529-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-56529-3_11.

Chalmers:2017:HSR

[Cha17k] Alan F. Chalmers. Hydrostatics and the scientific revolution. In One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 187– 191. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56529-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-56529-3_12.

Chalmers:2017:LCS

[Cha17l] Alan F. Chalmers. Liquids: A challenge for Seventeenth-Century mechanics. In One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–14. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56529-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-56529-3_1.

Chalmers:2017:NHL

[Cha17m] Alan F. Chalmers. Newton’s hydrostatics: Liquids as continua. In One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 135– 160. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56529-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-56529-3_9.

Chalmers:2017:OHY

[Cha17n] Alan F. Chalmers. One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56528-1 (hardcover), 3-319-56529-X (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). ix + 197 pp. LCCN QA905 .C43 2017. URL http:/

/link.springer.com/10.1007/
978-3-319-56529-3.

Chalmers:2017:PEL

[Cha17o] Alan F. Chalmers. Pascal’s equilibrium of liquids. In One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 83–98. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56529-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-56529-3_6.

Chen:2005:VPE

[Che05] Xiang Chen. Visual photometry in the Early 19th Century: A “good” science with “wrong” measurements. In Jed Z. Buchwald and Allan Franklin, editors, Wrong for the Right Reasons, volume 11 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 161–183. Springer, Dordrecht, The Netherlands, 2005. ISBN 1-4020-3048-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3048-7_8.

Chemla:2015:DTA

[Che15] Karine Chemla. Describing texts for algorithms: How they prescribe operations and integrate cases. Reflections based on ancient Chinese mathematical sources. In Karine Chemla and Jacques Virbel, editors, Texts, Textual Acts and the History of Science, volume 42 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 317–384. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-16444-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-16444-1_9.

Carolin:2006:NPM

[CL06] Luís Miguel Carolin and Henrique Leitão. Natural philosophy and mathematics in Portuguese universities, 1550–1650. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 153–168. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3975-1_11.

Cocroft:2006:FWW

[Coc06] Wayne D. Cocroft. First World War explosives manufacture: The British experience. In Roy Macleod and Jeffrey Allan Johnson, editors, Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 31–46. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5490-4. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-5490-4_3.

Corry:2004:AMA

[Cor04a] Leo Corry. The axiomatic method in action: 1900–1905. In Leo Corry, editor, David Hilbert and the Axiomatization of Physics (1898–1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 119–184. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2004. ISBN 1-4020-2778-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-1-4020-2778-9_4.

Corry:2004:AHE

[Cor04b] Leo Corry. Axiomatization in Hilbert’s early career. In Leo Corry, editor, David Hilbert and the Axiomatization of Physics (1898–1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 83–118. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2004. ISBN 1-4020-2778-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-1-4020-2778-9_3.

Corry:2004:BM

[Cor04c] Leo Corry. Back matter. In David Hilbert and the Axiomatization of Physics (1898–1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 445–513. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2004. ISBN 1-4020-2778-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-1-4020-2778-9/1.pdf.

Corry:2004:DHA

[Cor04d] Leo Corry. David Hilbert and the Axiomatization of Physics (1898– 1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2004. ISBN 1-4020-2777-X (hardcover), 1-4020-2778-8 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xvii + 513 pp. LCCN QA29.H5

C67 2004. URL http://www.loc.
gov/catdir/enhancements/
fy0823/2006280837-d.html;
http://www.loc.gov/catdir/
enhancements/fy0823/
2006280837-t.html.

Corry:2004:EMT

[Cor04e] Leo Corry. Einstein and Mie: Two pillars of Hilbert’s unified theory. In Leo Corry, editor, David Hilbert and the Axiomatization of Physics (1898–1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes:

New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 287–316. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2004. ISBN 1-4020-2778-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-1-4020-2778-9_7.

Corry:2004:E

[Cor04f] Leo Corry. Epilogue. In Leo Corry, editor, David Hilbert and the Axiomatization of Physics (1898– 1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 409–443. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2004. ISBN 1-4020-2778-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-1-4020-2778-9_10.

Corry:2004:FP

[Cor04g] Leo Corry. Foundations of physics: 1915–1916. In Leo Corry, editor, David Hilbert and the Axiomatization of Physics (1898– 1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 317–362. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2004. ISBN 1-4020-2778-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-2778-9_8.

Corry:2004:MER

[Cor04h] Leo Corry. From mechanical to electromagnetic reductionism: 1910–1914. In Leo Corry, editor, David Hilbert and the Axiomatization of Physics (1898– 1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 227–285. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2004. ISBN 1-4020-2778-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-2778-9_6.

Corry:2004:FM

[Cor04i] Leo Corry. Front matter. In David Hilbert and the Axiomatization of Physics (1898–1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xvii. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2004. ISBN 1-4020-2778-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-1-4020-2778-9/1.pdf.

Corry:2004:HG

[Cor04j] Leo Corry. Hilbert and GTR:

1916–1918. In Leo Corry, editor, David Hilbert and the Axiomatization of Physics (1898– 1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 363–407. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2004. ISBN 1-4020-2778-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-2778-9_9.

Corry:2004:I

[Cor04k] Leo Corry. Introduction. In Leo Corry, editor, David Hilbert and the Axiomatization of Physics (1898–1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–9. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2004. ISBN 1-4020-2778-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-2778-9_1.

Corry:2004:LNC

[Cor04l] Leo Corry. Late nineteenth century background. In Leo Corry, editor, David Hilbert and the Axiomatization of Physics (1898– 1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 11–81. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2004. ISBN 1-4020-2778-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-2778-9_2.

Corry:2004:MR

[Cor04m] Leo Corry. Minkowski and relativity: 1907–1909. In Leo Corry, editor, David Hilbert and the Axiomatization of Physics (1898– 1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 185–225. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2004. ISBN 1-4020-2778-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-2778-9_5.

Crisciani:2017:MQC

[Cri17] Chiara Crisciani. Medicine as queen: The Consilia of Bartolomeo da Montagnana. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 79–96. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-56514-9_5.

Crouch:2000:BWO

[Cro00] Tom D. Crouch. Blaming Wilbur and Orville: The Wright patent suits and the growth of American aeronautics. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 287– 300. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-011-4379-0_11.

Chemla:2015:PTA

[CV15a] Karine Chemla and Jacques Virbel. Prologue: Textual acts and the history of science. In Karine Chemla and Jacques Virbel, editors, Texts, Textual Acts and the History of Science, volume 42 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–46. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-16444-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-16444-1_1.

Chemla:2015:TTA

[CV15b] Karine Chemla and Jacques Virbel, editors. Texts, Textual Acts and the History of Science, volume 42 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2015. ISBN 3-319-16443-0 (hardcover), 3-319-16444-9 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). ix + 430 pp. LCCN P302 .T48 2015.

deCeglia:2003:AIN

[dC03] Francesco Paolo de Ceglia. Additio illa non videtur edenda: Giuseppe Biancani, reader of Galileo in an unedited censored text. In Mordechai Feingold, editor, The New Science and Jesuit Science: Seventeenth Century Perspectives, volume 6 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 159–186. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2003. ISBN 94-017-0361-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-017-0361-1_3.

DelNegro:2006:VPT

[Del06] Piero Del Negro. Venetian policy toward the University of Padua and scientific progress during the 18th Century. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 169–181. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3975-1_12.

Derry:2015:CBC

[Der15] Margaret E. Derry. Chicken breeding: The complex transition from traditional to genetic methods in the usa. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 18, pages 371–393. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-12185-7_18.

Dialetis:1999:SGS

[DGP99] Dimitris Dialetis, Kostas Gavroglu, and Manolis Patiniotis. The sciences in the Greek speaking regions during the 17th and 18th Centuries. In Kostas Gavroglu, editor, The Sciences in the European Periphery During the Enlightenment, volume 2 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 41–71. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 94-011-4770-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-011-4770-5_2.

Dijksterhuis:2005:LWC

[Dij05] Fokko Jan Dijksterhuis, editor.

Lenses and Waves: Christiaan Huygens and the Mathematical Science of Optics in the Seventeenth Century, volume 9 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2005. ISBN 1-4020-2697-8 (hardcover), 1-4020-2698-8 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). vii + 287 pp. LCCN QC352 .D555 2004.

Dijksterhuis:2017:MNI

[Dij17] Fokko Jan Dijksterhuis. Magi from the north: Instruments of fire and light in the Early Seventeenth Century. In Arianna Borrelli, Giora Hon, and Yaakov Zik, editors, The Optics of Giambattista Della Porta (ca. 1535– 1615): A Reassessment, volume 44 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 125–143. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-50215-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-50215-1_6.

deJong:2016:BAP

[dJ16] Teije de Jong. Babylonian astronomy 1880–1950: The players and the field. In Alexander Jones, Christine Proust, and John M. Steele, editors, A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science, volume 45 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 265–302. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-25865-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-25865-2_9.

deJong-Lambert:2012:AC

[dL12a] William deJong Lambert. Acquired characteristics. In The Cold War Politics of Genetic Research: An Introduction to the Lysenko Affair, volume 32 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 85–109. Springer Netherlands, Dordrecht, The Netherlands, 2012. ISBN 94-007-2840-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-007-2840-0_4.

deJong-Lambert:2012:BM

[dL12b] William deJong Lambert. Back matter. In The Cold War Politics of Genetic Research: An Introduction to the Lysenko Affair, volume 32 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 173–185. Springer Netherlands, Dordrecht, The Netherlands, 2012. ISBN 94-007-2840-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bbm:978-94-007-2840-0/1.pdf.

deJong-Lambert:2012:BWW

[dL12c] William deJong Lambert. Between world war and cold war. In The Cold War Politics of Genetic Research: An Introduction to the Lysenko Affair, volume 32 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 29–56. Springer Netherlands, Dordrecht, The Netherlands, 2012. ISBN 94-007-2840-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-007-2840-0_2.

deJong-Lambert:2012:CWP

[dL12d] William deJong Lambert. The Cold War Politics of Genetic Research: an Introduction to the Lysenko Affair, volume 32 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2012. ISBN 94-007-2839-5 (hardcover), 94-007-2840-9 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xxii + 185 pp. LCCN T14.7 .A73 vol. 32; HQ755.5.S65 D4 2012.

deJong-Lambert:2012:FM

[dL12e] William deJong Lambert. Front matter. In The Cold War Politics of Genetic Research: An Introduction to the Lysenko Affair, volume 32 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xxii. Springer Netherlands, Dordrecht, The Netherlands, 2012. ISBN 94-007-2840-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-94-007-2840-0/1.pdf.

deJong-Lambert:2012:MMM

[dL12f] William deJong Lambert. Mendelist– Morganists and Michurinism. In The Cold War Politics of Genetic Research: An Introduction to the Lysenko Affair, volume 32 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–27. Springer Netherlands, Dordrecht, The Netherlands, 2012. ISBN 94-007-2840-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-007-2840-0_1.

deJong-Lambert:2012:PWB

[dL12g] William deJong Lambert. Pigs wearing booties earn scorn of red press. In The Cold War Politics of Genetic Research: An Introduction to the Lysenko Affair, volume 32 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 141– 172. Springer Netherlands, Dordrecht, The Netherlands, 2012. ISBN 94-007-2840-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2840-0_6.

deJong-Lambert:2012:RL

[dL12h] William deJong Lambert. Reacting to Lysenko. In The Cold War Politics of Genetic Research: An Introduction to the Lysenko Affair, volume 32 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 57–83. Springer Netherlands, Dordrecht, The Netherlands, 2012. ISBN 94-007-2840-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-007-2840-0_3.

deJong-Lambert:2012:SA

[dL12i] William deJong Lambert. The “Spitzer Affair”. In The Cold War Politics of Genetic Research: An Introduction to the Lysenko Affair, volume 32 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 111– 140. Springer Netherlands, Dordrecht, The Netherlands, 2012. ISBN 94-007-2840-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2840-0_5.

Douglas:2000:WDA

[Dou00] Deborah G. Douglas. Who designs airports … engineers, architects, or city planners? Aspects of American airport design before World War II. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 301–322. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-011-4379-0_12.

deSousa:2010:SDB

[dS10] Filomena de Sousa. Shaping disciplinary boundaries: Scientific practice and politics in the Methodenstreit between the German Historical School and the Austrian School of Economics. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 221–240. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-3540-0_12.

Ducheyne:2012:BM

[Duc12a] Steffen Ducheyne. Back matter.

In “The main Business of natural Philosophy”: Isaac Newton’s Natural-Philosophical Methodology, volume 29 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 305–352. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2126-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-94-007-2126-5/1.pdf.

Ducheyne:2012:C

[Duc12b] Steffen Ducheyne. Conclusion.

In “The main Business of natural Philosophy”: Isaac Newton’s Natural-Philosophical Methodology, volume 29 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 299–303. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2126-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2126-5_7.

Ducheyne:2012:FLD

[Duc12c] Steffen Ducheyne. Facing the limits of deductions from phenomena: Newton’s quest for a mathematical– demonstrative optics. In “The main Business of natural Philosophy”: Isaac Newton’s Natural-Philosophical Methodology, volume 29 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 179–224. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2126-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2126-5_4.

Ducheyne:2012:FMa

[Duc12d] Steffen Ducheyne. Front matter. In “The main Business of natural Philosophy”: Isaac Newton’s Natural-Philosophical Methodology, volume 29 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xxv. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2126-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-007-2126-5/1.pdf.

Ducheyne:2012:FMb

[Duc12e] Steffen Ducheyne. Front matter. In “The main Business of natural Philosophy”: Isaac Newton’s Natural-Philosophical Methodology, volume 29 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 1. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2126-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-007-2126-5/1/1.
pdf.

Ducheyne:2012:FMc

[Duc12f] Steffen Ducheyne. Front matter.

In “The main Business of natural Philosophy”: Isaac Newton’s Natural-Philosophical Methodology, volume 29 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 53. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2126-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-007-2126-5/2/1.
pdf.

Ducheyne:2012:FMd

[Duc12g] Steffen Ducheyne. Front matter.

In “The main Business of natural Philosophy”: Isaac Newton’s Natural-Philosophical Methodology, volume 29 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 233. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2126-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/

                                        bfm:978-94-007-2126-5/3/1.
                                        pdf.

Ducheyne:2012:NCS

[Duc12h] Steffen Ducheyne. Newton and causes: Something borrowed and something new. In “The main Business of natural Philosophy”: Isaac Newton’s Natural-Philosophical Methodology, volume 29 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 3–52. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2126-9. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-2126-5_1.

Ducheyne:2012:MBN

[Duc12i] Steffen Ducheyne. “The main Business of natural Philosophy”: Isaac Newton’s Natural-Philosophical Methodology, volume 29 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2012. ISBN 94-007-2125-0 (hardcover), 94-007-2126-9 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xxv + 352 pp. LCCN QC6 .D783 2012.

Ducheyne:2012:TGP

[Duc12j] Steffen Ducheyne. “To treat of God from phenomena”. In “The main Business of natural Philosophy”: Isaac Newton’s Natural-Philosophical Methodology, volume 29 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 235–297. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2126-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2126-5_6.

Ducheyne:2012:UMPa

[Duc12k] Steffen Ducheyne. Uncovering the methodology of the Principia (I): The phase of model construction. In “The main Business of natural Philosophy”: Isaac Newton’s Natural-Philosophical Methodology, volume 29 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 55–106. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2126-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2126-5_2.

Ducheyne:2012:UMPb

[Duc12l] Steffen Ducheyne. Uncovering the methodology of the Principia (II): The phase of model application, theory formation and theory application. In “The main Business of natural Philosophy”: Isaac Newton’s Natural-Philosophical Methodology, volume 29 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 107–177.

Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2126-9. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-2126-5_3.

Ducheyne:2012:UMPc

[Duc12m] Steffen Ducheyne. Uncovering the methodology of the Principia (III): A brief chronology of Newton’s methodological itinerary. In “The main Business of natural Philosophy”: Isaac Newton’s Natural-Philosophical Methodology, volume 29 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 225–232. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2126-9. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-2126-5_5.

Dupre:2014:LAA

[Dup14] Sven Dupré, editor. Laboratories of Art: Alchemy and Art Technology from Antiquity to the 18th Century, volume 37 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2014. ISBN 3-319-05064-8 (hardcover), 3-319-05065-6 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xxiii + 200 pp. LCCN N7745.A4 .L33 2014.

Eamon:2017:TEG

[Eam17] William Eamon. A theater of experiments: Giambattista Della Porta and the scientific culture of Renaissance Naples. In Arianna Borrelli, Giora Hon, and Yaakov Zik, editors, The Optics of Giambattista Della Porta (ca. 1535– 1615): A Reassessment, volume 44 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 11–38. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-50215-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-50215-1_2.

Eddy:2007:AAC

[Edd07] Matthew D. Eddy. The Aberdeen Agricola: Chemical principles and practice in James Anderson’s Georgics and Geology. In Lawrence M. Principe, editor, New Narratives in Eighteenth-Century Chemistry: Contributions from the First Francis Bacon Workshop, 21–23 April 2005, California Institute of Technology, Pasadena, California, volume 18 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 139–156. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2007. ISBN 1-4020-6278-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-6278-0_7.

Ede:2017:SBP

[Ede17] Andrew Ede. Science born of poison, fire and smoke: Chemical warfare and the origins of big science. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 181–200. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-58436-2_11.

Feest:2010:HPEb

[Fee10a] Uljana Feest, editor. Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2010. ISBN 90-481-3539-7 (hardcover), 90-481-3540-0 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). vi + 320 pp. LCCN BD237 .H57 2010.

Feest:2010:HPEa

[Fee10b] Uljana Feest. Historical perspectives on Erklären and Verstehen: Introduction. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–13.

Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-3540-0_1.

Feffer:1997:EAC

[Fef97] Stuart M. Feffer. Ernst Abbe, Carl Zeiss, and the transformation of microscopical optics. In Jed Z. Buchwald, editor, Scientific Credibility and Technical Standards in 19th and early 20th century Germany and Britain, volume 1 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 23–66. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 1997. ISBN 94-009-1784-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-009-1784-2_2.

Feingold:2003:GCG

[Fei03a] Mordechai Feingold. The grounds for conflict: Grienberger, Grassi, Galileo, and posterity. In Mordechai Feingold, editor, The New Science and Jesuit Science: Seventeenth Century Perspectives, volume 6 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 121–157. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2003. ISBN 94-017-0361-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-94-017-0361-1_2.

Feingold:2003:NSJ

[Fei03b] Mordechai Feingold, editor. The New Science and Jesuit Science: Seventeenth Century Perspectives, volume 6 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2003. ISBN 90-481-6115-0 (hardcover), 94-017-0361-2 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). ix + 270 pp. LCCN ????

Ferguson:2000:AME

[Fer00] Robert G. Ferguson. Airframe manufacture and engineering exchange. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 259– 285. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-011-4379-0_10.

Finlay:2015:SPS

[Fin15] Mark R. Finlay. Science, promotion, and scandal: Soil bacteriology, legume inoculation, and the American campaign for soil improvement in the Progressive Era. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 205–229. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-12185-7_11.

Findlen:2017:DNK

[Fin17] Paula Findlen. The death of a naturalist: Knowledge and community in Late Renaissance Italy. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 155–195. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-56514-9_9.

Fedotova:2015:FCR

[FL15] Anastasia A. Fedotova and Marina V. Loskutova. Forests, climate, and the rise of scientific forestry in Russia: From local knowledge and natural history to modern experiments (1840s– early 1890s). In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 113–137. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-12185-7_7.

Feingold:2006:USE

[FNB06] Mordechai Feingold and Victor Navarro-Brotons, editors. Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3974-3 (hardcover), 1-4020-3975-1 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). viii + 309 pp. LCCN Q181 .U59 2006.

Franklin:2005:KUT

[Fra05] Allan Franklin. The Konopinski– Uhlenbeck theory of β decay: Its proposal and refutation. In Jed Z. Buchwald and Allan Franklin, editors, Wrong for the Right Reasons, volume 11 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 209–228. Springer, Dordrecht, The Netherlands, 2005. ISBN 1-4020-3048-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3048-7_10.

Frangsmyr:2012:PMN

[Frä12] Tore Frängsmyr. Peking Man:

New light on an old discovery. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 49–62. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2627-7_5.

Friedman:2014:GSU

[Fri14] David Friedman. Geometric survey and urban design: A project for the Rome of Paul IV (1555– 1559). In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 107–134. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-05998-3_5.

Galison:2000:AH

[Gal00] Peter Galison. An accident of history. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 3–43. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-94-011-4379-0_1.

Gavroglu:1999:SEP

[Gav99] Kostas Gavroglu, editor. The Sciences in the European Periphery During the Enlightenment, volume 2 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 1999. ISBN 0-7923-6563-1 (hardcover), 94-011-4770-1 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xi + 224 pp. LCCN B802.

Goethe:2015:GWL

[GBR15a] Norma B. Goethe, Philip Beeley, and David Rabouin, editors. G. W. Leibniz, Interrelations between Mathematics and Philosophy, volume 41 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2015. ISBN 94-017-9663-7 (hardcover), 94-017-9664-5 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). x + 210 pp. LCCN B2598 .G85 2015.

Goethe:2015:IBM

[GBR15b] Norma B. Goethe, Philip Beeley, and David Rabouin. The interrelations between mathematics and philosophy in Leibniz’s thought. In Norma B. Goethe, Philip Beeley, and David Rabouin, editors, G. W. Leibniz, Interrelations between Mathematics and Philosophy, volume 41 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 3–21. Springer Netherlands, Dordrecht, The Netherlands, 2015. ISBN 94-017-9664-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-017-9664-4_1.

Gerbino:2014:GOA

[Ger14a] Anthony Gerbino, editor. Geometrical Objects: Architecture and the Mathematical Sciences 1400– 1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2014. ISBN 3-319-05997-1 (hardcover), 3-319-05998-X (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xiv + 318 pp. LCCN NA190-1555.5.

Gerbino:2014:I

[Ger14b] Anthony Gerbino. Introduction. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–41. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-05998-3_1.

Grattan-Guinness:2012:REP

[GG12] Ivor Grattan-Guinness. On the role of the ecole polytechnique, 1794–1914, with especial reference to mathematics. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 13, pages 217–234. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-2627-7_13.

Gillispie:2012:PLH

[Gil12] Charles Coulston Gillispie. A professional life in the history of science*. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 15–24. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2627-7_2.

Ginoux:2017:ANT

[Gin17a] Jean-Marc Ginoux. Andronov’s notes: Toward the concept of self-oscillations. In History of Nonlinear Oscillations Theory in France (1880–1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 131–144. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-55239-2_5.

Ginoux:2017:BMa

[Gin17b] Jean-Marc Ginoux. Back matter. In History of Nonlinear Oscillations Theory in France (1880– 1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 103–104. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2.

ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bbm:978-3-319-55239-2/1/1.
                                        pdf.

Ginoux:2017:BMb

[Gin17c] Jean-Marc Ginoux. Back matter. In History of Nonlinear Oscillations Theory in France (1880– 1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 257–258. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bbm:978-3-319-55239-2/2/1.
                                        pdf.

Ginoux:2017:FIC

[Gin17d] Jean-Marc Ginoux. The first international conference on nonlinear processes: Paris 1933. In History of Nonlinear Oscillations Theory in France (1880–1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 165–176. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-55239-2_7.

Ginoux:2017:QPF

[Gin17e] Jean-Marc Ginoux. From quasiperiodic functions to recurrent motions. In History of Nonlinear Oscillations Theory in France (1880– 1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 13, pages 311– 330. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-55239-2_13.

Ginoux:2017:SDM

[Gin17f] Jean-Marc Ginoux. From the series-dynamo machine to the singing arc: Gérard-Lescuyer, Blondel, Poincaré. In History of Nonlinear Oscillations Theory in France (1880–1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 3–37. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-55239-2_1.

Ginoux:2017:FMa

[Gin17g] Jean-Marc Ginoux. Front matter. In History of Nonlinear Oscillations Theory in France (1880– 1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xxxvii. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-3-319-55239-2/1.pdf.

Ginoux:2017:FMb

[Gin17h] Jean-Marc Ginoux. Front matter. In History of Nonlinear Oscillations Theory in France (1880– 1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 1. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL

                                        https://link.springer.com/
                                        content/pdf/bfm:978-3-319-
                                        55239-2/1/1.pdf.

Ginoux:2017:FMc

[Gin17i] Jean-Marc Ginoux. Front matter. In History of Nonlinear Oscillations Theory in France (1880– 1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 105–108. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-3-319-55239-2/2/1.
                                        pdf.

Ginoux:2017:FMd

[Gin17j] Jean-Marc Ginoux. Front matter. In History of Nonlinear Oscillations Theory in France (1880– 1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 259–263. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-3-319-55239-2/3/1.
pdf.

Ginoux:2017:GWF

[Gin17k] Jean-Marc Ginoux. The Great War and the first triode designs: Abraham, Bloch, Blondel, Van der Pol. In History of Nonlinear Oscillations Theory in France (1880–1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 39–65. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-55239-2_2.

Ginoux:2017:HHS

[Gin17l] Jean-Marc Ginoux. Hadamard and his seminary: At the crossroads of ideas and theories. In History of Nonlinear Oscillations Theory in France (1880–1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 14, pages 331– 338. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-55239-2_14.

Ginoux:2017:HNO

[Gin17m] Jean-Marc Ginoux. History of Nonlinear Oscillations Theory in France (1880–1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55238-4 (hardcover), 3-319-55239-2 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xxxvii + 381 pp. LCCN QA867.5 .G56 2017.

Ginoux:2017:KBM

[Gin17n] Jean-Marc Ginoux. The Krylov– Bogolyubov method: Towards a nonlinear mechanics. In History of Nonlinear Oscillations Theory in France (1880–1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 291–304. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-55239-2_11.

Ginoux:2017:MPS

[Gin17o] Jean-Marc Ginoux. The Mandel’shtam– Papalexi School: The “Van der Pol–Poincaré” method. In History of Nonlinear Oscillations Theory in France (1880–1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 305–310. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-55239-2_12.

Ginoux:2017:PRO

[Gin17p] Jean-Marc Ginoux. The paradigm of relaxation oscillations in France. In History of Nonlinear Oscillations Theory in France (1880– 1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 177– 255. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-55239-2_8.

Ginoux:2017:PLM

[Gin17q] Jean-Marc Ginoux. The Poincaré– Lindstedt method: The incompatibility with radio engineering. In History of Nonlinear Oscillations Theory in France (1880– 1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 265– 273. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-55239-2_9.

Ginoux:2017:RVP

[Gin17r] Jean-Marc Ginoux. Response to Van der Pol’s and Andronov’s work in France. In History of Nonlinear Oscillations Theory in France (1880–1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 145–163. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-55239-2_6.

Ginoux:2017:VPL

[Gin17s] Jean-Marc Ginoux. Van der Pol’s lectures: Towards the concept of relaxation oscillations. In History of Nonlinear Oscillations Theory in France (1880–1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 109–130. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-55239-2_4.

Ginoux:2017:VPM

[Gin17t] Jean-Marc Ginoux. Van der Pol’s method: A simple and classic solution. In History of Nonlinear Oscillations Theory in France (1880– 1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 275– 289. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-55239-2_10.

Ginoux:2017:VPP

[Gin17u] Jean-Marc Ginoux. Van der Pol’s prototype equation: Existence and uniqueness of the periodic solution: Cartan, Van der Pol, Liénard. In History of Nonlinear Oscillations Theory in France (1880–1940), volume 49 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 67–101. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-55239-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-55239-2_3.

Glatigny:2014:EOA

[Gla14] Pascal Dubourg Glatigny. Epistemological obstacles to the analysis of structures: Giovanni Bottari’s aversion to a mathematical assessment of Saint-Peter’s Dome (1743). In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 203–214. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-05998-3_9.

Gliboff:2015:BBP

[Gli15] Sander J. Gliboff. Breeding better peas, pumpkins, and peasants: The practical Mendelism of Erich Tschermak. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 19, pages 395–413. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-12185-7_19.

Grasshoff:2003:HKK

[GM03] Gerd Graßhoff and Michael May. Hans Krebs’ and Kurt Henseleit’s laboratory notebooks and their discovery of the urea cycle —- reconstructed with computer models. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 13, pages 269–294. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/0-306-48152-9_13.

Goeing:2014:ALS

[Goe14a] Anja-Silvia Goeing. Appendix:

The letter of Sassolo da Prato about Vittorino, translated into English by James Astorga. In Summus Mathematicus et Omnis Humanitatis Pater: The Vitae of Vittorino da Feltre and the Spirit of Humanism, volume 33 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 141–154. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-7531-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-7531-2_6.

Goeing:2014:BM

[Goe14b] Anja-Silvia Goeing. Back matter. In Summus Mathematicus et Omnis Humanitatis Pater: The Vitae of Vittorino da Feltre and the Spirit of Humanism, volume 33 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 155–176. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-7531-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bbm:978-94-007-7531-2/1.pdf.

Goeing:2014:BHP

[Goe14c] Anja-Silvia Goeing. Between history and praise: Approaches on understanding humanist biographies. In Summus Mathematicus et Omnis Humanitatis Pater: The Vitae of Vittorino da Feltre and the Spirit of Humanism, volume 33 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 135–140. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-7531-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-7531-2_5.

Goeing:2014:SMO

[Goe14d] Anja-Silvia Goeing. Summus Mathematicus et Omnis Humanitatis Pater: The Vitae of Vittorino da Feltre and the Spirit of Humanism, volume 33 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2014. ISBN 94-007-7530-X (hardcover), 94-007-7531-8 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). viii + 176 pp. LCCN HM623 .G64 2014.

Goeing:2014:CES

[Goe14e] Anja-Silvia Goeing. The concept of education in the second generation of Vitae and portraits of Vittorino da Feltre. In Summus Mathematicus et Omnis Humanitatis Pater: The Vitae of Vittorino da Feltre and the Spirit of Humanism, volume 33 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 95–133. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-7531-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-7531-2_4.

Goeing:2014:FM

[Goe14f] Anja-Silvia Goeing. Front matter. In Summus Mathematicus et Omnis Humanitatis Pater: The Vitae of Vittorino da Feltre and the Spirit of Humanism, volume 33 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–viii. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-7531-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-007-7531-2/1.pdf.

Goeing:2014:I

[Goe14g] Anja-Silvia Goeing. Introduction. In Summus Mathematicus et Omnis Humanitatis Pater: The Vitae of Vittorino da Feltre and the Spirit of Humanism, volume 33 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–11. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-7531-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-7531-2_1.

Goeing:2014:SPC

[Goe14h] Anja-Silvia Goeing. Sassolo da Prato’s correspondence with Leonardo Dati, ca. 1443–1444. In Summus Mathematicus et Omnis Humanitatis Pater: The Vitae of Vittorino da Feltre and the Spirit of Humanism, volume 33 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 57–93. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-7531-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-7531-2_3.

Goeing:2014:SVF

[Goe14i] Anja-Silvia Goeing. The sources on Vittorino da Feltre. In Summus Mathematicus et Omnis Humanitatis Pater: The Vitae of Vittorino da Feltre and the Spirit of Humanism, volume 33 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 13–55. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-7531-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-7531-2_2.

Goldschmidt:2019:BM

[Gol19a] Asaf Goldschmidt. Back matter. In Medical practice in twelfthcentury China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], pages 301–322. ISBN 3-030-06102-7. ISSN 1385-0180 (print), 2215-0064 (electronic).

LCCN ???? URL https://link.
springer.com/content/pdf/
bfm:978-3-030-06103-6/1.pdf.

Goldschmidt:2019:CNa

[Gol19b] Asaf Goldschmidt. Case number 1. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 2, pages 29–32. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_2.

Goldschmidt:2019:CNj

[Gol19c] Asaf Goldschmidt. Case number 10. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 11, pages 59–61. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_11.

Goldschmidt:2019:CNk

[Gol19d] Asaf Goldschmidt. Case number 11. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 12, pages 63–66. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

                                        ????  URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_12.

Goldschmidt:2019:CNl

[Gol19e] Asaf Goldschmidt. Case number 12. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 13, pages 67–68. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

                                        ????  URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_13.

Goldschmidt:2019:CNm

[Gol19f] Asaf Goldschmidt. Case number 13. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 14, pages 69–71. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

                                        ????  URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_14.

Goldschmidt:2019:CNn

[Gol19g] Asaf Goldschmidt. Case number 14. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 15, pages 73–75. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_15.

Goldschmidt:2019:CNo

[Gol19h] Asaf Goldschmidt. Case number 15. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 16, pages 77–79. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_16.

Goldschmidt:2019:CNp

[Gol19i] Asaf Goldschmidt. Case number 16. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 17, pages 81–83. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_17.

Goldschmidt:2019:CNq

[Gol19j] Asaf Goldschmidt. Case number 17. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 18, pages 85–87. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_18.

Goldschmidt:2019:CNr

[Gol19k] Asaf Goldschmidt. Case number 18. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 19, pages 89–91. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_19.

Goldschmidt:2019:CNs

[Gol19l] Asaf Goldschmidt. Case number 19. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 20, pages 93–94. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_20.

Goldschmidt:2019:CNb

[Gol19m] Asaf Goldschmidt. Case number 2. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 3, pages 33–34. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_3.

Goldschmidt:2019:CNt

[Gol19n] Asaf Goldschmidt. Case number 20. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 21, pages 95–96. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_21.

Goldschmidt:2019:CNu

[Gol19o] Asaf Goldschmidt. Case number 21. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 22, pages 97–99. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_22.

Goldschmidt:2019:CNv

[Gol19p] Asaf Goldschmidt. Case number 22. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 23, pages 101–103. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_23.

Goldschmidt:2019:CNw

[Gol19q] Asaf Goldschmidt. Case number 23. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 24, pages 105–107. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_24.

Goldschmidt:2019:CNx

[Gol19r] Asaf Goldschmidt. Case number 24. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 25, pages 109–110. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_25.

Goldschmidt:2019:CNy

[Gol19s] Asaf Goldschmidt. Case number 25. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 26, pages 111–112. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_26.

Goldschmidt:2019:CNz

[Gol19t] Asaf Goldschmidt. Case number 26. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 27, pages 113–114. CODEN ????

ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_27.

Goldschmidt:2019:CNba

[Gol19u] Asaf Goldschmidt. Case number 27. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 28, pages 115–116. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_28.

Goldschmidt:2019:CNbb

[Gol19v] Asaf Goldschmidt. Case number 28. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 29, pages 117–118. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_29.

Goldschmidt:2019:CNbc

[Gol19w] Asaf Goldschmidt. Case number 29. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 30, pages 119–120. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_30.

Goldschmidt:2019:CNc

[Gol19x] Asaf Goldschmidt. Case number 3. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 4, pages 35–36. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

                                        ????  URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_4.

Goldschmidt:2019:CNbd

[Gol19y] Asaf Goldschmidt. Case number 30. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 31, pages 121–125. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_31.

Goldschmidt:2019:CNbe

[Gol19z] Asaf Goldschmidt. Case number 31. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 32, pages 127–128. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_32.

Goldschmidt:2019:CNbf

[Gol19-27] Asaf Goldschmidt. Case number 32. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 33, pages 129–130. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_33.

Goldschmidt:2019:CNbg

[Gol19-28] Asaf Goldschmidt. Case number 33. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 34, pages 131–132. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_34.

Goldschmidt:2019:CNbh

[Gol19-29] Asaf Goldschmidt. Case number 34. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 35, pages 133–134. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_35.

Goldschmidt:2019:CNbi

[Gol19-30] Asaf Goldschmidt. Case number 35. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 36, pages 135–136. CODEN ????

ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_36.

Goldschmidt:2019:CNbj

[Gol19-31] Asaf Goldschmidt. Case number 36. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 37, pages 137–139. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_37.

Goldschmidt:2019:CNbk

[Gol19-32] Asaf Goldschmidt. Case number 37. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 38, pages 141–142. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_38.

Goldschmidt:2019:CNbl

[Gol19-33] Asaf Goldschmidt. Case number 38. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 39, pages 143–144. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_39.

Goldschmidt:2019:CNbm

[Gol19-34] Asaf Goldschmidt. Case number 39. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 40, pages 145–146. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_40.

Goldschmidt:2019:CNd

[Gol19-35] Asaf Goldschmidt. Case number 4. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 5, pages 37–40. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_5.

Goldschmidt:2019:CNbn

[Gol19-36] Asaf Goldschmidt. Case number 40. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 41, pages 147–148. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_41.

Goldschmidt:2019:CNbo

[Gol19-37] Asaf Goldschmidt. Case number 41. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 42, pages 149–150. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_42.

Goldschmidt:2019:CNbp

[Gol19-38] Asaf Goldschmidt. Case number 42. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 43, pages 151–152. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_43.

Goldschmidt:2019:CNbq

[Gol19-39] Asaf Goldschmidt. Case number 43. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 44, pages 153–155. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_44.

Goldschmidt:2019:CNbr

[Gol19-40] Asaf Goldschmidt. Case number 44. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 45, pages 157–159. CODEN ????

ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_45.

Goldschmidt:2019:CNbs

[Gol19-41] Asaf Goldschmidt. Case number 45. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 46, pages 161–163. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_46.

Goldschmidt:2019:CNbt

[Gol19-42] Asaf Goldschmidt. Case number 46. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 47, pages 165–166. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_47.

Goldschmidt:2019:CNbu

[Gol19-43] Asaf Goldschmidt. Case number 47. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 48, pages 167–168. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/book/10.1007/
978-3-030-06103-6.

Goldschmidt:2019:CNbv

[Gol19-44] Asaf Goldschmidt. Case number 48. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 49, pages 169–170. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_49.

Goldschmidt:2019:CNbw

[Gol19-45] Asaf Goldschmidt. Case number 49. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 50, pages 171–173. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_50.

Goldschmidt:2019:CNe

[Gol19-46] Asaf Goldschmidt. Case number 5. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 6, pages 41–42. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

                                        ????  URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_6.

Goldschmidt:2019:CNbx

[Gol19-47] Asaf Goldschmidt. Case number 50. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 51, pages 175–177. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_51.

Goldschmidt:2019:CNby

[Gol19-48] Asaf Goldschmidt. Case number 51. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 52, pages 179–181. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_52.

Goldschmidt:2019:CNbz

[Gol19-49] Asaf Goldschmidt. Case number 52. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 53, pages 183–186. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_53.

Goldschmidt:2019:CNca

[Gol19-50] Asaf Goldschmidt. Case number 53. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 54, pages 187–189. CODEN ????

ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_54.

Goldschmidt:2019:CNcb

[Gol19-51] Asaf Goldschmidt. Case number 54. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 55, pages 191–192. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_55.

Goldschmidt:2019:CNcc

[Gol19-52] Asaf Goldschmidt. Case number 55. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 56, pages 193–194. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_56.

Goldschmidt:2019:CNcd

[Gol19-53] Asaf Goldschmidt. Case number 56. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 57, pages 195–196. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_57.

Goldschmidt:2019:CNce

[Gol19-54] Asaf Goldschmidt. Case number 57. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 58, pages 197–198. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_58.

Goldschmidt:2019:CNcf

[Gol19-55] Asaf Goldschmidt. Case number 58. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 59, pages 199–200. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_59.

Goldschmidt:2019:CNcg

[Gol19-56] Asaf Goldschmidt. Case number 59. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 60, pages 201–202. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_60.

Goldschmidt:2019:CNf

[Gol19-57] Asaf Goldschmidt. Case number 6. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 7, pages 43–45. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_7.

Goldschmidt:2019:CNch

[Gol19-58] Asaf Goldschmidt. Case number 60. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 61, pages 203–204. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_61.

Goldschmidt:2019:CNci

[Gol19-59] Asaf Goldschmidt. Case number 61. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 62, pages 205–206. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_62.

Goldschmidt:2019:CNcj

[Gol19-60] Asaf Goldschmidt. Case number 62. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 63, pages 207–208. CODEN ????

ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_63.

Goldschmidt:2019:CNck

[Gol19-61] Asaf Goldschmidt. Case number 63. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 64, pages 209–210. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_64.

Goldschmidt:2019:CNcl

[Gol19-62] Asaf Goldschmidt. Case number 64. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 65, pages 211–212. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_65.

Goldschmidt:2019:CNcm

[Gol19-63] Asaf Goldschmidt. Case number 65. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 66, pages 213–216. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_66.

Goldschmidt:2019:CNcn

[Gol19-64] Asaf Goldschmidt. Case number 66. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 67, pages 217–219. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_67.

Goldschmidt:2019:CNco

[Gol19-65] Asaf Goldschmidt. Case number 67. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 68, pages 221–222. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_68.

Goldschmidt:2019:CNcp

[Gol19-66] Asaf Goldschmidt. Case number 68. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 69, pages 223–226. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_69.

Goldschmidt:2019:CNcq

[Gol19-67] Asaf Goldschmidt. Case number 69. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 70, pages 227–228. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_70.

Goldschmidt:2019:CNg

[Gol19-68] Asaf Goldschmidt. Case number 7. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 8, pages 47–49. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_8.

Goldschmidt:2019:CNcr

[Gol19-69] Asaf Goldschmidt. Case number 70. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 71, pages 229–230. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_71.

Goldschmidt:2019:CNcs

[Gol19-70] Asaf Goldschmidt. Case number 71. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 72, pages 231–232. CODEN ????

ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_72.

Goldschmidt:2019:CNct

[Gol19-71] Asaf Goldschmidt. Case number 72. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 73, pages 233–234. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_73.

Goldschmidt:2019:CNcu

[Gol19-72] Asaf Goldschmidt. Case number 73. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 74, pages 235–237. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_74.

Goldschmidt:2019:CNcv

[Gol19-73] Asaf Goldschmidt. Case number 74. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 75, pages 239–242. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_75.

Goldschmidt:2019:CNcw

[Gol19-74] Asaf Goldschmidt. Case number 75. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 76, pages 243–244. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_76.

Goldschmidt:2019:CNcx

[Gol19-75] Asaf Goldschmidt. Case number 76. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 77, pages 245–247. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_77.

Goldschmidt:2019:CNcy

[Gol19-76] Asaf Goldschmidt. Case number 77. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 78, pages 249–250. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_78.

Goldschmidt:2019:CNcz

[Gol19-77] Asaf Goldschmidt. Case number 78. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 79, pages 251–253. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_79.

Goldschmidt:2019:CNda

[Gol19-78] Asaf Goldschmidt. Case number 79. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 80, pages 255–257. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_80.

Goldschmidt:2019:CNh

[Gol19-79] Asaf Goldschmidt. Case number 8. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 9, pages 51–54. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_9.

Goldschmidt:2019:CNdb

[Gol19-80] Asaf Goldschmidt. Case number 80. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 81, pages 259–261. CODEN ????

ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_81.

Goldschmidt:2019:CNdc

[Gol19-81] Asaf Goldschmidt. Case number 81. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 82, pages 263–266. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_82.

Goldschmidt:2019:CNdd

[Gol19-82] Asaf Goldschmidt. Case number 82. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 83, pages 267–270. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_83.

Goldschmidt:2019:CNde

[Gol19-83] Asaf Goldschmidt. Case number 83. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 84, pages 271–274. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_84.

Goldschmidt:2019:CNdf

[Gol19-84] Asaf Goldschmidt. Case number 84. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 85, pages 275–277. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_85.

Goldschmidt:2019:CNdg

[Gol19-85] Asaf Goldschmidt. Case number 85. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 86, pages 279–281. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_86.

Goldschmidt:2019:CNdh

[Gol19-86] Asaf Goldschmidt. Case number 86. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 87, pages 283–285. CODEN ???? ISBN 3-030-06102-7. ISSN ????

                                        LCCN ???? URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-06103-6_87.

Goldschmidt:2019:CNdi

[Gol19-87] Asaf Goldschmidt. Case number 87. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 88, pages 287–289. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_88.

Goldschmidt:2019:CNdj

[Gol19-88] Asaf Goldschmidt. Case number 88. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 89, pages 291–293. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_89.

Goldschmidt:2019:CNdk

[Gol19-89] Asaf Goldschmidt. Case number 89. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 90, pages 295–297. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_90.

Goldschmidt:2019:CNi

[Gol19-90] Asaf Goldschmidt. Case number 9. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 10, pages 55–57. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_10.

Goldschmidt:2019:CNdl

[Gol19-91] Asaf Goldschmidt. Case number 90. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 91, pages 299–300. CODEN ???? ISBN 3-030-06102-7. ISSN ????

LCCN ???? URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_91.

Goldschmidt:2019:FMa

[Gol19-92] Asaf Goldschmidt. Front matter. In Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], pages i– xxiv. ISBN 3-030-06102-7. ISSN 1385-0180 (print), 2215-0064 (electronic). LCCN ???? URL https:

//link.springer.com/book/10.
1007/978-3-030-06103-6.

Goldschmidt:2019:FMb

[Gol19-93] Asaf Goldschmidt. Front matter. In Medical practice in twelfthcentury China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], page 1. ISBN 3-030-06102-7. ISSN 1385-0180 (print), 2215-0064 (electronic). LCCN

????  URL https://link.
springer.com/content/pdf/
bfm:978-3-030-06103-6/1.pdf.

Goldschmidt:2019:FMc

[Gol19-94] Asaf Goldschmidt. Front matter. In Medical practice in twelfthcentury China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], page 27. ISBN 3-030-06102-7. ISSN 1385-0180 (print), 2215-0064 (electronic). LCCN

????  URL https://link.
springer.com/content/pdf/
bfm:978-3-030-06103-6/1.pdf.

Goldschmidt:2019:I

[Gol19-95] Asaf Goldschmidt. Introduction.

In Medical practice in twelfthcentury China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders [Gol19-96], chapter 1, pages 3– 25. CODEN ???? ISBN 3-030-06102-7. ISSN ???? LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-3-030-06103-6_1.

Goldschmidt:2019:MPT

[Gol19-96] Asaf Goldschmidt. Medical practice in twelfth-century China: a translation of Xu Shuwei’s ninety discussions [cases] on cold damage disorders, volume 54 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2019. ISBN 3-030-06102-7. ISSN 1385-0180 (print), 2215-0064 (electronic). 315 (est.) pp.

LCCN ???? URL https://link.
springer.com/content/pdf/
bfm:978-3-030-06103-6/1.pdf.

Good:2017:JHG

[Goo17] Gregory A. Good. John Herschel’s geology: The Cape of Good Hope in the 1830s. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 135–150. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-58436-2_8.

Gorman:2003:MMS

[Gor03] Michael John Gorman. Mathematics and modesty in the Society of Jesus: The problems of Christoph Grienberger. In Mordechai Feingold, editor, The New Science and Jesuit Science: Seventeenth Century Perspectives, volume 6 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–120. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2003. ISBN 94-017-0361-2. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-017-0361-1_1.

Gordin:2012:WCT

[Gor12] Michael D. Gordin. The weekday chemist: The training of Aleksandr Borodin. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 137–164. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2627-7_9.

Goulding:2010:BM

[Gou10a] Robert Goulding. Back matter. In Defending Hypatia: Ramus, Savile, and the Renaissance Rediscovery of Mathematical History, volume 25 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 179–201. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3542-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-90-481-3542-4/1.pdf.

Goulding:2010:DHR

[Gou10b] Robert Goulding. Defending Hypatia: Ramus, Savile, and the Renaissance Rediscovery of Mathematical History, volume 25 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2010. ISBN 90-481-3541-9 (hardcover), 90-481-3542-7 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xx + 201 pp. LCCN QA24 .G68 2009.

Goulding:2010:PPS

[Gou10c] Robert Goulding. From Plato to Pythagoras: The Scholae mathematicae. In Defending Hypatia: Ramus, Savile, and the Renaissance Rediscovery of Mathematical History, volume 25 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 35–74. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3542-7. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-3542-4_3.

Goulding:2010:FM

[Gou10d] Robert Goulding. Front matter. In Defending Hypatia: Ramus, Savile, and the Renaissance Rediscovery of Mathematical History, volume 25 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xx. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3542-7. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bfm:978-90-481-3542-4/1.pdf.

Goulding:2010:LL

[Gou10e] Robert Goulding. Lineages of learning. In Defending Hypatia: Ramus, Savile, and the Renaissance Rediscovery of Mathematical History, volume 25 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–18. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3542-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-3542-4_1.

Goulding:2010:PLE

[Gou10f] Robert Goulding. The puzzling lives of Euclid. In Defending Hypatia: Ramus, Savile, and the Renaissance Rediscovery of Mathematical History, volume 25 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 117–142. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3542-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-3542-4_5.

Goulding:2010:RHM

[Gou10g] Robert Goulding. Ramus and the history of mathematics. In Defending Hypatia: Ramus, Savile, and the Renaissance Rediscovery of Mathematical History, volume 25 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 19–33. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3542-7. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-3542-4_2.

Goulding:2010:RHB

[Gou10h] Robert Goulding. Rending Hypatia: The body of the elements. In Defending Hypatia: Ramus, Savile, and the Renaissance Rediscovery of Mathematical History, volume 25 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 143–178. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3542-7. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-3542-4_6.

Goulding:2010:BAO

[Gou10i] Robert Goulding. “To bring Alexandria to Oxford:” Henry Savile’s 1570 lectures on Ptolemy. In Defending Hypatia: Ramus, Savile, and the Renaissance Rediscovery of Mathematical History, volume 25 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 75–115. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3542-7. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-3542-4_4.

Galison:2000:AFT

[GR00] Peter Galison and Alex Roland, editors. Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2000. ISBN 0-7923-6037-0 (hardcover), 0-7923-6742-1 (paperback), 94-011-4379-X (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xvi + 383 pp. LCCN ????

Granada:2011:JKD

[Gra11] Miguel A. Granada. Johannes Kepler and David Fabricius: Their discussion on the Nova of 1604. In Patrick J. Boner, editor, Change and Continuity in Early Modern Cosmology, volume 27 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 67–92. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 94-007-0037-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-0037-6_5.

Grafton:2017:MMA

[Gra17] Anthony Grafton. A medical man among ecclesiastical historians: John Caius, Matthew Parker and the history of Cambridge University. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 113–127. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-56514-9_7.

Grene:2005:DHB

[Gre05] Marjorie Grene. Descartes and the heart beat: A conservative innovation. In Jed Z. Buchwald and Allan Franklin, editors, Wrong for the Right Reasons, volume 11 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 91–97. Springer, Dordrecht, The Netherlands, 2005. ISBN 1-4020-3048-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-3048-7_5.

Grosholz:2015:LMP

[Gro15] Emily R. Grosholz. Leibniz’s mathematical and philosophical analysis of time. In Norma B. Goethe, Philip Beeley, and David Rabouin, editors, G. W. Leibniz, Interrelations between Mathematics and Philosophy, volume 41 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 75–88. Springer Netherlands, Dordrecht, The Netherlands, 2015. ISBN 94-017-9664-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-94-017-9664-4_4.

Guttler:2015:DLM

[Güt15] Nils Güttler. Drawing the line:

Mapping cultivated plants and seeing nature in Nineteenth-Century plant geography. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 27–52. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-12185-7_3.

Hamerla:2006:ASR

[Ham06] Ralph R. Hamerla, editor. An American Scientist on the Research Frontier: Edward Morley, Community, and Radical Ideas in Nineteenth-Century Science, volume 13 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4088-1 (hardcover), 1-4020-4089-X (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xiii + 260 pp. LCCN Q143.M82 H36 2006.

Hamm:2017:TLA

[Ham17] Ernie Hamm. Trevor Levere, affinities that matter. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–7. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-58436-2_1.

Harwood:2015:DMT

[Har15] Jonathan Harwood. Did Mendelism transform plant breeding? Genetic theory and breeding practice, 1900–1945. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 17, pages 345–370. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-12185-7_17.

Hashimoto:2000:WTE

[Has00] Takehiko Hashimoto. The wind tunnel and the emergence of aeronautical research in Britain. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 223–239. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-011-4379-0_8.

Hatch:2011:DMC

[Hat11] Robert Alan Hatch. Discovering Mira Ceti: Celestial change and cosmic continuity. In Patrick J. Boner, editor, Change and Continuity in Early Modern Cosmology, volume 27 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 153–176. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 94-007-0037-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-0037-6_9.

Haug:2014:AIN

[Hau14] Henrike Haug. Artificial interventions in the natural form of things: Shared metallogenetical concepts of goldsmiths and alchemists. In Sven Dupré, editor, Laboratories of Art: Alchemy and Art Technology from Antiquity to the 18th Century, volume 37 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 79–103. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05065-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-05065-2_4.

Heeffer:2017:UIG

[Hee17] Albrecht Heeffer. Using invariances in geometrical diagrams: Della Porta, Kepler and Descartes on refraction. In Arianna Borrelli, Giora Hon, and Yaakov Zik, editors, The Optics of Giambattista Della Porta (ca. 1535–1615): A Reassessment, volume 44 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 145–168. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-50215-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-50215-1_7.

Heidarzadeh:2008:AA

[Hei08a] Tofigh Heidarzadeh. After Aristotle. In A History of Physical Theories of Comets, From Aristotle to Whipple, volume 19 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 21–51. Springer Netherlands, Dordrecht, The Netherlands, 2008. ISBN 1-4020-8323-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-1-4020-8323-5_2.

Heidarzadeh:2008:AN

[Hei08b] Tofigh Heidarzadeh. After Newton. In A History of Physical Theories of Comets, From Aristotle to Whipple, volume 19 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 125–150. Springer Netherlands, Dordrecht, The Netherlands, 2008. ISBN 1-4020-8323-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-1-4020-8323-5_5.

Heidarzadeh:2008:ATC

[Hei08c] Tofigh Heidarzadeh. Aristotle’s theory of comets. In A History of Physical Theories of Comets, From Aristotle to Whipple, volume 19 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–19. Springer Netherlands, Dordrecht, The Netherlands, 2008. ISBN 1-4020-8323-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-1-4020-8323-5_1.

Heidarzadeh:2008:BM

[Hei08d] Tofigh Heidarzadeh. Back matter. In A History of Physical Theories of Comets, From Aristotle to Whipple, volume 19 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 253– 275. Springer Netherlands, Dordrecht, The Netherlands, 2008. ISBN 1-4020-8323-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bbm:978-1-4020-8323-5/1.pdf.

Heidarzadeh:2008:CNP

[Hei08e] Tofigh Heidarzadeh. Comets in Newtonian physics. In A History of Physical Theories of Comets, From Aristotle to Whipple, volume 19 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 89– 124. Springer Netherlands, Dordrecht, The Netherlands, 2008. ISBN 1-4020-8323-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-1-4020-8323-5_4.

Heidarzadeh:2008:CLC

[Hei08f] Tofigh Heidarzadeh. Comets in the Laplacian cosmos. In A History of Physical Theories of Comets, From Aristotle to Whipple, volume 19 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 187– 209. Springer Netherlands, Dordrecht, The Netherlands, 2008. ISBN 1-4020-8323-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-1-4020-8323-5_7.

Heidarzadeh:2008:CPL

[Hei08g] Tofigh Heidarzadeh. Comets in the post–Laplacian era. In A History of Physical Theories of Comets, From Aristotle to Whipple, volume 19 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 211– 251. Springer Netherlands, Dordrecht, The Netherlands, 2008. ISBN 1-4020-8323-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-8323-5_8.

Heidarzadeh:2008:BN

[Hei08h] Tofigh Heidarzadeh. From Brahe to Newton. In A History of Physical Theories of Comets, From Aristotle to Whipple, volume 19 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 53–87. Springer Netherlands, Dordrecht, The Netherlands, 2008. ISBN 1-4020-8323-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-1-4020-8323-5_3.

Heidarzadeh:2008:FM

[Hei08i] Tofigh Heidarzadeh. Front matter. In A History of Physical Theories of Comets, From Aristotle to Whipple, volume 19 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i– xiii. Springer Netherlands, Dordrecht, The Netherlands, 2008. ISBN 1-4020-8323-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-1-4020-8323-5/1.pdf.

Heidarzadeh:2008:HPT

[Hei08j] Tofigh Heidarzadeh. A History of Physical Theories of Comets, From Aristotle to Whipple, volume 19 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2008. ISBN 1-4020-8322-X (hardcover), 1-4020-8323-8 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xiii + 275 pp. LCCN QB721.H45 2008.

Heidarzadeh:2008:NNT

[Hei08k] Tofigh Heidarzadeh. Non– Newtonian theories of comets. In A History of Physical Theories of Comets, From Aristotle to Whipple, volume 19 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 151–186. Springer Netherlands, Dordrecht, The Netherlands, 2008. ISBN 1-4020-8323-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-1-4020-8323-5_6.

Heidelberger:2010:MKM

[Hei10a] Michael Heidelberger. From Mill via von Kries to Max Weber: Causality, explanation, and understanding. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 13, pages 241–265. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-3540-0_13.

Heilen:2010:PDT

[Hei10b] Stephan Heilen. Ptolemy’s doctrine of the terms and its reception. In Alexander Jones, editor, Ptolemy in Perspective: Use and Criticism of his Work from Antiquity to the Nineteenth Century, volume 23 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 45–93. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-2788-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-2788-7_3.

Hessenbruch:1999:SPM

[Hes99] Arne Hessenbruch. The spread of precision measurement in Scandinavia 1660–1800. In Kostas Gavroglu, editor, The Sciences in the European Periphery During the Enlightenment, volume 2 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 179–224. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 94-011-4770-1. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-011-4770-5_5.

Heyman:2014:GMA

[Hey14] Jacques Heyman. Geometry, mechanics, and analysis in architecture. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 193–201. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-05998-3_8.

Hon:2008:AP

[HG08a] Giora Hon and Bernard R. Goldstein. The aesthetic path. In From Summetria to Symmetry: The Making of a Revolutionary Scientific Concept, volume 20 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 93–109. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2008. ISBN 1-4020-8448-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-1-4020-8448-5_3.

Hon:2008:ACS

[HG08b] Giora Hon and Bernard R. Goldstein. The ancient concept of symmetry in scientific contexts in early modern times and its association with harmony. In From Summetria to Symmetry: The Making of a Revolutionary Scientific Concept, volume 20 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 157–176. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2008. ISBN 1-4020-8448-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-8448-5_5.

Hon:2008:BM

[HG08c] Giora Hon and Bernard R. Goldstein. Back matter. In From Summetria to Symmetry: The Making of a Revolutionary Scientific Concept, volume 20 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 295–335. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2008. ISBN 1-4020-8448-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-1-4020-8448-5/1.pdf.

Hon:2008:SSM

[HG08d] Giora Hon and Bernard R. Goldstein. From Summetria to Symmetry: The Making of a Revolutionary Scientific Concept, volume 20 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2008. ISBN 1-4020-8447-1 (hardcover), 1-4020-8448-X (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xvi + 335 pp. LCCN Q172.5.S95 H66 2008; T14.7 .A73 vol. 20.

Hon:2008:FMa

[HG08e] Giora Hon and Bernard R. Goldstein. Front matter. In From Summetria to Symmetry: The Making of a Revolutionary Scientific Concept, volume 20 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages ii–xvi. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2008. ISBN 1-4020-8448-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bfm:978-1-4020-8448-5/1.pdf.

Hon:2008:FMb

[HG08f] Giora Hon and Bernard R. Goldstein. Front matter. In From Summetria to Symmetry: The Making of a Revolutionary Scientific Concept, volume 20 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 1. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2008. ISBN 1-4020-8448-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bfm:978-1-4020-8448-5/1/1.
                                        pdf.

Hon:2008:FMc

[HG08g] Giora Hon and Bernard R. Goldstein. Front matter. In From Summetria to Symmetry: The Making of a Revolutionary Scientific Concept, volume 20 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 67. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2008. ISBN 1-4020-8448-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-1-4020-8448-5/2/1.
pdf.

Hon:2008:FMd

[HG08h] Giora Hon and Bernard R. Goldstein. Front matter. In From Summetria to Symmetry: The Making of a Revolutionary Scientific Concept, volume 20 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 177. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2008. ISBN 1-4020-8448-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-1-4020-8448-5/3/1.
pdf.

Hon:2008:I

[HG08i] Giora Hon and Bernard R. Goldstein. Introduction. In From Summetria to Symmetry: The Making of a Revolutionary Scientific Concept, volume 20 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–65. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2008. ISBN 1-4020-8448-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-1-4020-8448-5_1.

Hon:2008:LRD

[HG08j] Giora Hon and Bernard R. Goldstein. Legendre’s revolutionary definition of symmetry as a scientific concept (1794). In From Summetria to Symmetry: The Making of a Revolutionary Scientific Concept, volume 20 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 221–261. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2008. ISBN 1-4020-8448-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-1-4020-8448-5_8.

Hon:2008:LEI

[HG08k] Giora Hon and Bernard R. Goldstein. Leonhard Euler (1707– 1783) and Immanuel Kant (1724– 1804): “revolutions” that did not happen. In From Summetria to Symmetry: The Making of a Revolutionary Scientific Concept, volume 20 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 203–219. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2008. ISBN 1-4020-8448-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.

springer.com/chapter/10.
1007/978-1-4020-8448-5_7.

Hon:2008:MP

[HG08l] Giora Hon and Bernard R. Goldstein. The mathematical path. In From Summetria to Symmetry: The Making of a Revolutionary Scientific Concept, volume 20 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 69–91. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2008. ISBN 1-4020-8448-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-8448-5_2.

Hon:2008:NASa

[HG08m] Giora Hon and Bernard R. Goldstein. New aesthetic sensibilities in Italian and French architecture. In From Summetria to Symmetry: The Making of a Revolutionary Scientific Concept, volume 20 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 111–155. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2008. ISBN 1-4020-8448-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-8448-5_4.

Hon:2008:NASb

[HG08n] Giora Hon and Bernard R. Goldstein. New applications of symmetry in mathematics and physics:

1788–1815. In From Summetria to Symmetry: The Making of a Revolutionary Scientific Concept, volume 20 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 263–293. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2008. ISBN 1-4020-8448-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-8448-5_9.

Hon:2008:TSN

[HG08o] Giora Hon and Bernard R. Goldstein. The treatment of symmetry in natural history (1738– 1815). In From Summetria to Symmetry: The Making of a Revolutionary Scientific Concept, volume 20 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 179–201. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2008. ISBN 1-4020-8448-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-8448-5_6.

Hoyningen-Huene:2006:CDV

[HH06] Paul Hoyningen-Huene. Context of discovery versus context of justification and Thomas Kuhn. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 119–131. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-4251-5_8.

Hiebert:2014:BCN

[Hie14a] Erwin Hiebert. Berlin: Capital of the New German Reich. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 16, pages 125–130. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-06602-8_16.

Hiebert:2014:EEM

[Hie14b] Erwin Hiebert. Early experience in music-making. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 27– 32. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-06602-8_4.

Hiebert:2014:EHG

[Hie14c] Erwin Hiebert. Encounter with the Helmholtz Group in Berlin. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 85–89. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-06602-8_10.

Hiebert:2014:EMT

[Hie14d] Erwin Hiebert. Encounters with music theorists in Japan. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 15, pages 111–122. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-06602-8_15.

Hiebert:2014:E

[Hie14e] Erwin Hiebert. The enharmonium. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 95– 97. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-06602-8_12.

Hiebert:2014:EOE

[Hie14f] Erwin Hiebert. Expert opinions. Evaluating the enharmonium. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 13, pages 99–106. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-06602-8_13.

Hiebert:2014:PEC

[Hie14g] Erwin Hiebert. From physiology to energy conservation. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 21–25. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-06602-8_3.

Hiebert:2014:FMa

[Hie14h] Erwin Hiebert. Front matter. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xxiii. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-3-319-06602-8/1.pdf.

Hiebert:2014:FMb

[Hie14i] Erwin Hiebert. Front matter. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 1. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (electronic).

URL https://link.springer.
com/content/pdf/bfm:978-3-
319-06602-8/1/1.pdf.

Hiebert:2014:FMc

[Hie14j] Erwin Hiebert. Front matter. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 79. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (electronic).

URL https://link.springer.
com/content/pdf/bfm:978-3-
319-06602-8/2/1.pdf.

Hiebert:2014:FMd

[Hie14k] Erwin Hiebert. Front matter. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 123. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (electronic).

URL https://link.springer.
com/content/pdf/bfm:978-3-
319-06602-8/3/1.pdf.

Hiebert:2014:HLP

[Hie14l] Erwin Hiebert. The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2014. ISBN 3-319-06601-3 (hardcover), 3-319-06602-1 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xxiii + 269 pp. LCCN ML3805 .H54 2014.

Hiebert:2014:HSJ

[Hie14m] Erwin Hiebert. The Helmholtz setting in the Johannes Müller Circle in Berlin. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 3–10. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-06602-8_1.

Hiebert:2014:I

[Hie14n] Erwin Hiebert. Introduction. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 81–84. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-06602-8_9.

Hiebert:2014:JIH

[Hie14o] Erwin Hiebert. Just intonation and the harmonium. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 59–77. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-06602-8_8.

Hiebert:2014:NDP

[Hie14p] Erwin Hiebert. New directions in physiology in the Johannes Müller Circle in Berlin. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 11–19. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-06602-8_2.

Hiebert:2014:PS

[Hie14q] Erwin Hiebert. The Papendick Sammlung. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 91– 93. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-06602-8_11.

Hiebert:2014:PAC

[Hie14r] Erwin Hiebert. Physiological acoustics and combination tones. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 33–38. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-06602-8_5.

Hiebert:2014:PCM

[Hie14s] Erwin Hiebert. Physiological causes of musical harmony. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 39–45. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-06602-8_6.

Hiebert:2014:STP

[Hie14t] Erwin Hiebert. Sensations of tone as the physiological basis for the theory of music. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 47–57. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-06602-8_7.

Hiebert:2014:BV

[Hie14u] Erwin Hiebert. With Bruckner in Vienna. In The Helmholtz Legacy in Physiological Acoustics, volume 39 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 14, pages 107–110.

Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-06602-1. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-06602-8_14.

Hirai:2017:IMD

[Hir17] Hiro Hirai. Imagination, maternal desire and embryology in Thomas Fienus. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 239–253. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-56514-9_11.

Hoffmann:2003:PS

[Hof03] Christoph Hoffmann. The pocket schedule. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 183–202. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9.

ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/0-306-48152-9_9.

Holmes:2003:LNI

[Hol03] Frederic L. Holmes. Laboratory notebooks and investigative pathways. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 14, pages 295–308. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/0-306-48152-9_14.

Homburg:2006:OSF

[Hom06] Ernst Homburg. Operating on several fronts: The trans-national activities of Royal Dutch/Shell, 1914–1918. In Roy Macleod and Jeffrey Allan Johnson, editors, Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 123–144. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5490-4. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-5490-4_8.

Hong:1997:SCJ

[Hon97] Sungook Hong. Syntony and credibility: John Ambrose Fleming, Guglielmo Marconi, and the Maskelyne Affair. In Jed Z. Buchwald, editor, Scientific Credibility and Technical Standards in 19th and early 20th century Germany and Britain, volume 1 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 157–176. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 1997. ISBN 94-009-1784-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-009-1784-2_5.

Horan:2015:NCC

[Hor15] Joseph Horan. Napoleonic cotton cultivation: A case study in scientific expertise and agricultural innovation in France and Italy, 1806–1814. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 73–91. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-12185-7_5.

Howard:2006:LWF

[How06] Don Howard. Lost wanderers in the forest of knowledge: Some thoughts on the discovery–justification distinction. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 3–22. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-4251-5_2.

Hoyrup:2016:OWH

[Høy16] Jens Høyrup. As the outsider walked in the historiography of Mesopotamian mathematics until Neugebauer. In Alexander Jones, Christine Proust, and John M. Steele, editors, A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science, volume 45 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 165–195. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-25865-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-25865-2_5.

Holmes:2003:RBR

[HRR03] Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors. Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1039-7 (hardcover), 0-306-48152-9 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xv + 325 pp. LCCN Q225.5 .R475

2003. URL http://www.loc.gov/
catdir/enhancements/fy0814/
2002041330-d.html;
http://www.loc.gov/catdir/
enhancements/fy0814/
2002041330-t.html.

Hubbard:2017:FEE

[Hub17] Jennifer Hubbard. Fishing an extreme environment: Science, sovereignty and Hudson Bay. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 13, pages 219– 254. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-58436-2_13.

Huppert:2014:PMD

[Hup14] Ann C. Huppert. Practical mathematics in the drawings of Baldassarre Peruzzi and Antonio da Sangallo the Younger. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 79–106. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-05998-3_4.

Jackson:1997:BDL

[Jac97] Myles W. Jackson. Buying the dark lines of the solar spectrum: Joseph Von Fraunhofer’s standard for the manufacture of optical glass. In Jed Z. Buchwald, editor, Scientific Credibility and Technical Standards in 19th and early 20th century Germany and Britain, volume 1 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–22. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 1997. ISBN 94-009-1784-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-009-1784-2_1.

Jacquart:2017:MST

[Jac17] Danielle Jacquart. Medicine in some Thirteenth-Century biblical commentaries, with a flashback on Augustine’s De genesi ad litteram. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 31–56. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-56514-9_3.

Jahnert:2019:BM

[Jäh19a] Martin Jähnert. Back matter. In Practicing the Correspondence Principle in the Old Quantum Theory: a Transformation through Implementation [Jäh19i], pages 271–293. ISBN 3-030-13299-4, 3-030-13300-1 (e-book), 3-030-13301-X, 3-030-13302-8. LCCN QC174.17.C65 J34 2019.

Jahnert:2019:C

[Jäh19b] Martin Jähnert. Conclusion. In Practicing the Correspondence Principle in the Old Quantum Theory: a Transformation through Implementation [Jäh19i], chapter 8, pages 255–269. CO-DEN ???? ISBN 3-030-13299-4, 3-030-13300-1 (e-book), 3-030-13301-X, 3-030-13302-8. ISSN ???? LCCN QC174.17.C65

J34 2019. URL https://link.
springer.com/chapter/10.
1007/978-3-030-13300-9_8.

Jahnert:2019:CRI

[Jäh19c] Martin Jähnert. Copenhagen reactions: The intensity problem in Copenhagen, 1924–1925. In Practicing the Correspondence Principle in the Old Quantum Theory: a Transformation through Implementation [Jäh19i], chapter 7, pages 211–253. CO-DEN ???? ISBN 3-030-13299-4, 3-030-13300-1 (e-book), 3-030-13301-X, 3-030-13302-8. ISSN ???? LCCN QC174.17.C65

J34 2019. URL https://link.
springer.com/chapter/10.
1007/978-3-030-13300-9_7.

Jahnert:2019:CPC

[Jäh19d] Martin Jähnert. The correspondence principle in Copenhagen 1913–1923: Origin, formulation and consolidation. In Practicing the Correspondence Principle in the Old Quantum Theory: a Transformation through Implementation [Jäh19i], chapter 2, pages 13–51. CO-DEN ???? ISBN 3-030-13299-4, 3-030-13300-1 (e-book), 3-030-13301-X, 3-030-13302-8. ISSN ???? LCCN QC174.17.C65

J34 2019. URL https://link.
springer.com/chapter/10.
1007/978-3-030-13300-9_2.

Jahnert:2019:CPQ

[Jäh19e] Martin Jähnert. The correspondence principle in the quantum network 1918–1926. In Practicing the Correspondence Principle in the Old Quantum Theory: a Transformation through Implementation [Jäh19i], chapter 3, pages 53–67. CO-DEN ???? ISBN 3-030-13299-4, 3-030-13300-1 (e-book), 3-030-13301-X, 3-030-13302-8. ISSN ???? LCCN QC174.17.C65

J34 2019. URL https://link.
springer.com/chapter/10.
1007/978-3-030-13300-9_3.

Jahnert:2019:FSA

[Jäh19f] Martin Jähnert. Fertilizer on a sandy acreage: Franck, Hund and the Ramsauer effect. In Practicing the Correspondence Principle in the Old Quantum Theory: a Transformation through Implementation [Jäh19i], chapter 5, pages 127–163. CO-DEN ???? ISBN 3-030-13299-4, 3-030-13300-1 (e-book), 3-030-13301-X, 3-030-13302-8. ISSN ???? LCCN QC174.17.C65

J34 2019. URL https://link.
springer.com/chapter/10.
1007/978-3-030-13300-9_5.

Jahnert:2019:FM

[Jäh19g] Martin Jähnert. Front matter. In Practicing the Correspondence Principle in the Old Quantum Theory: a Transformation through Implementation [Jäh19i], pages i–xiii. ISBN 3-030-13299-4, 3-030-13300-1 (e-book), 3-030-13301-X, 3-030-13302-8. LCCN QC174.17.C65

J34 2019. URL https://link.
springer.com/book/10.1007/
978-3-030-13300-9.

Jahnert:2019:I

[Jäh19h] Martin Jähnert. Introduction.

In Practicing the Correspondence Principle in the Old Quantum Theory: a Transformation through Implementation [Jäh19i], chapter 1, pages 1–12. CODEN ???? ISBN 3-030-13299-4, 3-030-13300-1 (e-book), 3-030-13301-X, 3-030-13302-8. ISSN ???? LCCN QC174.17.C65

J34 2019. URL https://link.
springer.com/chapter/10.
1007/978-3-030-13300-9_1.

Jahnert:2019:PCP

[Jäh19i] Martin Jähnert. Practicing the Correspondence Principle in the Old Quantum Theory: a Transformation through Implementation, volume 56 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2019. ISBN 3-030-13299-4, 3-030-13300-1 (e-book), 3-030-13301-X, 3-030-13302-8. xiii + 293 + 21 pp. LCCN QC174.17.C65 J34 2019.

Jahnert:2019:CCA

[Jäh19j] Martin Jähnert. That I cannot conceive of after the results of your dissertation: Fritz Reiche and the F -sum rule. In Practicing the Correspondence Principle in the Old Quantum Theory: a Transformation through Implementation [Jäh19i], chapter 6, pages 165–209. CODEN ???? ISBN 3-030-13299-4, 3-030-13300-1 (e-book), 3-030-13301-X, 3-030-13302-8. ISSN ???? LCCN QC174.17.C65

                                        J34 2019. URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-13300-9_6.

Jahnert:2019:UMW

[Jäh19k] Martin Jähnert. Using the magic wand: Sommerfeld, multiplet intensities and the correspondence principle. In Practicing the Correspondence Principle in the Old Quantum Theory: a Transformation through Implementation [Jäh19i], chapter 4, pages 69–125. CODEN ???? ISBN 3-030-13299-4, 3-030-13300-1 (e-book), 3-030-13301-X, 3-030-13302-8. ISSN ???? LCCN QC174.17.C65

                                        J34 2019. URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-13300-9_4.

Jakab:2000:AAM

[Jak00] Peter L. Jakab. Aerospace in adolescence: McCook Field and the beginnings of modern flight research. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 45– 66. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-011-4379-0_2.

Jesseph:2015:LEI

[Jes15] Douglas M. Jesseph. Leibniz on the elimination of infinitesimals.

In Norma B. Goethe, Philip Beeley, and David Rabouin, editors, G. W. Leibniz, Interrelations between Mathematics and Philosophy, volume 41 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 189– 205. Springer Netherlands, Dordrecht, The Netherlands, 2015. ISBN 94-017-9664-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-017-9664-4_9.

Johnson:2006:WVL

[JM06] Jeffrey Allan Johnson and Roy Macleod. The war the victors lost: The dilemmas of chemical disarmament, 1919–1926. In Roy Macleod and Jeffrey Allan Johnson, editors, Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 13, pages 221–245. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5490-4. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-5490-4_13.

Jungnickel:2017:BM

[JM17a] Christa Jungnickel and Russell McCormmach. Back matter. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 409–460. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bbm:978-3-319-49565-1/1.pdf.

Jungnickel:2017:CO

[JM17b] Christa Jungnickel and Russell McCormmach. Concluding observations. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 16, pages 397–408. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-49565-1_16.

Jungnickel:2017:CLC

[JM17c] Christa Jungnickel and Russell McCormmach. Connecting laws: Careers and theories in the 1840s. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 153–190. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-49565-1_7.

Jungnickel:2017:EPU

[JM17d] Christa Jungnickel and Russell McCormmach. Establishing physics at the universities. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 51–72. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-49565-1_2.

Jungnickel:2017:FC

[JM17e] Christa Jungnickel and Russell McCormmach. Foundations and connections. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 15, pages 349–396. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-49565-1_15.

Jungnickel:2017:FM

[JM17f] Christa Jungnickel and Russell McCormmach. Front matter. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xxxi. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-3-319-49565-1/1.pdf.

Jungnickel:2017:GPB

[JM17g] Christa Jungnickel and Russell McCormmach. German physicists before and around 1830. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 73–112. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017.

ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-49565-1_3.

Jungnickel:2017:KCW

[JM17h] Christa Jungnickel and Russell McCormmach. Kirchhoff, Clausius, Weber, and connectedness. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 205–239. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-49565-1_9.

Jungnickel:2017:MP

[JM17i] Christa Jungnickel and Russell McCormmach. Mathematicians and physicists. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 191–204. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-49565-1_8.

Jungnickel:2017:MTP

[JM17j] Christa Jungnickel and Russell McCormmach. Methods of theoretical physics. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 289–314. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-49565-1_12.

Jungnickel:2017:OPT

[JM17k] Christa Jungnickel and Russell McCormmach. Ordinary professorships for theoretical physics. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 13, pages 315–333. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-49565-1_13.

Jungnickel:2017:PRAb

[JM17l] Christa Jungnickel and Russell McCormmach. Physical research in the Annalen and in the Fortschritte. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 14, pages 335–347. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-49565-1_14.

Jungnickel:2017:PRAa

[JM17m] Christa Jungnickel and Russell McCormmach. Physical research in the Annalen and other journals around 1870. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 241–255. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-49565-1_10.

Jungnickel:2017:PRP

[JM17n] Christa Jungnickel and Russell McCormmach. Physics research in “Poggendorff’s Annalen” in the 1840s. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 137–151. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-49565-1_6.

Jungnickel:2017:PTP

[JM17o] Christa Jungnickel and Russell McCormmach. Positions in theoretical physics. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 257–288. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-49565-1_11.

Jungnickel:2017:PNP

[JM17p] Christa Jungnickel and Russell McCormmach. Promoting a new physics: Earth magnetism at Göttingen. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 113–125. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-49565-1_4.

Jungnickel:2017:RTU

[JM17q] Christa Jungnickel and Russell McCormmach. Reforms in teaching university physics: Development of the seminar and the laboratory in the 1830s and 1840s. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 127–136. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-49565-1_5.

Jungnickel:2017:SPH

[JM17r] Christa Jungnickel and Russell McCormmach. The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-49564-X (hardcover), 3-319-49565-8 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xxxi + 460 pp. LCCN QC9.G3 J86 2017.

Jungnickel:2017:TCT

[JM17s] Christa Jungnickel and Russell McCormmach. Toward a characterization of theoretical physics in Germany. In The Second Physicist: On the History of Theoretical Physics in Germany, volume 48 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–50. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2017. ISBN 3-319-49565-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-49565-1_1.

Johnson:2006:TMM

[Joh06] Jeffrey Allan Johnson. Technological mobilization and munitions production: Comparative perspectives on Germany and Austria. In Roy Macleod and Jeffrey Allan Johnson, editors, Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–20. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5490-4. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/1-4020-5490-4_1.

Jones:2005:OWS

[Jon05] Alexander Jones. “In order that we should not ourselves appear to be adjusting our estimates … to make them fit some predetermined amount”. In Jed Z. Buchwald and Allan Franklin, editors, Wrong for the Right Reasons, volume 11 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 17–39. Springer, Dordrecht, The Netherlands, 2005. ISBN 1-4020-3048-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3048-7_2.

Jones:2010:ARA

[Jon10a] Alexander Jones. Ancient rejection and adoption of Ptolemy’s frame of reference for longitudes. In Alexander Jones, editor, Ptolemy in Perspective: Use and Criticism of his Work from Antiquity to the Nineteenth Century, volume 23 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 11–44. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-2788-2. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-2788-7_2.

Jones:2010:PPU

[Jon10b] Alexander Jones, editor. Ptolemy in Perspective: Use and Criticism of his Work from Antiquity to the Nineteenth Century, volume 23 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2010. ISBN 90-481-2787-4 (hardcover), 90-481-2788-2 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xv + 229 pp. LCCN Q124.95 J66 2010.

Jones:2016:MJO

[JPS16] Alexander  Jones,  Chris-

tine Proust, and John M. Steele, editors. A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science, volume 45 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2016. ISBN 3-319-25863-X (hardcover), 3-319-25865-6 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xi + 342 pp. LCCN Q125 .M38 2016.

Kallinen:2006:LPV

[Kal06] Maija Kallinen. Lectures and practices. The variety of mathematical and mechanical teaching at the University of Uppsala in the 17th Century. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 111–125. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3975-1_8.

Kassler:2018:A

[Kas18a] Jamie C. Kassler. Appendix.

In Newton’s Sensorium: Anatomy of a Concept, volume 53 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 139–151. Springer International Publishing AG, Cham, Switzerland, 2018. ISBN 3-319-72053-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-72053-1_6.

Kassler:2018:BM

[Kas18b] Jamie C. Kassler. Back matter. In Newton’s Sensorium: Anatomy of a Concept, volume 53 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 153– 190. Springer International Publishing AG, Cham, Switzerland, 2018. ISBN 3-319-72053-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bbm:978-3-319-72053-1/1.pdf.

Kassler:2018:C

[Kas18c] Jamie C. Kassler. Conclusion. In Newton’s Sensorium: Anatomy of a Concept, volume 53 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 127–138. Springer International Publishing AG, Cham, Switzerland, 2018. ISBN 3-319-72053-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-72053-1_5.

Kassler:2018:FMa

[Kas18d] Jamie C. Kassler. Front matter. In Newton’s Sensorium: Anatomy of a Concept, volume 53 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i– xxv. Springer International Publishing AG, Cham, Switzerland, 2018. ISBN 3-319-72053-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bfm:978-3-319-72053-1/1.pdf.

Kassler:2018:FMb

[Kas18e] Jamie C. Kassler. Front matter. In Newton’s Sensorium: Anatomy of a Concept, volume 53 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 1. Springer International Publishing AG, Cham, Switzerland, 2018.

ISBN 3-319-72053-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-3-319-72053-1/1/1.
pdf.

Kassler:2018:FMc

[Kas18f] Jamie C. Kassler. Front matter. In Newton’s Sensorium: Anatomy of a Concept, volume 53 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 29. Springer International Publishing AG, Cham, Switzerland, 2018. ISBN 3-319-72053-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-3-319-72053-1/2/1.
pdf.

Kassler:2018:FMd

[Kas18g] Jamie C. Kassler. Front matter. In Newton’s Sensorium: Anatomy of a Concept, volume 53 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 59. Springer International Publishing AG, Cham, Switzerland, 2018. ISBN 3-319-72053-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-3-319-72053-1/3/1.
pdf.

Kassler:2018:FMe

[Kas18h] Jamie C. Kassler. Front matter. In Newton’s Sensorium: Anatomy of a Concept, volume 53 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 97. Springer International Publishing AG, Cham, Switzerland, 2018. ISBN 3-319-72053-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-3-319-72053-1/4/1.
pdf.

Kassler:2018:SNT

[Kas18i] Jamie C. Kassler. I The Sensorium in Newton’s texts. In Newton’s Sensorium: Anatomy of a Concept, volume 53 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 3–27. Springer International Publishing AG, Cham, Switzerland, 2018. ISBN 3-319-72053-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-72053-1_1.

Kassler:2018:IHSa

[Kas18j] Jamie C. Kassler. Ii the human Sensorium in context. In Newton’s Sensorium: Anatomy of a Concept, volume 53 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 31– 57. Springer International Publishing AG, Cham, Switzerland, 2018. ISBN 3-319-72053-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-72053-1_2.

Kassler:2018:IHSb

[Kas18k] Jamie C. Kassler. III The human Sensorium in wider context. In Newton’s Sensorium: Anatomy of a Concept, volume 53 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 61–95. Springer International Publishing AG, Cham, Switzerland, 2018. ISBN 3-319-72053-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-72053-1_3.

Kassler:2018:IGD

[Kas18l] Jamie C. Kassler. IV Generalising to the divine Sensorium. In Newton’s Sensorium: Anatomy of a Concept, volume 53 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 99– 125. Springer International Publishing AG, Cham, Switzerland, 2018. ISBN 3-319-72053-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-72053-1_4.

Kassler:2018:NSA

[Kas18m] Jamie C. Kassler. Newton’s Sensorium: Anatomy of a Concept, volume 53 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2018. ISBN 3-319-72052-X (hardcover), 3-319-72053-8 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xxv + 190 pp. LCCN Q6 .K37 2018.

Keller:2015:OOS

[Kel15] Agathe Keller. Ordering operations in square root extractions, analyzing some early medieval Sanskrit mathematical texts with the help of speech act theory. In Karine Chemla and Jacques Virbel, editors, Texts, Textual Acts and the History of Science, volume 42 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 183–218. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-16444-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-16444-1_5.

Kieffer:2014:LAA

[Kie14] Fanny Kieffer. The laboratories of art and alchemy at the Uffizi Gallery in Renaissance Florence: Some material aspects. In Sven Dupré, editor, Laboratories of Art: Alchemy and Art Technology from Antiquity to the 18th Century, volume 37 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 105–127. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05065-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-05065-2_5.

Kim:2002:BM

[Kim02a] Dong-Won Kim. Back matter.

In Leadership and Creativity: A History of the Cavendish Laboratory, 1871–1919, volume 5 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 187–225. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2002. ISBN 94-017-2055-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-94-017-2055-7/1.pdf.

Kim:2002:BCT

[Kim02b] Dong-Won Kim. The beginning of the Cavendish traditions, 1871–1879. In Leadership and Creativity: A History of the Cavendish Laboratory, 1871– 1919, volume 5 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–25. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2002. ISBN 94-017-2055-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-017-2055-7_1.

Kim:2002:ECS

[Kim02c] Dong-Won Kim. The emergence of the Cavendish School, 1895–1900. In Leadership and Creativity: A History of the Cavendish Laboratory, 1871–1919, volume 5 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 93–118. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2002. ISBN 94-017-2055-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-017-2055-7_4.

Kim:2002:EE

[Kim02d] Dong-Won Kim. The end of an era, 1914–1919. In Leadership and Creativity: A History of the Cavendish Laboratory, 1871–1919, volume 5 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 175–186. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2002. ISBN 94-017-2055-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-017-2055-7_6.

Kim:2002:FM

[Kim02e] Dong-Won Kim. Front matter. In Leadership and Creativity: A History of the Cavendish Laboratory, 1871–1919, volume 5 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xxiv. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2002. ISBN 94-017-2055-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-017-2055-7/1.pdf.

Kim:2002:JJTa

[Kim02f] Dong-Won Kim. J. J. Thomson’s first ten years at the Cavendish, 1885–1894. In Leadership and Creativity: A History of the Cavendish Laboratory, 1871–1919, volume 5 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 51–92. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2002. ISBN 94-017-2055-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-017-2055-7_3.

Kim:2002:JJTb

[Kim02g] Dong-Won Kim. J. J. Thomson’s leadership and the development of the Cavendish School, 1901–1914. In Leadership and Creativity: A History of the Cavendish Laboratory, 1871–1919, volume 5 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 119–174. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2002. ISBN 94-017-2055-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-017-2055-7_5.

Kim:2002:LCH

[Kim02h] Dong-Won Kim. Leadership and Creativity: A History of the Cavendish Laboratory, 1871–1919, volume 5 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2002. ISBN 90-481-5956-3 (hardcover), 94-017-2055-X (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xxiv + 225 pp. LCCN QC51.G72 K45 2002.

Kim:2002:RD

[Kim02i] Dong-Won Kim. Rayleigh’s directorship, 1880–1884. In Leadership and Creativity: A History of the Cavendish Laboratory, 1871– 1919, volume 5 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 26–50. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2002. ISBN 94-017-2055-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-017-2055-7_2.

Kim:2012:NTI

[Kim12] Yung Sik Kim. “Natural theology of industry” in Seventeenth-Century China?: Ideas about the role of heaven in production techniques in Song Yingxing’s Heaven’s work in opening things (tiangong kaiwu). In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 197–214.

Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2627-7_12.

Kinder:2006:TVI

[Kin06] Sebastian Kinder. Transforming a village into an industrial town: The Royal Prussian Powder Plant in Kirchmöser (Brandenburg). In Roy Macleod and Jeffrey Allan Johnson, editors, Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 47–60. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5490-4. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-5490-4_4.

Kisacky:2014:BRC

[Kis14] Jeanne Kisacky. Breathing room:

Calculating an architecture of air. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400– 1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 247–279. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014.

ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-05998-3_11.

Klaiber:2014:AME

[Kla14] Susan Klaiber. Architecture and mathematics in early modern religious orders. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 137–164. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-05998-3_6.

Klein:2007:ACE

[Kle07]  Ursula Klein.  Apothecary-

chemists in Eighteenth-Century Germany. In Lawrence M. Principe, editor, New Narratives in Eighteenth-Century Chemistry: Contributions from the First Francis Bacon Workshop, 21–23 April 2005, California Institute of Technology, Pasadena, California, volume 18 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 97–137. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2007. ISBN 1-4020-6278-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-6278-0_6.

Kleeberg:2010:VBN

[Kle10]  Bernhard Kleeberg.  Vestiges

of the book of nature: Religious experience and hermeneutic practices in Protestant German theology, ca. 1900. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 37–59. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-3540-0_3.

Kleve:2012:PPL

[Kle12] Knut Kleve. The puzzle picture of Lucretius: A thriller from Herculaneum. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 63–78. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2627-7_6.

Knight:2017:PCW

[Kni17] David Knight. Poetry, chemistry, and wisdom. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 81–94. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-58436-2_5.

Knoeff:2007:PCA

[Kno07] Rina Knoeff. Practicing chemistry “after the hippocratical manner”: Hippocrates and the importance of chemistry for Boerhaave’s medicine. In Lawrence M. Principe, editor, New Narratives in Eighteenth-Century Chemistry: Contributions from the First Francis Bacon Workshop, 21–23 April 2005, California Institute of Technology, Pasadena, California, volume 18 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 63–76. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2007. ISBN 1-4020-6278-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-1-4020-6278-0_4.

Knobloch:2015:AET

[Kno15] Eberhard Knobloch. Analyticité, équipollence et théorie des courbes chez Leibniz. (French) [Analyticity, equipollence and curve theory in Leibniz]. In Norma B. Goethe, Philip Beeley, and David Rabouin, editors, G. W. Leibniz, Interrelations between Mathematics and Philosophy, volume 41 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 89–110. Springer Netherlands, Dordrecht, The Netherlands, 2015. ISBN 94-017-9664-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-017-9664-4_5.

Langlinay:2006:KW

[Lan06] Erik Langlinay. Kuhlmann at war, 1914–1924. In Roy Macleod and Jeffrey Allan Johnson, editors, Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 145–166. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5490-4. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-5490-4_9.

Langins:2017:PWW

[Lan17] Janis Langins. Poetry in war and war in nature. From Vauban to Naturphilosophie to Clausewitz. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 117–134. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-58436-2_7.

Lavelle:2015:AIC

[Lav15] Peter Lavelle. Agricultural improvement at China’s first agricultural experiment stations. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 16, pages 323–344. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-12185-7_16.

Leary:2010:IEV

[Lea10] David E. Leary. Instead of Erklären and Verstehen: William James on human understanding. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 121–140.

Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-3540-0_7.

Lee:2015:MCT

[Lee15] Victoria Lee. Mold cultures:

Traditional industry and microbial studies in Early Twentieth-Century Japan. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 231–252. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-12185-7_12.

Leitao:2003:JMP

[Lei03] Henrique Leitão. Jesuit mathematical practice in Portugal, 1540–1759. In Mordechai Feingold, editor, The New Science and Jesuit Science: Seventeenth Century Perspectives, volume 6 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 229–247. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2003.

ISBN 94-017-0361-2. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-017-0361-1_5.

Levere:2007:DTB

[Lev07] Trevor H. Levere. Dr. Thomas Beddoes (1760–1808): Chemistry, medicine, and books in the French and chemical revolutions. In Lawrence M. Principe, editor, New Narratives in Eighteenth-Century Chemistry: Contributions from the First Francis Bacon Workshop, 21–23 April 2005, California Institute of Technology, Pasadena, California, volume 18 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 157–176. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2007. ISBN 1-4020-6278-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-1-4020-6278-0_8.

Levey:2015:CII

[Lev15] Samuel Levey. Comparability of infinities and infinite multitude in Galileo and Leibniz. In Norma B. Goethe, Philip Beeley, and David Rabouin, editors, G. W. Leibniz, Interrelations between Mathematics and Philosophy, volume 41 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 157–187. Springer Netherlands, Dordrecht, The Netherlands, 2015. ISBN 94-017-9664-5.

ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-017-9664-4_8.

Loskutova:2015:RAE

[LF15] Marina V. Loskutova and Anastasia A. Fedotova. The rise of applied entomology in the Russian Empire: Governmental, public, and academic responses to insect pest outbreaks from 1840 to 1894. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 139–162. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-12185-7_8.

Loettgers:2003:ECB

[Loe03] Andrea Loettgers. Exploring contents and boundaries of experimental practice in laboratory notebooks: Samuel Pierpont Langley and the mapping of the infrared region of the solar spectrum. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 159–182. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/0-306-48152-9_8.

Maclean:2017:MCA

[Mac17] Ian Maclean. A medical collection anatomized: The Catalogus bibliothecae Hieremiae Martii (1572). In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 129–151. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-56514-9_8.

Martelli:2014:AAD

[Mar14] Matteo Martelli. The alchemical art of dyeing: The fourfold division of alchemy and the Enochian tradition. In Sven Dupré, editor, Laboratories of Art: Alchemy and Art Technology from Antiquity to the 18th Century, volume 37 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–22. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05065-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-05065-2_1.

Matz:2015:NSP

[Mat15] Brendan Matz. Nutrition science and the practice of animal feeding in Germany, 1850–1880. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 163–181. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-12185-7_9.

Mauskopf:2007:RLS

[Mau07] Seymour Mauskopf. Reflections:

“a likely story”. In Lawrence M. Principe, editor, New Narratives in Eighteenth-Century Chemistry: Contributions from the First Francis Bacon Workshop, 21–23 April 2005, California Institute of Technology, Pasadena, California, volume 18 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 177–193. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2007. ISBN 1-4020-6278-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-1-4020-6278-0_9.

Mauskopf:2012:CHS

[Mau12] Seymour H. Mauskopf. A career in the history of science as a student of Charles Gillispie. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 25–35. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-2627-7_3.

McCormmach:2012:BM

[McC12a] Russell McCormmach. Back matter. In Weighing the World: the Reverend John Michell of Thornhill, volume 28 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 459–488. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2022-X. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bbm:978-94-007-2022-0/1.pdf.

McCormmach:2012:C

[McC12b] Russell McCormmach. Cambridge. In Weighing the World: the Reverend John Michell of Thornhill, volume 28 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 21–52. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2022-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2022-0_2.

McCormmach:2012:ER

[McC12c] Russell McCormmach. Early researches. In Weighing the World: the Reverend John Michell of Thornhill, volume 28 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 53–107. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2022-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2022-0_3.

McCormmach:2012:FMa

[McC12d] Russell McCormmach. Front matter. In Weighing the World: the Reverend John Michell of Thornhill, volume 28 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xvii. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2022-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bfm:978-94-007-2022-0/1.pdf.

McCormmach:2012:FMb

[McC12e] Russell McCormmach. Front matter. In Weighing the World: the Reverend John Michell of Thornhill, volume 28 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 1. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2022-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bfm:978-94-007-2022-0/1/1.
                                        pdf.

McCormmach:2012:FMc

[McC12f] Russell McCormmach. Front matter. In Weighing the World: the Reverend John Michell of Thornhill, volume 28 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 305. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2022-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bfm:978-94-007-2022-0/2/1.
                                        pdf.

McCormmach:2012:H

[McC12g] Russell McCormmach. Home. In Weighing the World: the Reverend John Michell of Thornhill, volume 28 of Archimedes:

New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 3–19. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2022-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2022-0_1.

McCormmach:2012:LR

[McC12h] Russell McCormmach. Late researches. In Weighing the World: the Reverend John Michell of Thornhill, volume 28 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 203–304. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2022-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2022-0_6.

McCormmach:2012:L

[McC12i] Russell McCormmach. Letters.

In Weighing the World: the Reverend John Michell of Thornhill, volume 28 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 307–457. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2022-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2022-0_7.

McCormmach:2012:Tb

[McC12j] Russell McCormmach. Thornhill. In Weighing the World: the Reverend John Michell of Thornhill, volume 28 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 159–201. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2022-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-2022-0_5.

McCormmach:2012:Ta

[McC12k] Russell McCormmach. Transitions. In Weighing the World: the Reverend John Michell of Thornhill, volume 28 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 109–157. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2022-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-2022-0_4.

McCormmach:2012:WWR

[McC12l] Russell McCormmach. Weighing the World: the Reverend John Michell of Thornhill, volume 28 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2012. ISBN 94-007-2021-1 (hardcover), 94-007-2022-X (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xvii + 488 pp. LCCN T14.7 .A73 vol. 28; Q143.M52 M33 2012.

McCormmach:2014:AP

[McC14a] Russell McCormmach. Aspects of personality. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 97–105. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-02438-7_5.

McCormmach:2014:A

[McC14b] Russell McCormmach. Autism. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 135–148. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-02438-7_8.

McCormmach:2014:AB

[McC14c] Russell McCormmach. Autism, background. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 16, pages 289–292. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-02438-7_16.

McCormmach:2014:CGF

[McC14d] Russell McCormmach. Cavendish and Grey family trees. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 251–257. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-02438-7_12.

McCormmach:2014:CLH

[McC14e] Russell McCormmach. Cavendishes, Lowthers, and Holker Hall. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 15, pages 275–287. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-02438-7_15.

McCormmach:2014:CHC

[McC14f] Russell Mc-Cormmach. Cavendish’s houses: Clapham Common and Bedford Square. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 14, pages 263–274. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-02438-7_14.

McCormmach:2014:C

[McC14g] Russell McCormmach. Conclusion. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 149–164. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-02438-7_9.

McCormmach:2014:E

[McC14h] Russell McCormmach. Eccentricity. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 115–134. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-02438-7_7.

McCormmach:2014:EL

[McC14i] Russell McCormmach. Everyday life. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 25– 57. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-02438-7_3.

McCormmach:2014:FMa

[McC14j] Russell McCormmach. Front matter. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xvii. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-3-319-02438-7/1.pdf.

McCormmach:2014:FMb

[McC14k] Russell McCormmach. Front matter. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 1. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/content/
                                        pdf/bfm:978-3-319-02438-7/1/
                                        1.pdf.

McCormmach:2014:FMc

[McC14l] Russell McCormmach. Front matter. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 165. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/content/
                                        pdf/bfm:978-3-319-02438-7/2/
                                        1.pdf.

McCormmach:2014:FMd

[McC14m] Russell McCormmach. Front matter. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 229. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/content/
pdf/bfm:978-3-319-02438-7/3/
1.pdf.

McCormmach:2014:J

[McC14n] Russell McCormmach. Journal of 1785. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 231–250. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-02438-7_11.

McCormmach:2014:LBB

[McC14o] Russell McCormmach. Letters between brothers: Henry and Frederick Cavendish. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 13, pages 259–261. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-02438-7_13.

McCormmach:2014:LNP

[McC14p] Russell McCormmach. Life of natural philosophy. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 59–96. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-02438-7_4.

McCormmach:2014:NAS

[McC14q] Russell McCormmach. Normality, abnormality, and scientists. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 17–24. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-02438-7_2.

McCormmach:2014:P

[McC14r] Russell McCormmach. The person. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 3–16. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-02438-7_1.

McCormmach:2014:PHC

[McC14s] Russell McCormmach. The Personality of Henry Cavendish — A Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2014. ISBN 3-319-02437-X (hardcover), 3-319-02438-8 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xvii + 310 pp. LCCN QD22.C4 M33 2014.

McCormmach:2014:T

[McC14t] Russell McCormmach. The traveler. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 167–228. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-02438-7_10.

McCormmach:2014:VC

[McC14u] Russell McCormmach. Views of Cavendish. In The Personality of Henry Cavendish — a Great Scientist with Extraordinary Peculiarities, volume 36 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 107–114. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-02438-8. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-02438-7_6.

McOuat:2017:CAI

[McO17] Gordon McOuat. “These can not all have an interest for England”: Symmetry, beauty and the trouble with Romanticism in Britain. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 163– 180. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-58436-2_10.

McVaugh:2017:MC

[McV17] Michael McVaugh. In a Montpellier classroom. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 57– 76. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-56514-9_4.

Melville:2016:ANR

[Mel16] Duncan J. Melville. After Neugebauer: Recent developments in Mesopotamian mathematics. In Alexander Jones, Christine Proust, and John M. Steele, editors, A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science, volume 45 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 237–263. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-25865-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-25865-2_8.

Meli:2017:GBI

[Mel17] Domenico Bertoloni Meli. Gerardus Blasius and the Illustrated Amsterdam Observationes from Nicolaas Tulp to Frederik Ruysch. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 255–297. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-56514-9_12.

Meskens:2013:AAB

[Mes13a] Ad Meskens. The Antwerp arithmetic books. In Practical Mathematics in a Commercial Metropolis: Mathematical Life in Late 16th Century Antwerp, volume 31 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 57–95. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5721-2. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.

springer.com/chapter/10.
1007/978-94-007-5721-9_5.

Meskens:2013:ATH

[Mes13b] Ad Meskens. The arithmetic teacher and his school. In Practical Mathematics in a Commercial Metropolis: Mathematical Life in Late 16th Century Antwerp, volume 31 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 35–56. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5721-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-5721-9_4.

Meskens:2013:AN

[Mes13c] Ad Meskens. The art of navigation. In Practical Mathematics in a Commercial Metropolis: Mathematical Life in Late 16th Century Antwerp, volume 31 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 139–160. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5721-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-5721-9_8.

Meskens:2013:BM

[Mes13d] Ad Meskens. Back matter. In Practical Mathematics in a Commercial Metropolis: Mathematical Life in Late 16th Century Antwerp, volume 31 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 213–258. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5721-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-94-007-5721-9/1.pdf.

Meskens:2013:BF

[Mes13e] Ad Meskens. Ballistics and fortifications. In Practical Mathematics in a Commercial Metropolis: Mathematical Life in Late 16th Century Antwerp, volume 31 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 197–210. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5721-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-007-5721-9_11.

Meskens:2013:CF

[Mes13f] Ad Meskens. The Coignet family. In Practical Mathematics in a Commercial Metropolis: Mathematical Life in Late 16th Century Antwerp, volume 31 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 9–23. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013.

ISBN 94-007-5721-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-5721-9_2.

Meskens:2013:C

[Mes13g] Ad Meskens. Conclusion. In Practical Mathematics in a Commercial Metropolis: Mathematical Life in Late 16th Century Antwerp, volume 31 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 211–212. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5721-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-007-5721-9_12.

Meskens:2013:FM

[Mes13h] Ad Meskens. Front matter. In Practical Mathematics in a Commercial Metropolis: Mathematical Life in Late 16th Century Antwerp, volume 31 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xv. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5721-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-007-5721-9/1.pdf.

Meskens:2013:IM

[Mes13i] Ad Meskens. Instrument makers. In Practical Mathematics in a Commercial Metropolis: Mathematical Life in Late 16th Century Antwerp, volume 31 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 113–137. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5721-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-5721-9_7.

Meskens:2013:I

[Mes13j] Ad Meskens. Introduction. In Practical Mathematics in a Commercial Metropolis: Mathematical Life in Late 16th Century Antwerp, volume 31 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–8. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5721-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-5721-9_1.

Meskens:2013:MW

[Mes13k] Ad Meskens. Mapping the world. In Practical Mathematics in a Commercial Metropolis: Mathematical Life in Late 16th Century Antwerp, volume 31 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 161–177. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5721-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-5721-9_9.

Meskens:2013:PHN

[Mes13l] Ad Meskens. Peter Heyns and the nymphs of the laurel tree. In Practical Mathematics in a Commercial Metropolis: Mathematical Life in Late 16th Century Antwerp, volume 31 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 25–34. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5721-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-007-5721-9_3.

Meskens:2013:PMC

[Mes13m] Ad Meskens. Practical Mathematics in a Commercial Metropolis: Mathematical Life in Late 16th Century Antwerp, volume 31 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2013. ISBN 94-007-5720-4 (hardcover), 94-007-5721-2 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xv + 258 pp. LCCN QA24 .M47 2013. Drawings by Paul Tytgat.

Meskens:2013:S

[Mes13n] Ad Meskens. Stargazing. In Practical Mathematics in a Commercial Metropolis: Mathematical Life in Late 16th Century Antwerp, volume 31 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 179–196. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5721-2. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-007-5721-9_10.

Meskens:2013:WG

[Mes13o] Ad Meskens. Wine gauging. In Practical Mathematics in a Commercial Metropolis: Mathematical Life in Late 16th Century Antwerp, volume 31 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 97–112. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2013. ISBN 94-007-5721-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-5721-9_6.

Miller:2017:MLM

[Mil17] David Philip Miller. “Men of letters” and “men of press copies”: The cultures of James Watt’s copying machine. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 65– 79. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-58436-2_4.

Mittenhuber:2010:TTM

[Mit10] Florian Mittenhuber. The tradition of texts and maps in Ptolemy’s geography. In Alexander Jones, editor, Ptolemy in Perspective: Use and Criticism of his Work from Antiquity to the Nineteenth Century, volume 23 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 95–119. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-2788-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-2788-7_4.

Macleod:2006:FFC

[MJ06] Roy Macleod and Jeffrey Allan Johnson, editors. Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5489-0 (hardcover), 1-4020-5490-4 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xix + 277 pp. LCCN TP20 .F765 2006.

Manning:2017:I

[MK17a] Gideon Manning and Cynthia Klestinec. Introduction. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–14. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-56514-9_1.

Manning:2017:PPP

[MK17b] Gideon Manning and Cynthia Klestinec, editors. Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56513-3 (hardcover), 3-319-56514-1 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xliv + 279 pp. LCCN R131 .P765 2017.

Maienschein:2012:CGD

[ML12] Jane Maienschein and Manfred D. Laubichler. Charles Gillispie in the Digital Age. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 37–45. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2627-7_4.

Mazliak:2009:BM

[MT09a] Laurent Mazliak and Rossana Tazzioli. Back matter. In Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 189–194. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2009. ISBN 90-481-2740-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-90-481-2740-5/1.pdf.

Mazliak:2009:CP

[MT09b] Laurent Mazliak and Rossana Tazzioli. The case of pérès. In Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 155–164. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2009. ISBN 90-481-2740-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-2740-5_7.

Mazliak:2009:CRB

[MT09c] Laurent Mazliak and Rossana Tazzioli. Cultural relations between Italy and France. In Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 181–188. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2009. ISBN 90-481-2740-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-2740-5_10.

Mazliak:2009:FP

[MT09d] Laurent Mazliak and Rossana Tazzioli. French propaganda. In Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 171–179. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2009. ISBN 90-481-2740-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-2740-5_9.

Mazliak:2009:FMa

[MT09e] Laurent Mazliak and Rossana Tazzioli. Front matter. In Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–ix. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2009. ISBN 90-481-2740-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-90-481-2740-5/1.pdf.

Mazliak:2009:FMb

[MT09f] Laurent Mazliak and Rossana Tazzioli. Front matter. In Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 1. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2009. ISBN 90-481-2740-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-90-481-2740-5/1/1.
pdf.

Mazliak:2009:FMc

[MT09g] Laurent Mazliak and Rossana Tazzioli. Front matter. In Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 33. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2009. ISBN 90-481-2740-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bfm:978-90-481-2740-5/2/1.
                                        pdf.

Mazliak:2009:FMd

[MT09h] Laurent Mazliak and Rossana Tazzioli. Front matter. In Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 153. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2009. ISBN 90-481-2740-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bfm:978-90-481-2740-5/3/1.
                                        pdf.

Mazliak:2009:IFE

[MT09i] Laurent Mazliak and Rossana Tazzioli. Italy and France at the eve of WW1. In Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 13–16. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2009. ISBN 90-481-2740-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-2740-5_3.

Mazliak:2009:MWV

[MT09j] Laurent Mazliak and Rossana Tazzioli. Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 90-481-2739-4 (hardcover), 90-481-2740-8 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). ix + 194 pp. LCCN QA29.V65 M395 2009.

Mazliak:2009:PAJ

[MT09k] Laurent Mazliak and Rossana Tazzioli. Persuading the American Jewish community. In Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 165–169. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2009. ISBN 90-481-2740-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-2740-5_8.

Mazliak:2009:P

[MT09l] Laurent Mazliak and Rossana Tazzioli. The protagonists. In Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 9–12. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2009. ISBN 90-481-2740-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-2740-5_2.

Mazliak:2009:VVH

[MT09m] Laurent Mazliak and Rossana Tazzioli. Vito Volterra and his institutional activity. In Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 17–19. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2009. ISBN 90-481-2740-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-2740-5_4.

Mazliak:2009:VII

[MT09n] Laurent Mazliak and Rossana Tazzioli. Volterra and Italian interventionism between 1914 and 1915. In Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 21–32. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2009. ISBN 90-481-2740-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-2740-5_5.

Mazliak:2009:VEB

[MT09o] Laurent Mazliak and Rossana Tazzioli. Volterra’s exchanges with Borel, Hadamard and Picard in World War One. In Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 35–152. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2009. ISBN 90-481-2740-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-2740-5_6.

Mazliak:2009:WBA

[MT09p] Laurent Mazliak and Rossana Tazzioli. What is this book about? In Mathematicians at war: Volterra and his French colleagues in World War I, volume 22 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 3–7. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2009. ISBN 90-481-2740-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-2740-5_1.

Muller:2010:EHH

[Mül10] Philipp Müller. Explaining history. Hippolyte Taine’s philosophy of historical science. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 81–99. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-3540-0_5.

Navarro-Brotons:2006:CAS

[NB06] Victor Navarro-Brotóns. The cultivation of astronomy in Spanish universities in the latter half of the 16th Century. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 83–98. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3975-1_6.

Neven:2014:TAA

[Nev14] Sylvie Neven. Transmission of alchemical and artistic knowledge in German mediaeval and premodern recipe books. In Sven Dupré, editor, Laboratories of Art: Alchemy and Art Technology from Antiquity to the 18th Century, volume 37 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 23–51. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05065-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-05065-2_2.

Nieto-Galan:1999:ISM

[NG99] Agustí Nieto-Galan. The images of science in modern spain. In Kostas Gavroglu, editor, The Sciences in the European Periphery During the Enlightenment, volume 2 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 73–94. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 94-011-4770-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-011-4770-5_3.

Nieto-Galan:2006:SEC

[NGRR06] Agustí Nieto-Galan and Antoni Roca-Rosell. Scientific education and the crisis of the university in 18th Century Barcelona. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 18, pages 273–288. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-3975-1_18.

Niazi:2014:BM

[Nia14a] Kaveh Niazi. Back matter. In Qut.b al-Dīn Shīrāzī and the Configuration of the Heavens: A Comparison of Texts and Models, volume 35 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 155–189. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-6999-7. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bbm:978-94-007-6999-1/1.pdf.

Niazi:2014:CR

[Nia14b] Kaveh Niazi. Concluding remarks. In Qut.b al-Dīn Shīrāzī and the Configuration of the Heavens: A Comparison of Texts and Models, volume 35 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 145–153. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-6999-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-94-007-6999-1_6.

Niazi:2014:FM

[Nia14c] Kaveh Niazi. Front matter. In Qut.b al-Dīn Shīrāzī and the Configuration of the Heavens: A Comparison of Texts and Models, volume 35 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xiii. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-6999-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-007-6999-1/1.pdf.

Niazi:2014:MI

[Nia14d] Kaveh Niazi. The Mongols in Iran. In Qut.b al-Dīn Shīrāzī and the Configuration of the Heavens: A Comparison of Texts and Models, volume 35 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 21–60. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-6999-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-007-6999-1_2.

Niazi:2014:PVA

[Nia14e] Kaveh Niazi. Persian vs. Arabic:

Language as determinant of content in shīrāzī’s works on Hay’a. In Qut.b al-Dīn Shīrāzī and the Configuration of the Heavens: A Comparison of Texts and Models, volume 35 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 123–143. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-6999-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-94-007-6999-1_5.

Niazi:2014:PAS

[Nia14f] Kaveh Niazi. The principal astronomical sources. In Qut.b al-Dīn Shīrāzī and the Configuration of the Heavens: A Comparison of Texts and Models, volume 35 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 85–122. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-6999-7. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-6999-1_4.

Niazi:2014:PBS

[Nia14g] Kaveh Niazi. Purpose and background of study. In Qut.b al-Dīn Shīrāzī and the Configuration of the Heavens: A Comparison of Texts and Models, volume 35 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–19. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-6999-7. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-6999-1_1.

Niazi:2014:QAD

[Nia14h] Kaveh Niazi. Qut.b al-Dīn Shīrāzī and the Configuration of the Heavens: A Comparison of Texts and Models, volume 35 of Archimedes:

New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2014. ISBN 94-007-6998-9 (hardcover), 94-007-6999-7 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xiii + 189 pp. LCCN Q174.8 .N53 2014.

Niazi:2014:SL

[Nia14i] Kaveh Niazi. Shīrāzī’s life. In Qut.b al-Dīn Shīrāzī and the Configuration of the Heavens: A Comparison of Texts and Models, volume 35 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 61–84. Springer Science+Business Media Dordrecht, Dordrecht, The Netherlands, 2014. ISBN 94-007-6999-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-6999-1_3.

Nickelsen:2006:BM

[Nic06a] Kärin Nickelsen. Back matter. In Draughtsmen, Botanists and Nature: The Construction of Eighteenth-Century Botanical Illustrations, volume 15 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 273–295. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4820-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bbm:978-1-4020-4820-3/1.pdf.

Nickelsen:2006:CBIb

[Nic06b] Kärin Nickelsen. The construction of botanical illustrations. In Draughtsmen, Botanists and Nature: The Construction of Eighteenth-Century Botanical Illustrations, volume 15 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 229–264. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4820-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-1-4020-4820-3_7.

Nickelsen:2006:CBIa

[Nic06c] Kärin Nickelsen. The content of botanical illustrations. In Draughtsmen, Botanists and Nature: The Construction of Eighteenth-Century Botanical Illustrations, volume 15 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 71–105. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4820-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-1-4020-4820-3_3.

Nickelsen:2006:DBN

[Nic06d] Kärin Nickelsen. Draughtsmen, Botanists and Nature: the Construction of Eighteenth-century Botanical Illustrations, volume 15 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4819-X (hardcover), 1-4020-4820-3 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). ix + 295 pp. LCCN QK98.15 .N53

2006. URL http://www.loc.gov/
catdir/enhancements/fy0824/
2007464743-d.html;
http://www.loc.gov/catdir/
enhancements/fy0824/
2007464743-t.html.

Nickelsen:2006:FM

[Nic06e] Kärin Nickelsen. Front matter. In Draughtsmen, Botanists and Nature: The Construction of Eighteenth-Century Botanical Illustrations, volume 15 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–ix. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4820-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-1-4020-4820-3/1.pdf.

Nickelsen:2006:I

[Nic06f] Kärin Nickelsen. Introduction.

In Draughtsmen, Botanists and Nature: The Construction of Eighteenth-Century Botanical Illustrations, volume 15 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–18. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4820-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-1-4020-4820-3_1.

Nickelsen:2006:LT

[Nic06g] Kärin Nickelsen. Links with tradition. In Draughtsmen, Botanists and Nature: The Construction of Eighteenth-Century Botanical Illustrations, volume 15 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 185–228. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4820-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-1-4020-4820-3_6.

Nickelsen:2006:MBI

[Nic06h] Kärin Nickelsen. The making of botanical illustrations. In Draughtsmen, Botanists and Nature: The Construction of Eighteenth-Century Botanical Illustrations, volume 15 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 19–70. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4820-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-1-4020-4820-3_2.

Nickelsen:2006:MM

[Nic06i] Kärin Nickelsen. Methods and materials. In Draughtsmen, Botanists and Nature: The Construction of Eighteenth-Century Botanical Illustrations, volume 15 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 265–271. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4820-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-1-4020-4820-3_8.

Nickelsen:2006:RBI

[Nic06j] Kärin Nickelsen. The role of botanical illustrations. In Draughtsmen, Botanists and Nature: The Construction of Eighteenth-Century Botanical Illustrations, volume 15 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 107–148. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4820-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-4820-3_4.

Nickelsen:2006:VL

[Nic06k] Kärin Nickelsen. Visual language. In Draughtsmen, Botanists and Nature: The Construction of Eighteenth-Century Botanical Illustrations, volume 15 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 149–184. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4820-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-1-4020-4820-3_5.

Nickles:2006:HAC

[Nic06l] Thomas Nickles. Heuristic appraisal: Context of discovery or justification? In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 159–182. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-4251-5_10.

Newman:2003:CLN

[NP03]  William  R.  New-

man and Lawrence M. Principe. The chymical laboratory notebooks of George Starkey. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 25– 41. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/0-306-48152-9_2.

Newman:2005:ACS

[NP05]  William  R.  New-

man and Lawrence M. Principe. Alchemy and the changing significance of analysis. In Jed Z. Buchwald and Allan Franklin, editors, Wrong for the Right Reasons, volume 11 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 73–89. Springer, Dordrecht, The Netherlands, 2005. ISBN 1-4020-3048-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3048-7_4.

Nutton:2017:JCH

[Nut17] Vivian Nutton. John Caius, historian. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 97– 111. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-56514-9_6.

Oldroyd:2012:GHR

[Old12] David Oldroyd. The genesis of historical research on the history of geology, with thoughts about Kirwan, de Luc, and Whiggery. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 167–177. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2627-7_10.

Olesko:1997:PTC

[Ole97] Kathryn M. Olesko. Precision, tolerance, and consensus: Local cultures in German and British resistance standards. In Jed Z. Buchwald, editor, Scientific Credibility and Technical Standards in 19th and early 20th century Germany and Britain, volume 1 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 117–156. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 1997. ISBN 94-009-1784-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-009-1784-2_4.

Oghina-Pavie:2015:RPB

[OP15] Cristiana Oghina-Pavie. Rose and pear breeding in Nineteenth-Century France: The practice and science of diversity. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 53– 72. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-12185-7_4.

Ossendrijver:2016:TBM

[Oss16] Mathieu Ossendrijver. Translating Babylonian mathematical astronomy: Neugebauer and beyond. In Alexander Jones, Christine Proust, and John M. Steele, editors, A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science, volume 45 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 333–342. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-25865-6. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-25865-2_11.

Palmerino:2003:TJR

[Pal03] Carla Rita Palmerino. Two Jesuit responses to Galileo’s science of motion: Honore Fabri and Pierre le Cazre. In Mordechai Feingold, editor, The New Science and Jesuit Science: Seventeenth Century Perspectives, volume 6 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 187–227. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2003. ISBN 94-017-0361-2. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-017-0361-1_4.

Pancaldi:2006:WCI

[Pan06] Giuliano Pancaldi. Wartime chemistry in Italy: Industry, the military, and the professors. In Roy Macleod and Jeffrey Allan Johnson, editors, Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 61–74. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5490-4. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/1-4020-5490-4_5.

Pantalony:2009:ASR

[Pan09a] David Pantalony. Altered Sensations: Rudolph Koenig’s Acoustical Workshop in Nineteenth-Century Paris, volume 24 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 90-481-2815-3 (hardcover), 90-481-2816-1 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xxxvi + 372 pp. LCCN ML3805 .P368 2009.

Pantalony:2009:BM

[Pan09b] David Pantalony. Back matter.

In Altered Sensations: Rudolph Koenig’s Acoustical Workshop in Nineteenth-Century Paris, volume 24 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 167–372. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 90-481-2816-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bbm:978-90-481-2816-7/1.pdf.

Pantalony:2009:CR

[Pan09c] David Pantalony. Constructing a reputation, 1866–1879. In Altered Sensations: Rudolph Koenig’s Acoustical Workshop in Nineteenth-Century Paris, volume 24 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 83–108. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 90-481-2816-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-90-481-2816-7_5.

Pantalony:2009:ENA

[Pan09d] David Pantalony. Expanding the North American market, 1871– 1882. In Altered Sensations: Rudolph Koenig’s Acoustical Workshop in Nineteenth-Century Paris, volume 24 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 109– 132. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 90-481-2816-1. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-2816-7_6.

Pantalony:2009:FS

[Pan09e] David Pantalony. The Faraday of sound. In Altered Sensations: Rudolph Koenig’s Acoustical Workshop in Nineteenth-Century Paris, volume 24 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 133–165. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 90-481-2816-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-90-481-2816-7_7.

Pantalony:2009:FM

[Pan09f] David Pantalony. Front matter. In Altered Sensations: Rudolph Koenig’s Acoustical Workshop in Nineteenth-Century Paris, volume 24 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xxxvi. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 90-481-2816-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-90-481-2816-7/1.pdf.

Pantalony:2009:HHS

[Pan09g] David Pantalony. Hermann von Helmholtz and the sensations of tone. In Altered Sensations: Rudolph Koenig’s Acoustical Workshop in Nineteenth-Century Paris, volume 24 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 19–36. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 90-481-2816-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-90-481-2816-7_2.

Pantalony:2009:MI

[Pan09h] David Pantalony. The market and its influences. In Altered Sensations: Rudolph Koenig’s Acoustical Workshop in Nineteenth-Century Paris, volume 24 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 65–81. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 90-481-2816-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-90-481-2816-7_4.

Pantalony:2009:T

[Pan09i] David Pantalony. Training. In Altered Sensations: Rudolph Koenig’s Acoustical Workshop in Nineteenth-Century Paris, volume 24 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–17. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 90-481-2816-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-90-481-2816-7_1.

Pantalony:2009:TW

[Pan09j] David Pantalony. Transformations in the workshop. In Altered Sensations: Rudolph Koenig’s Acoustical Workshop in Nineteenth-Century Paris, volume 24 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 37–63. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 90-481-2816-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-90-481-2816-7_3.

Pantalony:2017:CDR

[Pan17] David Pantalony. Collectors, displays and replicas in context: What we can learn from provenance research in science museums. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 14, pages 255–275. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7.

ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-58436-2_14.

Parnes:2003:ACT

[Par03] Ohad Parnes. From agents to cells: Theodor Schwann’s research notes of the years 1835–1838. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 119– 140. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/0-306-48152-9_6.

Peckhaus:2006:PDB

[Pec06] Volker Peckhaus. Psychologism and the distinction between discovery and justification. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 99–116. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-4251-5_7.

Pepe:2006:UAS

[Pep06] Luigi Pepe. Universities, academies, and sciences in Italy in the Modern Age. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 141–151. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-3975-1_10.

Peset:2006:ERS

[Pes06] José Luis Peset. Enlightenment and renovation in the Spanish university. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 15, pages 231–239. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-3975-1_15.

Phillips:2010:EDC

[Phi10] Denise Phillips. Epistemological distinctions and cultural politics: Educational reform and the Naturwissenschaft/ Geisteswissenschaft distinction in Nineteenth-Century Germany. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 15–35. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-3540-0_2.

Phillips:2015:PPA

[Phi15] Denise Phillips. Plants and places:

Agricultural knowledge and plant geography in Germany, 1750– 1810. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 9–26. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-12185-7_2.

Pinero:2006:FMV

[Piñ06] José M. López Piñero. The Faculty of Medicine of Valencia: Its position in Renaissance Europe. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 65–82. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-3975-1_5.

Pincock:2010:AUE

[Pin10] Christopher Pincock. Accounting for the unity of experience in Dilthey, Rickert, Bradley and Ward. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 187–206. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-3540-0_10.

Phillips:2015:I

[PK15a] Denise Phillips and Sharon Kingsland. Introduction. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–8.

Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-12185-7_1.

Phillips:2015:NPH

[PK15b] Denise Phillips and Sharon Kingsland, editors. New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2015. ISBN 3-319-12184-7 (hardcover), 3-319-12185-5 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). vii + 509 pp. LCCN QH307.2 .N49 2015.

Potthast:2006:CJD

[Pot06] Thomas Potthast. Contexts of justifying and discovering the nature of ecosystems: From concepts to objects and vice versa. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 197–214. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-4251-5_12.

Powers:2007:CPH

[Pow07] John C. Powers. Chemistry without principles: Herman Boerhaave on instruments and elements. In Lawrence M. Principe, editor, New Narratives in Eighteenth-Century Chemistry: Contributions from the First Francis Bacon Workshop, 21–23 April 2005, California Institute of Technology, Pasadena, California, volume 18 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 45–61. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2007. ISBN 1-4020-6278-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-1-4020-6278-0_3.

Principe:2007:NNE

[Pri07a] Lawrence M. Principe, editor. New Narratives in Eighteenth-Century Chemistry: Contributions from the First Francis Bacon Workshop, 21–23 April 2005, California Institute of Technology, Pasadena, California, volume 18 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2007. ISBN 1-4020-6273-7 (hardcover), 1-4020-6278-8 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xi + 200 pp. LCCN QD14 .N49 2007.

Principe:2007:RNN

[Pri07b] Lawrence M. Principe. A revolution nobody noticed? Changes in early Eighteenth-Century chymistry. In Lawrence M. Principe, editor, New Narratives in Eighteenth-Century Chemistry: Contributions from the First Francis Bacon Workshop, 21–23 April 2005, California Institute of Technology, Pasadena, California, volume 18 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–22. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2007. ISBN 1-4020-6278-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-1-4020-6278-0_1.

Principe:2014:GCA

[Pri14] Lawrence M. Principe. Goldsmiths and chymists: The activity of artisans within alchemical circles. In Sven Dupré, editor, Laboratories of Art: Alchemy and Art Technology from Antiquity to the 18th Century, volume 37 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 157–179. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05065-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-05065-2_7.

Probst:2015:LRS

[Pro15a] Siegmund Probst. Leibniz as reader and second inventor: The cases of Barrow and Mengoli. In Norma B. Goethe, Philip Beeley, and David Rabouin, editors, G. W. Leibniz, Interrelations between Mathematics and Philosophy, volume 41 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 111– 134. Springer Netherlands, Dordrecht, The Netherlands, 2015. ISBN 94-017-9664-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-017-9664-4_6.

Proust:2015:TSL

[Pro15b] Christine Proust. A tree-structured list in a mathematical series text from Mesopotamia. In Karine Chemla and Jacques Virbel, editors, Texts, Textual Acts and the History of Science, volume 42 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 281–316. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-16444-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-16444-1_8.

Proust:2016:MPI

[Pro16] Christine Proust. Mathematical and philological insights on cuneiform texts. Neugebauer’s correspondence with fellow Assyriologists. In Alexander Jones, Christine Proust, and John M. Steele, editors, A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science, volume 45 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 207–235. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-25865-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-25865-2_7.

Rabouin:2015:DBS

[Rab15] David Rabouin. The difficulty of being simple: On some interactions between mathematics and Philosophy in Leibniz’s analysis of notions. In Norma B. Goethe, Philip Beeley, and David Rabouin, editors, G. W. Leibniz, Interrelations between Mathematics and Philosophy, volume 41 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 49–72. Springer Netherlands, Dordrecht, The Netherlands, 2015. ISBN 94-017-9664-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-94-017-9664-4_3.

Ragep:2010:IRP

[Rag10] F. Jamil Ragep. Islamic reactions to Ptolemy’s imprecisions. In Alexander Jones, editor, Ptolemy in Perspective: Use and Criticism of his Work from Antiquity to the Nineteenth Century, volume 23 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 121–134. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-2788-2. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-2788-7_5.

Rasch:2006:SMK

[Ras06] Manfred Rasch. Science and the military: The Kaiser Wilhelm Foundation for Military–Technical Science. In Roy Macleod and Jeffrey Allan Johnson, editors, Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 179–202. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5490-4. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-5490-4_11.

Raynaud:2016:BM

[Ray16a] Dominique Raynaud. Back matter. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 215–297. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-3-319-42721-8/1.pdf.

Raynaud:2016:MHR

[Ray16b] Dominique Raynaud. De Mesa’s hypothesis regarding the arithmetic construction of perspective. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 191–213. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-42721-8_11.

Raynaud:2016:FFR

[Ray16c] Dominique Raynaud. Fact and fiction regarding Masaccio’s trinity fresco. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 53–67. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-42721-8_4.

Raynaud:2016:FMa

[Ray16d] Dominique Raynaud. Front matter. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i– xi. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/content/pdf/
                                        bfm:978-3-319-42721-8/1.pdf.

Raynaud:2016:FMb

[Ray16e] Dominique Raynaud. Front matter. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 13. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/content/
pdf/bfm:978-3-319-42721-8/1/
1.pdf.

Raynaud:2016:FMc

[Ray16f] Dominique Raynaud. Front matter. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 69. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/content/
pdf/bfm:978-3-319-42721-8/2/
1.pdf.

Raynaud:2016:FMd

[Ray16g] Dominique Raynaud. Front matter. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 131. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/content/
pdf/bfm:978-3-319-42721-8/3/
1.pdf.

Raynaud:2016:HPC

[Ray16h] Dominique Raynaud. The Hauck– Panofsky Conjecture regarding curvilinear perspective. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 161–175. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-42721-8_9.

Raynaud:2016:IAH

[Ray16i] Dominique Raynaud. Ibn al-Haytham on binocular vision. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 71–93. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-42721-8_5.

Raynaud:2016:KBR

[Ray16j] Dominique Raynaud. Knowledge and beliefs regarding linear perspective. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 15–35. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-42721-8_2.

Raynaud:2016:LIA

[Ray16k] Dominique Raynaud. The legacy of Ibn al-Haytham. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 95–114. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-42721-8_6.

Raynaud:2016:PNA

[Ray16l] Dominique Raynaud. Perspectiva naturalis/artificialis. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–12. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-42721-8_1.

Raynaud:2016:PTP

[Ray16m] Dominique Raynaud. The properties of two-point perspective. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 133–159. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-42721-8_8.

Raynaud:2016:RTP

[Ray16n] Dominique Raynaud. The rejection of the two-point perspective system. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 115–129. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-42721-8_7.

Raynaud:2016:SBV

[Ray16o] Dominique Raynaud. Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2016. ISBN 3-319-42720-2 (hardcover), 3-319-42721-0 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xi + 297 pp. LCCN QP487 .R396 2016.

Raynaud:2016:UEP

[Ray16p] Dominique Raynaud. Understanding errors in perspective. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 37– 52. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-42721-8_3.

Raynaud:2016:WCC

[Ray16q] Dominique Raynaud. The White– Carter Conjecture on synthetic perspective. In Studies on Binocular Vision: Optics, Vision and Perspective from the Thirteenth to the Seventeenth Centuries, volume 47 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 177–189. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-42721-0. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-42721-8_10.

Renn:2003:HCF

[RD03] Jürgen Renn and Peter Damerow.

The hanging chain: A forgotten “discovery” buried in Galileo’s notes on Mition. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–24. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/0-306-48152-9_1.

Rheinberger:2003:CCE

[Rhe03] Hans-Jörg Rheinberger. Carl Correns’ experiments with Pisum, 1896–1899. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 221–252. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/0-306-48152-9_11.

Richardson:2006:FSS

[Ric06] Alan Richardson. Freedom in a scientific society: Reading the context of Reichenbach’s contexts. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 41–54. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-4251-5_4.

Ritter:2016:ONA

[Rit16] Jim Ritter. Otto Neugebauer and Ancient Egypt. In Alexander Jones, Christine Proust, and John M. Steele, editors, A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science, volume 45 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 127–163. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-25865-6. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-25865-2_4.

Robadey:2015:WDG

[Rob15] Anne Robadey. A work on the degree of generality revealed in the organization of enumerations: Poincaré’s classification of singular points of differential equations. In Karine Chemla and Jacques Virbel, editors, Texts, Textual Acts and the History of Science, volume 42 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 385–419. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-16444-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-16444-1_10.

Roland:2000:PIR

[Rol00] Alex Roland. Pools of invention: The role of patents in the development of American aircraft, 1917–1997. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 13, pages 323– 345. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-011-4379-0_13.

Rosinska:2006:MAU

[Ros06] Grazyna Rosińska. “Mathematics for astronomy” at universities in Copernicus’ time: Modern attitudes toward ancient problems. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 9–28. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/1-4020-3975-1_2.

Rothman:2011:CCK

[Rot11] Aviva Rothman. From cosmos to confession: Kepler and the connection between astronomical and religious truth. In Patrick J. Boner, editor, Change and Continuity in Early Modern Cosmology, volume 27 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 115–133. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 94-007-0037-7. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-0037-6_7.

Rowe:2016:GGN

[Row16] David E. Rowe. From Graz to Göttingen: Neugebauer’s early intellectual journey. In Alexander Jones, Christine Proust, and John M. Steele, editors, A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science, volume 45 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–59. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-25865-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-25865-2_1.

Renn:2003:EIR

[RS03] Jürgen Renn and Tilman Sauer.

Errors and insights: Reconstructing the genesis of general relativity from Einstein’s Zürich Notebook. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 253–268. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/0-306-48152-9_12.

Rupke:2012:FDB

[Rup12] Nicolaas A. Rupke. Five discourses of Bible and science 1750–2000. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 179–195. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-2627-7_11.

Rutkin:2010:UAP

[Rut10] H. Darrel Rutkin. The use and abuse of Ptolemy’s Tetrabiblos in Renaissance and Early Modern Europe: Two case studies (Giovanni Pico della Mirandola and Filippo Fantoni). In Alexander Jones, editor, Ptolemy in Perspective: Use and Criticism of his Work from Antiquity to the Nineteenth Century, volume 23 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 135–149. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-2788-2. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-2788-7_6.

Rutkin:2019:ATA

[Rut19a] H. Darrel Rutkin. Astrology and theology in Albertus Magnus and Thomas Aquinas: Fate, divination and Providence. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], chapter 5, pages 173– 234. CODEN ???? ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN ???? LCCN B108-5802; QB25. URL https://

link.springer.com/chapter/
10.1007/978-3-030-10779-6_5.

Rutkin:2019:ATR

[Rut19b] H. Darrel Rutkin. Astrology and theology/religion in Roger Bacon. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], chapter 6, pages 235– 270. CODEN ???? ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN ???? LCCN B108-5802; QB25. URL https://

link.springer.com/chapter/
10.1007/978-3-030-10779-6_6.

Rutkin:2019:ASP

[Rut19c] H. Darrel Rutkin. Astrology in society, politics and culture. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], chapter 11, pages 423–463. CO-DEN ???? ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN ???? LCCN B108-

5802; QB25. URL https://link.
springer.com/chapter/10.
1007/978-3-030-10779-6_11.

Rutkin:2019:BM

[Rut19d] H. Darrel Rutkin. Back matter. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], pages 465– 515. ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN 1385-0180 (print), 2215-0064 (electronic). LCCN B108-5802; QB25.

Rutkin:2019:DAR

[Rut19e] H. Darrel Rutkin. Defending astrology: Roger Bacon and the Speculum Astronomiae. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], chapter 4, pages 117–167. CO-DEN ???? ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN ???? LCCN B108-

                                        5802; QB25. URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-10779-6_4.

Rutkin:2019:DCA

[Rut19f] H. Darrel Rutkin. Disciplinary configurations: Astrology and the mathematical disciplines. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], chapter 9, pages 367–383. CO-DEN ???? ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN ???? LCCN B108-

                                        5802; QB25. URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-030-10779-6_9.

Rutkin:2019:FMa

[Rut19g] H. Darrel Rutkin. Front matter. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], pages i–lxxxix. ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN 1385-0180 (print), 2215-0064 (electronic). LCCN B108-5802; QB25. URL https:

//link.springer.com/book/10.
1007/978-3-030-10779-6.

Rutkin:2019:FMb

[Rut19h] H. Darrel Rutkin. Front matter. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], pages 1–3. ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN 1385-0180 (print), 2215-0064 (electronic). LCCN B108-5802; QB25.

Rutkin:2019:FMc

[Rut19i] H. Darrel Rutkin. Front matter. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], pages 169– 171. ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN 1385-0180 (print), 2215-0064 (electronic). LCCN B108-5802; QB25.

Rutkin:2019:FMd

[Rut19j] H. Darrel Rutkin. Front matter. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], pages 271– 273. ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN 1385-0180 (print), 2215-0064 (electronic). LCCN B108-5802; QB25.

Rutkin:2019:FMe

[Rut19k] H. Darrel Rutkin. Front matter. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], pages 363– 365. ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN 1385-0180 (print), 2215-0064 (electronic). LCCN B108-5802; QB25.

Rutkin:2019:IAT

[Rut19l] H. Darrel Rutkin. Imagines astronomicae (talismans) in the Speculum astronomiae, Albertus Magnus and Thomas Aquinas. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], chapter 7, pages 275–315. CO-DEN ???? ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN ???? LCCN B108-

5802; QB25. URL https://link.
springer.com/chapter/10.
1007/978-3-030-10779-6_7.

Rutkin:2019:IFU

[Rut19m] H. Darrel Rutkin. Institutional foundations: The universities. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], chapter 10, pages 385–421. CO-DEN ???? ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN ???? LCCN B108-

5802; QB25. URL https://link.
springer.com/chapter/10.
1007/978-3-030-10779-6_10.

Rutkin:2019:LNP

[Rut19n] H. Darrel Rutkin. Ligamentum naturalis philosophiae et metaphysicae: Astrology and Aristotelian natural philosophy. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], chapter 1, pages 5–39. CODEN ???? ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN ???? LCCN B108-

5802; QB25. URL https://link.
springer.com/chapter/10.
1007/978-3-030-10779-6_1.

Rutkin:2019:MPM

[Rut19o] H. Darrel Rutkin. Mathematizing the picture: Mathematics, perspectiva and astrology in Roger Bacon and Albertus Magnus. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], chapter 2, pages 41–92. CO-DEN ???? ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN ???? LCCN B108-

5802; QB25. URL https://link.
springer.com/chapter/10.
1007/978-3-030-10779-6_2.

Rutkin:2019:NPF

[Rut19p] H. Darrel Rutkin. The natural philosophical foundations for astrological revolutions: Albertus Magnus’s commentary on the pseudo-Aristotelian De Causis Proprietatum Elementorum. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], chapter 3, pages 93–115. CODEN ???? ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN ???? LCCN B108-

5802; QB25. URL https://link.
springer.com/chapter/10.
1007/978-3-030-10779-6_3.

Rutkin:2019:OVS

[Rut19q] H. Darrel Rutkin. Opera et verba sapientiae: Astrology and magic in Roger Bacon. In Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800 [Rut19r], chapter 8, pages 317– 362. CODEN ???? ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN ???? LCCN B108-5802; QB25. URL https://

link.springer.com/chapter/
10.1007/978-3-030-10779-6_8.

Rutkin:2019:SAA

[Rut19r] H. Darrel Rutkin. Sapientia Astrologica: Astrology, Magic and Natural Knowledge, ca. 1250–1800, volume 55 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2019. ISBN 3-030-10778-7, 3-030-10779-5 (e-book), 3-030-10780-9. ISSN 1385-0180 (print), 2215-0064 (electronic). lxxxix + 515 + 9 pp. LCCN B108-5802; QB25.

Simoes:1999:CKE

[SCD99] Ana Simões, Ana Carneiro, and Maria Paula Diogo. Constructing knowledge: Eighteenth-Century Portugal and the new sciences. In Kostas Gavroglu, editor, The Sciences in the European Periphery During the Enlightenment, volume 2 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–40. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 94-011-4770-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-94-011-4770-5_1.

Schatzberg:2000:WAW

[Sch00] Eric Schatzberg. Wooden airplanes in World War II: National comparisons and symbolic culture. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 183–205. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-011-4379-0_6.

Schafer:2006:AVD

[Sch06a] Lothar Schäfer. Autonomy versus development: Duhem on progress in science. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 79–97. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-4251-5_6.

Schickore:2006:FPC

[Sch06b] Jutta Schickore. A forerunner? — perhaps, but not to the context distinction. William Whewell’s Germano–Cantabrigian history of the fundamental ideas. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 57–77. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-4251-5_5.

Schiemann:2006:IJD

[Sch06c] Gregor Schiemann. Inductive justification and discovery. On Hans Reichenbach’s foundation of the autonomy of the philosophy of science. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 23–39. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-4251-5_3.

Schiemann:2009:HHM

[Sch09] Gregor Schiemann, editor. Hermann von Helmholtz’s Mechanism: The Loss of Certainty: a Study on the Transition from Classical to Modern Philosophy of Nature, volume 17 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2009. ISBN 1-4020-5629-X (hardcover), 1-4020-5630-3 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). x + 282 pp. LCCN B3279.H54 S3513 2009.

Schmaus:2010:UEF

[Sch10] Warren Schmaus. Understanding and explanation in France: From Maine de Biran’s Méthode Psychologique to Durkheim’s Les Formes Élémentaires de la vie Religieuse. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 101–120. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-3540-0_6.

Schabas:2017:MFT

[Sch17] Margaret Schabas. More food for thought: Mill, Coleridge and the dismal science of economics. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 151– 162. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-58436-2_9.

Sturm:2006:HCW

[SG06] Thomas Sturm and Gerd Gigerenzer. How can we use the distinction between discovery and justification? On the weaknesses of the strong programme in the sociology of science. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 133–158. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-4251-5_9.

Shapiro:2003:NON

[Sha03] Alan E. Shapiro. Newton’s optical notebooks: Public versus private data. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 43–65. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/0-306-48152-9_3.

Shapiro:2005:SEN

[Sha05] Alan E. Shapiro. Skating on the edge: Newton’s investigation of chromatic dispersion and achromatic prisms and lenses. In Jed Z. Buchwald and Allan Franklin, editors, Wrong for the Right Reasons, volume 11 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 99–125. Springer, Dordrecht, The Netherlands, 2005. ISBN 1-4020-3048-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3048-7_6.

Sibum:2003:NNK

[Sib03] H. Otto Sibum. Narrating by numbers: Keeping an account of Early 19th Century laboratory experiences. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 141– 158. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/0-306-48152-9_7.

Siraisi:2017:LL

[Sir17] Nancy Siraisi. A life of learning. In Gideon Manning and Cynthia Klestinec, editors, Professors, Physicians and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, volume 50 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 15–28. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56514-1. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-56514-9_2.

Sivin:2015:BM

[Siv15a] Nathan Sivin. Back matter. In Health Care in Eleventh-Century China, volume 43 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 193–223. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-20427-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bbm:978-3-319-20427-7/1.pdf.

Sivin:2015:CM

[Siv15b] Nathan Sivin. Classical medicine. In Health Care in Eleventh-Century China, volume 43 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 53–91. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-20427-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-20427-7_4.

Sivin:2015:C

[Siv15c] Nathan Sivin. Conclusions. In Health Care in Eleventh-Century China, volume 43 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 183–192. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-20427-0. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-20427-7_8.

Sivin:2015:FM

[Siv15d] Nathan Sivin. Front matter. In Health Care in Eleventh-Century China, volume 43 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xvii. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-20427-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/content/pdf/
                                        bfm:978-3-319-20427-7/1.pdf.

Sivin:2015:HCE

[Siv15e] Nathan Sivin. Health Care in Eleventh-Century China, volume 43 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2015. ISBN 3-319-20426-2 (hardcover), 3-319-20427-0 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xvii + 223 pp. LCCN RA527.

Sivin:2015:HCM

[Siv15f] Nathan Sivin. Health care, medicine, and Chinese society. In Health Care in Eleventh-Century China, volume 43 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–8. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-20427-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-20427-7_1.

Sivin:2015:QE

[Siv15g] Nathan Sivin. The question of efficacy. In Health Care in Eleventh-Century China, volume 43 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 31–52. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-20427-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-20427-7_3.

Sivin:2015:SDV

[Siv15h] Nathan Sivin. Some definitions and viewpoints. In Health Care in Eleventh-Century China, volume 43 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 9–30. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-20427-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-20427-7_2.

Sivin:2015:TER

[Siv15i] Nathan Sivin. Therapy in elite religions. In Health Care in Eleventh-Century China, volume 43 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 129–170. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-20427-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-20427-7_6.

Sivin:2015:TPR

[Siv15j] Nathan Sivin. Therapy in popular religion. In Health Care in Eleventh-Century China, volume 43 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 93–128. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-20427-0. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-20427-7_5.

Sivin:2015:TSR

[Siv15k] Nathan Sivin. Therapy in the state religion. In Health Care in Eleventh-Century China, volume 43 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 171–181. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-20427-0. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-20427-7_7.

Smith:2000:ETE

[SM00] George E. Smith and David A.

Mindell. The emergence of the turbofan engine. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 107–155. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-011-4379-0_4.

Sendra-Mocholi:2006:BUS

[SM06] Cristina Sendra-Mocholí. Botany in university studies in the late 18th Century. The case of Valencia University. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 17, pages 259–272. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3975-1_17.

Smith:2005:WWN

[Smi05] George E. Smith. Was wrong Newton bad Newton? In Jed Z. Buchwald and Allan Franklin, editors, Wrong for the Right Reasons, volume 11 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 127–160. Springer, Dordrecht, The Netherlands, 2005. ISBN 1-4020-3048-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-3048-7_7.

Smith:2006:OTW

[Smi06] John Kenly Smith, Jr. Organizing for total war: DuPont and smokeless powder in World War I. In Roy Macleod and Jeffrey Allan Johnson, editors, Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 167–178. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5490-4. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-5490-4_10.

Smith:2010:BTR

[Smi10] Roger Smith. British thought on the relations between the natural sciences and the humanities, c. 1870–1910. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 161–185. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.

springer.com/chapter/10.
1007/978-90-481-3540-0_9.

Smith:2017:GDP

[Smi17] A. Mark Smith. Giambattista Della Porta’s theory of vision in the De refractione of 1593: Sources, problems, implications. In Arianna Borrelli, Giora Hon, and Yaakov Zik, editors, The Optics of Giambattista Della Porta (ca. 1535–1615): A Reassessment, volume 44 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 97– 123. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-50215-8. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-50215-1_5.

Sowards:2015:RFD

[Sow15] Adam M. Sowards. Rexford F.

Daubenmire and the ecology of place: Applied ecology in the Mid-Twentieth-Century American West. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 15, pages 297–322. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-12185-7_15.

Schickore:2006:IRC

[SS06a] Jutta Schickore and Friedrich Steinle. Introduction: Revisiting the context distinction. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 7–19. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-4251-5_1.

Schickore:2006:RDJ

[SS06b] Jutta Schickore and Friedrich Steinle, editors. Revisiting Discovery and Justification: Historical and Philosophical Perspectives on the Context Distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2006. ISBN 1-4020-4250-7 (hardcover), 1-4020-4251-5 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xix + 232 pp. LCCN Q175 .R418 2006. URL

                                        https://www.springer.com/us/
                                        book/9781402048197.

Siegmund-Schultze:2016:PWO

[SS16]  Reinhard  Siegmund-

Schultze. “Not in possession of any Weltanschauung”: Otto Neugebauer’s flight from Nazi Germany and his search for objectivity in mathematics, in reviewing, and in history. In Alexander Jones, Christine Proust, and John M. Steele, editors, A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science, volume 45 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 61–106. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-25865-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-25865-2_2.

Steele:2000:BM

[Ste00a] John M. Steele. Back matter. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 293–323. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-94-015-9528-5/1.pdf.

Steele:2000:C

[Ste00b] John M. Steele. China. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 161–215. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-015-9528-5_6.

Steele:2000:CSS

[Ste00c] John M. Steele. Chinese sunrise and sunset times. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 263–267. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-015-9528-5_10.

Steele:2000:D

[Ste00d] John M. Steele. Discussion. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 231–236. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-015-9528-5_8.

Steele:2000:FMa

[Ste00e] John M. Steele. Front matter. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i– xi. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/content/pdf/
bfm:978-94-015-9528-5/1.pdf.

Steele:2000:FMb

[Ste00f] John M. Steele. Front matter. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 1. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-015-9528-5/1/1.
pdf.

Steele:2000:FMc

[Ste00g] John M. Steele. Front matter. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 19. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-015-9528-5/2/1.
pdf.

Steele:2000:FMd

[Ste00h] John M. Steele. Front matter. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 159. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-015-9528-5/3/1.
pdf.

Steele:2000:FMe

[Ste00i] John M. Steele. Front matter. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 229. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-015-9528-5/4/1.
pdf.

Steele:2000:FMf

[Ste00j] John M. Steele. Front matter. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, page 237. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-94-015-9528-5/5/1.
pdf.

Steele:2000:GRW

[Ste00k] John M. Steele. The Greco– Roman world. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 85–105. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-015-9528-5_3.

Steele:2000:I

[Ste00l] John M. Steele. Introduction. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 3–17. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-015-9528-5_1.

Steele:2000:INE

[Ste00m] John M. Steele. The Islamic Near East. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 107–124. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-015-9528-5_4.

Steele:2000:J

[Ste00n] John M. Steele. Japan. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 217–228. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.

springer.com/chapter/10.
1007/978-94-015-9528-5_7.

Steele:2000:LBE

[Ste00o] John M. Steele. Late Babylonian eclipse records. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 239–262. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-015-9528-5_9.

Steele:2000:LMR

[Ste00p] John M. Steele. Late Medieval and Renaissance Europe. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 125–158. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-015-9528-5_5.

Steele:2000:M

[Ste00q] John M. Steele. Mesopotamia.

In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 21–83. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-015-9528-5_2.

Steele:2000:OPE

[Ste00r] John M. Steele. Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2000. ISBN 90-481-5454-5 (hardcover), 94-015-9528-3 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xi + 323 pp. LCCN ????

Steele:2000:TCE

[Ste00s] John M. Steele. Timed Chinese eclipse records. In Observations and Predictions of Eclipse Times by Early Astronomers, volume 4 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 269–292. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2000. ISBN 94-015-9528-3. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-015-9528-5_11.

Steinle:2003:PSP

[Ste03] Friedrich Steinle. The practice of studying practice: Analyzing research records of Ampère and Faraday. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 93–118. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/0-306-48152-9_5.

Steen:2006:TEU

[Ste06a] Kathryn Steen. Technical expertise and U.S. mobilization, 1917– 18: High explosives and war gases. In Roy Macleod and Jeffrey Allan Johnson, editors, Frontline and Factory: Comparative Perspectives on the Chemical Industry at War, 1914–1924, volume 16 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 103–122. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-5490-4. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-5490-4_7.

Steinle:2006:CFL

[Ste06b] Friedrich Steinle. Concept formation and the limits of justification: “discovering” the two electricities. In Jutta Schickore and Friedrich Steinle, editors, Revisiting Discovery and Justification: Historical and philosophical perspectives on the context distinction, volume 14 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 11, pages 183–195. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-4251-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-4251-5_11.

Steele:2010:DML

[Ste10] J. M. Steele. Dunthorne, Mayer, and Lalande on the secular acceleration of the moon. In Alexander Jones, editor, Ptolemy in Perspective: Use and Criticism of his Work from Antiquity to the Nineteenth Century, volume 23 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 203–215. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-2788-2. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-2788-7_8.

Steele:2016:NAC

[Ste16] John M. Steele. Neugebauer’s astronomical cuneiform texts and its reception. In Alexander Jones, Christine Proust, and John M. Steele, editors, A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science, volume 45 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 303–332. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2016. ISBN 3-319-25865-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/978-3-319-25865-2_10.

Stewart:2017:MEB

[Ste17] Larry Stewart. At the medical edge or, the Beddoes Effect. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 47–64. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-58436-2_3.

Suber:2010:SSB

[Šub10] Daniel Šuber. Social science between neo-Kantianism and philosophy of life: The cases of Weber, Simmel, and Mannheim. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 14, pages 267–290. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-3540-0_14.

Suppe:2000:CNF

[Sup00] Frederick Suppe. The changing nature of flight and ground test instrumentation and data: 1940–1969. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 67– 105. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-011-4379-0_3.

Swerdlow:2005:PTL

[Swe05] N. M. Swerdlow. Ptolemy’s theories of the latitude of the planets in the Almagest, handy tables, and planetary hypotheses. In Jed Z. Buchwald and Allan Franklin, editors, Wrong for the Right Reasons, volume 11 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 41–71. Springer, Dordrecht, The Netherlands, 2005. ISBN 1-4020-3048-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/1-4020-3048-7_3.

Swerdlow:2010:TLK

[Swe10] N.  M.  Swerd-

low. Tycho, Longomontanus, and Kepler on Ptolemy’s solar observations and theory, precession of the equinoxes, and obliquity of the ecliptic. In Alexander Jones, editor, Ptolemy in Perspective: Use and Criticism of his Work from Antiquity to the Nineteenth Century, volume 23 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 151–202. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-2788-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-2788-7_7.

Swerdlow:2012:UPT

[Swe12] N. M. Swerdlow. Urania Propitia, Tabulae Rudophinae faciles redditae a Maria Cunitia: Beneficent Urania, the Adaptation of the Rudolphine Tables by Maria Cunitz. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 81–121. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-2627-7_7.

Teboul:2015:ESC

[Teb15] Michel Teboul. The enumeration structure of [Chinese characters] Ěryǎ’s “semantic lists”. In Karine Chemla and Jacques Virbel, editors, Texts, Textual Acts and the History of Science, volume 42 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 267–280. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-16444-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-16444-1_7.

Tessicini:2011:CGN

[Tes11] Dario Tessicini. Cornelius Gemma and the new star of 1572. In Patrick J. Boner, editor, Change and Continuity in Early Modern Cosmology, volume 27 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 51–65. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 94-007-0037-7. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-0037-6_4.

Tihon:2010:UAP

[Tih10] Anne Tihon. An unpublished astronomical papyrus contemporary with Ptolemy. In Alexander Jones, editor, Ptolemy in Perspective: Use and Criticism of his Work from Antiquity to the Nineteenth Century, volume 23 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–10. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-2788-2. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-2788-7_1.

Todes:2003:LIL

[Tod03] Daniel P. Todes. From lone investigator to laboratory chief: Ivan Pavlov’s research notebooks as a reflection of his managerial and interpretive style. In Frederic L. Holmes, Jürgen Renn, and Hans-Jörg Rheinberger, editors, Reworking the Bench: Research Notebooks in the History of Science, volume 7 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 203– 220. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-306-48152-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/0-306-48152-9_10.

Treitel:2015:ABN

[Tre15] Corinna Treitel. Artificial or biological? Nature, fertilizer, and the German origins of organic agriculture. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 183–203. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-12185-7_10.

Uebel:2010:OVO

[Ueb10] Thomas Uebel. Opposition to Verstehen in orthodox logical empiricism. In Uljana Feest, editor, Historical Perspectives on Erklären and Verstehen, volume 21 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 15, pages 291–309. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2010. ISBN 90-481-3540-0. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-3540-0_15.

Vermij:2011:JPH

[Ver11] Rienk Vermij. Johannes Phocylides Holwarda and the interpretation of new stars in the Dutch Republic. In Patrick J. Boner, editor, Change and Continuity in Early Modern Cosmology, volume 27 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 135–152. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 94-007-0037-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-94-007-0037-6_8.

Vetter:2015:RKE

[Vet15]  Jeremy Vetter.  Regionalizing

knowledge: The ecological approach of the USDA Office of Dryland Agriculture on the Great Plains. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 14, pages 277–296. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-12185-7_14.

Vincenti:2000:EEE

[Vin00] Walter G. Vincenti. Engineering experiment and engineering theory: The aerodynamics of wings at supersonic speeds, 1946–1948. In Peter Galison and Alex Roland, editors, Atmospheric Flight in the Twentieth Century, volume 3 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 157–180. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 94-011-4379-X. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-011-4379-0_5.

Virbel:2015:SAT

[Vir15a] Jacques Virbel. Speech act theory and instructional texts. In Karine Chemla and Jacques Virbel, editors, Texts, Textual Acts and the History of Science, volume 42 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 49–86. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-16444-9. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

                                        link.springer.com/chapter/
                                        10.1007/978-3-319-16444-1_2.

Virbel:2015:TE

[Vir15b] Jacques Virbel. Textual enumeration. In Karine Chemla and Jacques Virbel, editors, Texts, Textual Acts and the History of Science, volume 42 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 221–266. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015.

ISBN 3-319-16444-9. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-16444-1_6.

Wesley:2014:IRT

[Wes14] Morgan Wesley. Igne refutata:

Thermal analysis in the laboratory practices of John Dwight and Ehrenfried Walther von Tschirnhaus. In Sven Dupré, editor, Laboratories of Art: Alchemy and Art Technology from Antiquity to the 18th Century, volume 37 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 181–200. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05065-6. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://

link.springer.com/chapter/
10.1007/978-3-319-05065-2_8.

Wilson:2012:SCR

[Wil12] David B. Wilson. Simplicity in the Copernican revolution: Galileo, Descartes, Newton. In Jed Z. Buchwald, editor, A Master of Science History: Essays in Honor of Charles Coulston Gillispie, volume 30 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 8, pages 123–133. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2012. ISBN 94-007-2627-9. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-94-007-2627-7_8.

Wille:2015:CSM

[Wil15] Robert-Jan Wille. The coproduction of station morphology and agricultural management in the tropics: Transformations in botany at the Botanical Garden at Buitenzorg, Java 1880–1904. In Denise Phillips and Sharon Kingsland, editors, New Perspectives on the History of Life Sciences and Agriculture, volume 40 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 13, pages 253–275. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2015. ISBN 3-319-12185-5. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

                                        springer.com/chapter/10.
                                        1007/978-3-319-12185-7_13.

Wood:2006:CCC

[Woo06] Paul Wood. Candide in Caledonia: The culture of science in the Scottish universities, 1690– 1805. In Mordechai Feingold and Victor Navarro-Brotons, editors, Universities and Science in the Early Modern Period, volume 12 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 13, pages 183–199. Springer, Dordrecht, The Netherlands, 2006. ISBN 1-4020-3975-1. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.

springer.com/chapter/10.
1007/1-4020-3975-1_13.

Wuthrich:2011:BM

[Wüt11a] Adrian Wüthrich. Back matter. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages 191–208. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bbm:978-90-481-9228-1/1.pdf.

Wuthrich:2011:DNM

[Wüt11b] Adrian Wüthrich. The development of a new means of representation: Goals and milestones. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 7, pages 177–189. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-9228-1_7.

Wuthrich:2011:DEF

[Wüt11c] Adrian Wüthrich. The Dirac equation: Feynman’s great struggle. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 4, pages 65–111.

Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-9228-1_4.

Wuthrich:2011:FTS

[Wüt11d] Adrian Wüthrich. The field theoretical systematization of Feynman’s theory. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 6, pages 141–175. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-9228-1_6.

Wuthrich:2011:FPS

[Wüt11e] Adrian Wüthrich. Free propagation and successive scattering. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 5, pages 113–139. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-9228-1_5.

Wuthrich:2011:FM

[Wüt11f] Adrian Wüthrich. Front matter. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, pages i–xvii. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/content/pdf/
bfm:978-90-481-9228-1/1.pdf.

Wuthrich:2011:GFD

[Wüt11g] Adrian Wüthrich. The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2011. ISBN 90-481-9227-7 (hardcover), 90-481-9228-5 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xvii + 208 pp. LCCN QC794.6.F4 W88 2010.

Wuthrich:2011:IOU

[Wüt11h] Adrian Wüthrich. Introduction:

Origin, use and interpretation of Feynman diagrams. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 1, pages 1–21. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-90-481-9228-1_1.

Wuthrich:2011:QEF

[Wüt11i] Adrian Wüthrich. Quantum electrodynamics without Feynman diagrams. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 2, pages 23–50. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-9228-1_2.

Wuthrich:2011:QMH

[Wüt11j] Adrian Wüthrich. Quantum mechanics without a Hamiltonian operator. In The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 51–64. Springer Science+Business Media B.V., Dordrecht, The Netherlands, 2011. ISBN 90-481-9228-5. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-90-481-9228-1_3.

Yeomans:2014:JAS

[YKS14] David Yeomans, Jason M. Kelly, and Frank Salmon. James “Athenian” Stuart and the geometry of setting out. In Anthony Gerbino, editor, Geometrical Objects: Architecture and the Mathematical Sciences 1400–1800, volume 38 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 12, pages 281–312. Springer International Publishing Switzerland, Cham, Switzerland; Heidelberg, Germany; New York, NY, USA; Dordrecht, The Netherlands; London, UK, 2014. ISBN 3-319-05998-X. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-05998-3_12.

Zeller:2017:CCC

[Zel17] Suzanne Zeller. Context, connections and culture: The history of science in Canada as a field of study. In Jed Buchwald and Larry Stewart, editors, The Romance of Science: Essays in Honour of Trevor H. Levere, volume 52 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 15, pages 277– 299. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-58436-7. ISSN 1385-0180 (print), 2215-0064 (elec-

tronic). URL https://link.
springer.com/chapter/10.
1007/978-3-319-58436-2_15.

Zik:2017:CR

[ZH17a] Yaakov Zik and Giora Hon. Conclusion: A reassessment. In Arianna Borrelli, Giora Hon, and Yaakov Zik, editors, The Optics of Giambattista Della Porta (ca. 1535–1615): A Reassessment, volume 44 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 9, pages 201– 205. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-50215-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-50215-1_9.

Zik:2017:GDP

[ZH17b] Yaakov Zik and Giora Hon. Giambattista Della Porta: A magician or an optician? In Arianna Borrelli, Giora Hon, and Yaakov Zik, editors, The Optics of Giambattista Della Porta (ca. 1535– 1615): A Reassessment, volume 44 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 3, pages 39–55. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-50215-8. ISSN 1385-0180 (print), 2215-0064 (elec-

                                        tronic). URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-3-319-50215-1_3.